### v2

# Transfer Learning Between U.S. Presidential Elections: How Should We Learn From A 2020 Ad Campaign To Inform 2024 Ad Campaigns?

Abstract
For the 2024 U.S. presidential election, would negative, digital ads against Donald Trump impact voter turnout in Pennsylvania (PA), a key “tipping point” state? The gold standard to address this question, a randomized experiment where voters get randomized to different ads, yields unbiased estimates of the ad effect, but is very expensive. Instead, we propose a less-than-ideal, but significantly cheaper and faster framework based on transfer learning, where we transfer knowledge from a past ad experiment in 2020 to evaluate ads for 2024. A key component of our framework is a sensitivity analysis that quantifies the unobservable differences between 2020 and 2024 elections, where sensitivity parameters can be calibrated in a data-driven manner. We propose two estimators of the 2024 ad effect: a simple regression estimator with bootstrap, which we recommend for practitioners in this field, and an estimator based on the efficient influence function for broader applications. Using our framework, we estimate the effect of running a negative, digital ad campaign against Trump on voter turnout in PA for the 2024 election. Our findings indicate effect heterogeneity across counties of PA and among important subgroups stratified by gender, urbanicity, and education attainment.
Keywords: Causal inference; Generalizability; Sensitivity analysis; Transportability; Exponential tilting
1 Introduction
1.1 Motivation: Learning from past ad campaigns
In recent years, political campaigns have used randomized experiments to evaluated political ads (e.g., Gerber et al., (2011); Kalla and Broockman, (2018); Aggarwal et al., (2023)). For example, in the 2020 U.S. presidential election, Aggarwal et al., (2023) conducted a large-scale, randomized controlled trial (RCT) among $1,999,282$ registered voters from Pennsylvania (PA), Wisconsin (WI), Michigan (MI), North Carolina (NC) and Arizona (AZ). They found that a negative, digital ad campaign against President Donald Trump during the 2020 U.S. presidential election was ineffective in changing voter turnout.
The main empirical question we address in this paper is as follows: similar to 2020, would negative digital ads against Trump remain ineffective in changing voter turnout for the 2024 U.S. presidential election? In both 2020 and 2024 elections, Trump was the nominee for the Republican party and digital anti-Trump ads were used extensively. But, as Aggarwal et al., (2023) pointed out, 2020 was an exceptional election due to COVID-19 and their null results may not generalize to less exceptional elections. In particular, compared to 2020, voters faced new issues in 2024 including women’s rights, inflation, the Russo-Ukraine War and the Israel-Hamas War (Pew Research Center,, 2024; Ipsos Core Political,, 2024). Also, Loving and Smith, (2024) showed that after the attack on the U.S. Capitol Building on January 6, 2021, some Republican voters left their party, signaling a potential shift in voter demographics between the 2020 and the 2024 elections. In short, there were measurable and unmeasurable differences between 2020 and 2024 in terms of electoral contexts and voter demographics.
The ideal approach to answer the main empirical question is to re-run the randomized experiment by Aggarwal et al., (2023). While this approach can yield unbiased estimates of the ad effect irrespective of the differences between 2020 and 2024, ad campaigns are very expensive. For example, Aggarwal et al., (2023)’s experiment in 2020 costed \$8.9 million U.S. dollars (USD). More recently, one super political action committee for the Democratic Party, which was characterized as “an ad-making laboratory…testing thousands of messages, social media posts and ads in the 2024 race, ranking them in order of effectiveness”, spent \$450 million USD for the 2024 election (Schleifer and Goldmacher,, 2024).
Our approach to answer the empirical question is less than ideal, but significantly cheaper and faster. Specifically, we use transfer learning with sensitivity analysis to “transfer” knowledge from the existing, 2020 experiment by Aggarwal et al., (2023) while accounting for measurable and unmeasurable changes in electoral context and voter demographics between 2020 and 2024. More formally, transfer learning uses the overlapping measurements about voters in 2020 and 2024 (e.g., gender, age group, party affiliation, voting history) to “adjust” for measurable differences between the two elections. We then use parameters from a sensitivity analysis to quantify any unmeasured differences between the elections (e.g., changes in electoral context). We also propose a new, data-driven procedure to calibrate/benchmark the magnitude of the parameters from the sensitivity analysis based on sample splitting and design sensitivity (Rosenbaum,, 2004, 2020).
With our framework, we estimate the effect of negative, digital ads against Trump on voter turnout for the 2024 U.S. presidential election. We focus on roughly $4.9$ million registered voters in Pennsylvania (PA) as our target population. PA is not only the largest swing state in terms of electoral votes, but also the “tipping point state” for the 2024 U.S. presidential election (e.g., FiveThirtyEight, (2024)). We present a county-by-county analysis of the ad effect and a subgroup analysis among $20$ politically important subgroups. In the county-by-county analysis, we find that if 2020 and 2024 elections are similar with respect to some of the voters’ demographics, the negative digital ad campaign against Trump would decrease voter turnout in Fulton county, a heavily Republican-leaning county. But, the ads would remain ineffective in all other counties of PA for the 2024 election. Moreover, if there are slight, unmeasured differences between 2020 and 2024, the ad effects change from being insignificant to significant in 60 counties. In the subgroup analysis, we find that the negative ads can decrease voter turnout among female voters in rural areas with low college education and increase turnout among non-female voters in urban areas with high college education.
1.2 Related work and contributions
Our work builds upon several works on generalizing or transporting treatment effects from a source population to a target population under a sensitivity analysis framework (Nguyen et al.,, 2017; Colnet et al.,, 2021; Dahabreh et al.,, 2023; Zeng et al.,, 2023; Duong et al.,, 2023; Ek and Zachariah,, 2023; Huang, 2024b, ). Specifically, we work under the exponential tilting sensitivity model (Robins et al.,, 2000), which has been used in works on generalizability and transportability (Dahabreh et al.,, 2022), and make the following new contributions.
  • (a)
    We allow the source data to have more covariates than the target data. Not only was this the case in our own data analysis, but this setting is common when the source data is derived from a randomized experiment where detailed information about the study units is collected. Zeng et al., (2023) also considered this setup for a similar reason, but focused on efficient and minimax estimation.
  • (b)
    We propose a simple regression estimator with nonparametric, percentile bootstrap to estimate the treatment effect in the target population; see Section 4.1. Notably, while qualitatively suggested by several works in this area, we formally show one theoretically correct approach to conduct bootstrap-based inference for transfer learning. We recommend this analysis pipeline for practitioners because of its simplicity, theoretically attractive properties (e.g., consistency, asymptotic normality), and the estimator based on the efficient influence function (EIF) is not doubly robust; see below.
  • (c)
    We also propose an estimator based on the EIF. This result extends the novel results in Zeng et al., (2023) to the setting where sensitivity parameters are present. While this estimator is more widely applicable than that in (b), it is more complex, requires estimating four nuisance functions and is not doubly robust; see Section 4.2
  • (d)
    For either procedure (b) or (c), we propose a simple calibration procedure to generate interpretable, “reference” magnitudes of the sensitivity parameter. Unlike existing methods for calibration based on omitting a measured covariate (e.g., Hsu and Small, (2013); Cinelli and Hazlett, (2020); Ek and Zachariah, (2023); Huang, 2024b ), our calibration procedure uses the same covariates for both calibration and sensitivity analysis. The calibration procedure is inspired by a clever idea underlying design sensitivity (Rosenbaum,, 2004) and sample splitting where we create a data-driven “favorable” situation (Rosenbaum,, 2020, Chapter 15) by splitting the source data in a particular way; see Section 5.
While the listed contributions are directly motivated from the statistical challenges in our data analysis, we believe they can be meaningful in other contexts, notably in generalizing the results of a randomized trial to a target population with mis-matching covariates and unmeasurable differences between the populations. More broadly, we hope our analysis pipeline centered on sensitivity analysis with transfer learning is useful to practitioners who want a simple, theoretically valid approach for generalization or transportation tasks.
2 Transfer learning between elections
2.1 Setup: Observed data
Suppose we collect $n_{s}$ independent and identically distributed (i.i.d.) samples from a source population. For each study unit $i\in{\cal I}_{s}=\{1,\ldots,n_{s}\}$ in the source data, we observe the following:
[…]
The variable ${\bf X}_{i}\in{\cal X}$ is the pre-treatment covariate (e.g., voter demographics), $A_{i}$ is the binary treatment indicator (e.g., assigned to ad campaign against Trump or not), $Y_{i}$ is the binary outcome (e.g., voted or not), and $S_{i}$ indicates whether unit $i$ is from the source sample (i.e., $S_{i}=1$) or not (i.e., $S_{i}=0$). In our data analysis, the source data is from Aggarwal et al., (2023). Independently, we also collect $n_{t}$ i.i.d. samples from the target population and for each study unit $i\in{\cal I}_{t}=\{n_{s}+1,\ldots,n_{s}+n_{t}=n\}$ from the target data, we observe the following [^1] For notational convenience, we overload the notation ${\bf O}_{i}$ to represent the observed data from unit $i$. If the data is from the source, ${\bf O}_{i}=({\bf X}_{i}$, $A_{i}$, $Y_{i},S_{i}=1)$ and if the data is from the target, ${\bf O}_{i}=({\bf V}_{i},S_{i}=0)$.:
[…]
The variable ${\bf V}_{i}\in{\cal V}\subseteq{\cal X}$ is a subset of the covariates in $\cal{X}$. In our data analysis, the target data consists of registered voters from PA’s voter registration database and ${\bf V}_{i}$ is voter $i$’s demographic information in the database (e.g., age group, gender, party affiliation, voting history). Because ${\bf V}_{i}$ is present in both the source and the target data, we refer to it as the shared covariate. Figure 2.1 summarizes our data setup.
Figure 2.1: A visualization of the data setup.
We make some remarks about the setup. First, if the covariates are discrete, some modeling assumptions about the outcome regression or the propensity score in Sections 4.1 and 4.2 are automatically satisfied. In our data analysis, all covariates were discrete. Second, we allow ${\cal V}\subseteq{\cal X}$ because, as far as we are aware of, there is no publicly available dataset of the 2024 voter population that measured the same attributes about voters as the source data from 2020. In general, we find that if the source population is from a randomized controlled trial, the covariates from it (i.e., ${\bf X}_{i}$) are richer than those from the target population (i.e., ${\bf V}_{i}$); see Zeng et al., (2023) who echoed a similar sentiment. Third, while we focus on binary outcomes $Y_{i}$ due to our data analysis, our framework generalizes to a continuous outcome; see Section A of the Appendix. Fourth, similar to other works on transfer learning, we assume that the units in the source and the target data are independent and sampled from an infinite population in order to derive asymptotic properties of our estimators below. But, this may lead to conservative inference in some settings (Jin and Rothenhäusler,, 2024) and Section 7 discusses this issue in the context of our data analysis.
2.2 Setup: Causal estimands and nuisance functions
We use the counterfactual framework to define causal effects. Let $Y^{(a)}_{i}$ be the counterfactual outcome of unit $i$ when the treatment is, possibly contrary to fact, set to $a\in\{0,1\}$. The causal estimand of interest, denoted as $\theta$, is the average treatment effect in the target population (TATE):
$\theta=\theta_{1}-\theta_{0},\text{ where }\theta_{a}=\mathbb{E}\left[Y^{(a)}_{i}\mid S_{i}=0\right]\text{ and }a\in\{0,1\}.$
In our data analysis, $\theta$ is the average effect of a digital ad campaign against Trump on voter turnout in 2024 among registered PA voters. We remark that for a binary outcome, other measures of treatment effects are possible, such as the risk ratio $\theta_{1}/\theta_{0}$ and the odds ratio $[\theta_{1}/(1-\theta_{1})][\theta_{0}/(1-\theta_{0})]^{-1}$. While we focus on mean differences (i.e. $\theta_{1}-\theta_{0}$) like Aggarwal et al., (2023), our results are derived for $\theta_{a}$ and thus, can be extended to cover the risk ratio and the odds ratio; see Ye et al., (2023) for an example.
We define the following functions, often referred to as nuisance functions.
  • The propensity score in the source population: $\pi({\bf x})=\mathbb{P}(A_{i}=1\mid{\bf X}_{i}={\bf x},S_{i}=1)$, ${\bf x}\in{\cal X}$.
  • The outcome models in the source population for each level of treatment $a\in\{0,1\}$:
    • With all covariates ${\bf X}_{i}$: $\mu_{a}({\bf x})=\mathbb{E}(Y_{i}\mid{\bf X}_{i}={\bf x},A_{i}=a,S_{i}=1)$, ${\bf x}\in{\cal X}$.
    • With the shared covariates ${\bf V}_{i}$: $\rho_{a}({\bf v})=\mathbb{E}(\mu_{a}({\bf X}_{i})\mid{\bf V}_{i}={\bf v},S_{i}=1)$, ${\bf v}\in{\cal V}$.
  • The ratio of probability densities of ${\bf V}_{i}$ between the two populations: $w({\bf v})=p_{{\bf V}_{i}\mid S_{i}=0}({\bf v}\mid S_{i}=0)/p_{{\bf V}_{i}\mid S_{i}=1}({\bf v}\mid S_{i}=1)$ where $p_{{\bf V}_{i}\mid S_{i}=s}(\cdot)$ is the conditional density of ${\bf V}_{i}$ given $S_{i}=s$, $s=0,1$.
We conclude by defining the following notations for order and convergences. For two real sequences of numbers $b_{n}$ and $d_{n}$, we denote $b_{n}=O(b_{n})$ if $|b_{n}|\leq C|d_{n}|$ for a constant $C$ and denote $b_{n}\asymp d_{n}$ if $b_{n}=O(d_{n})$ and $d_{n}=O(b_{n})$. We use $\to_{p}$ to denote convergence in probability and $\to_{d}$ to denote convergence in distribution. For a sequence of random variables $Z_{n}$ and a real sequence of numbers $b_{n}$, we denote $Z_{n}=o_{p}(b_{n})$ if $Z_{n}/b_{n}\to_{p}0$. For a measurable and integrable function $f(\cdot)$, we denote its $L_{2}$ norm by $\lVert f({\bf O}_{i})\rVert=\sqrt{\mathbb{E}\{f^{2}({\bf O}_{i})\}}$.
2.3 Causal identification
To identify the TATE, it’s common to make two sets of assumptions (Stuart et al.,, 2011; Tipton,, 2013; Nguyen et al.,, 2017; Dahabreh et al.,, 2023; Zeng et al.,, 2023; Huang, 2024a, ; Huang, 2024b, ). The first set of assumptions ensures the identification of the average treatment effect (ATE) in the source population with the source data.
Assumption 2.1 (Identification of the ATE in the Source Population)
  1. 1.
    (Stable Unit Treatment Value Assumption, SUTVA, Rubin, (1980)): $Y_{i}=Y_{i}^{(A_{i})}$ if $S_{i}=1$.
  2. 2.
    (Strong Ignorability, Rosenbaum and Rubin, (1983)): $Y^{(1)}_{i},Y^{(0)}_{i}\perp\!\!\!\perp A_{i}\mid{\bf X}_{i},S_{i}=1$ and $10<\pi({\bf x})<$ 1 for ${\bf x}\in{\cal X}$.
Assumption 2.1 is automatically satisfied if the source data is from a randomized controlled trial, such as our source data from Aggarwal et al., (2023). Also, to identify the TATE, SUTVA is not necessarily for the target population (i.e. $S_{i}=0$). This is because identification is based on transferring information about the potential outcomes, not the observed outcomes.
The second set of assumptions ensures that we can generalize or transfer the identified ATE from the source population to the target population.
Assumption 2.2 (Positivity of $S_{i}$)
$\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})>0$ for ${\bf v}\in{\cal V}$; $\mathbb{P}(S_{i}=0)>0$.
Assumption 2.3 (Transportability)
$Y^{(1)}_{i},Y^{(0)}_{i}\perp\!\!\!\perp S_{i}\mid{\bf V}_{i}$.
The first part of Assumption 2.2 will be violated if there are some values of the shared covariates ${\bf V}_{i}$ that are only observed in the target population, for instance if Aggarwal et al., (2023) focused only on young voters and the target population consists of voters from all ages. The second part of Assumption 2.2 excludes the case where the target sample size is much smaller than the source sample size. Because both parts depend solely on observable quantities, Assumption 2.2 can be checked with data; see Figure 6.1 for an example.
Assumption 2.3, referred to as transportability, states that conditional on the shared covariates ${\bf V}_{i}$, the distributions of the potential outcomes are identical between the source and the target populations. The assumption is violated if the distribution of the potential outcomes differ between the source and the target populations after adjusting for ${\bf V}_{i}$. For example, if ${\bf V}_{i}$ only contains political party, Assumption 2.3 will be violated if within each political party, voter turnout under treatment or control is different between the 2020 and 2024 elections. Unfortunately, unlike Assumption 2.2, Assumption 2.3 depends on counterfactual quantities and cannot be checked with data. Furthermore, unlike strong ignorability in Assumption 2.1, we are not aware of a feasible experimental design to guarantee Assumption 2.3 in electoral contexts.222 A study design that satisfies Assumption 2.3 is to randomize the selection of study units into the source data (Tipton,, 2013; Tipton and Peck,, 2017). In our data analysis, this design implies Aggarwal et al., (2023) randomized voters to be either in their 2020 study or to be a registered voter in PA for the 2024 election and we believe that this design is impractical. This is the main motivation for us to embed sensitivity analysis within transfer learning so that our framework does not rely on Assumption 2.3.
Under Assumptions 2.1-2.3, the TATE can be identified (Zeng et al.,, 2023):
$\theta$ $=\mathbb{E}\left[\mathbb{E}\left\{\mu_{1}({\bf X}_{i})-\mu_{0}({\bf X}_{i})\mid{\bf V}_{i},S_{i}=1\right\}\mid S_{i}=0\right].$ (2.1)
In words, $\theta$ is identified by first averaging the conditional average treatment (CATE) effect in the source population (i.e., $\mu_{1}({\bf X}_{i})-\mu_{0}({\bf X}_{i})$) over the shared covariates ${\bf V}_{i}$ (i.e., the inner expectation in equation (2.1)) and second, averaging this quantity among units in the target population (i.e., the outer expectation in equation (2.1)). For efficient and minimax estimation of $\theta$ under Assumptions 2.1-2.3, see Zeng et al., (2023).
3 Sensitivity analysis of transportability
As discussed above, suppose transportability (i.e., Assumption 2.3) no longer holds even after conditioning on $V_{i}$ and we measure the departure from it by the sensitivity parameter $\Gamma_{a}\in(0,\infty)=\mathbb{R}^{+}$ for $a\in\{0,1\}$. Specifically, the parameter $\Gamma_{a}$ is defined as the odds ratio of counterfactual outcomes between the target and source populations for a given ${\bf v}\in{\cal V}$:
$\Gamma_{a}$ $=\frac{{\rm ODD}_{a}({\bf v},0)}{{\rm ODD}_{a}({\bf v},1)},\ {\rm ODD}_{a}({\bf v},s)=\frac{\mathbb{P}(Y^{(a)}_{i}=1\mid{\bf V}_{i}={\bf v},S_{i}=s)}{\mathbb{P}(Y^{(a)}_{i}=0\mid{\bf V}_{i}={\bf v},S_{i}=s)},\ s\in\{0,1\},{\bf v}\in{\cal V}.$ (3.1)
When $\Gamma_{a}=1$, i.e., the conditional distributions of $Y^{(a)}_{i}$ given ${\bf V}_{i}$ are identical between the source and target populations (i.e., Assumption 2.3 holds). As $\Gamma_{a}$ moves away from $1$, the degree of violation of transportability increases. For example, in our data analysis, $\Gamma_{1}=1.05$ means that the counterfactual odd of voting in 2024 is $1.05$ times that in 2020 when a registered voter, possibly contrary to fact, gets negative ads against Trump. Similarly, $\Gamma_{1}=0.95$ means the counterfactual odd of voting in 2024 is $0.95$ times that in 2020 when a registered voter, possibly contrary to fact, gets negative ads about Trump.
Similar to other sensitivity analyses, the sensitivity parameter $\Gamma_{a}$ is not identifiable. Instead, investigators identify and estimate the TATE for a given $\Gamma_{a}$ and in doing so, study the sensitivity of the TATE when transportability is violated. Specifically, for a given $\Gamma_{a}\in\mathbb{R}^{+}$, the expected counterfactual outcome under treatment level $a\in\{0,1\}$ is identified as follows.
Lemma 3.1 (Identification of TATE Under Sensitivity Model)
Suppose Assumptions 2.1 and 2.2 hold. For a given $\Gamma_{a}\in\mathbb{R}^{+}$, the expected counterfactual outcome under treatment level $a\in\{0,1\}$ is
$\mathbb{E}[Y^{(a)}_{i}\mid S_{i}=0]$ $=\mathbb{E}\left[\dfrac{\Gamma_{a}\rho_{a}({\bf V}_{i})}{\Gamma_{a}\rho_{a}({\bf V}_{i})+1-\rho_{a}({\bf V}_{i})}\Bigg{|}S_{i}=0\right]=\theta_{a}(\Gamma_{a}).$ (3.2)
To highlight the inclusion of the sensitivity analysis, we denote the mean counterfactual outcome under treatment level $a$ by $\theta_{a}(\Gamma_{a})$ and the TATE by $\theta_{1}(\Gamma_{1})-\theta_{0}(\Gamma_{0})$. Despite the expanded notation, the interpretation of $\theta_{a}(\Gamma_{a})$ as an average of the counterfactual outcome $Y^{(a)}_{i}$ in the target population remains the same regardless of the value of $\Gamma_{a}$. For example, in our data analysis, if $\Gamma_{1}=1$, $\theta_{1}(1)$ is the proportion of registered PA voters who would vote in the 2024 election if all voters were assigned to anti-Trump digital ads and transportability held. If $\Gamma_{1}=1.1$, $\theta_{1}(1.1)$ is the proportion of registered PA voters who would vote in the 2024 election if all voters were assigned to anti-Trump digital ads and transportatbility was violated by $\Gamma_{1}=1.1$.
We conclude this section with a couple of remarks on the sensitivity model (3.1). First, this model was first proposed by Robins et al., (2000) as a non-parametric (just) identified model for describing selection bias in missing data. The model was later called an exponential tilting model (Rotnitzky et al.,, 2001; Birmingham et al.,, 2003) and an extrapolation-factorization model (Linero and Daniels,, 2018). The model was also used to conduct sensitivity analysis for unmeasured confounding in causal inference (Franks et al.,, 2020; Scharfstein et al.,, 2021) and for violation of the transportability assumption in generalizability (Dahabreh et al.,, 2022). In particular, when ${\cal V}={\cal X}$, Lemma 3.1 recovers the identification of the TATE in Dahabreh et al., (2022). Second, we choose this model for sensitivity analysis as it (a) posits no testable implications on the data, (b) makes statistical inference tractable (e.g., asymptotic normality), and (c) has a simple, odds ratio interpretation. Third, the sensitivity model can be extended in various ways. For example, it can be extended to handle a continuous, counterfactual outcome where the sensitivity model tilts the entire density of the counterfactual outcome; see Section A of the Appendix where we discuss identification, estimation, and interpretation of the TATE under a sensitivity model for a continuous, counterfactual outcome. Also, at the expense of more sensitivity parameters, model (3.1) can be extended to allow $\Gamma_{a}$ to depend on ${\bf V}_{i}$ and $Y^{(a)}_{i}$; see Franks et al., (2020) and Scharfstein et al., (2021) for examples. Fourth, model (3.1) can be reformulated under a selection model (see Section A of the Appendix) or under an $R^{2}$-based model (Franks et al.,, 2020).
4 Estimation and inference
4.1 Outcome regression and percentile bootstrap
The analysis pipeline in this section is appropriate when ${\cal X}$ is discrete or, more generally, when the outcome regression model $\rho_{a}({\bf v})$ can be consistently estimated at a parametric rate. This is the case in our data analysis where voter’s demographics are discrete variables. Even if ${\cal X}$ is not discrete, we suggest investigators begin with this analysis since it is not only simple, but also the alternative analysis based on the efficient influence function (EIF) is not doubly robust; see Section 4.2.
From the identification equation (3.2), a natural estimator of $\theta_{a}(\Gamma_{a})$ would be a plug-in estimator that takes a weighted average of an estimator of the outcome regression function $\rho_{a}({\bf v})$, denoted as $\widehat{\rho}_{a}({\bf v})$, among the target sample. We call this estimator the outcome regression (OR) estimator:
$\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})=\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})}{\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})}.$ (4.1)
Also, from the definition of $\rho_{a}({\bf v})$ n Section 2.2, a simple estimator of $\widehat{\rho}_{a}({\bf v})$ is to regress $\widehat{\mu}_{a}({\bf x})$ on ${\bf v}$ using ordinary least squares (OLS) and $\widehat{\mu}_{a}({\bf x})$ is an estimate of $\mu_{a}({\bf x})$. If ${\cal X}$ is discrete, the OLS estimators of $\widehat{\mu}_{a}({\bf x})$ and $\widehat{\rho}_{a}({\bf v})$ can be expressed as
$\widehat{\mu}_{a}({\bf x})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}Y_{i}\mathds{1}(A_{i}=a,{\bf X}_{i}={\bf x})}{\sum_{i\in{\cal I}_{s}}\mathds{1}(A_{i}=a,{\bf X}_{i}={\bf x})},\ \widehat{\rho}_{a}({\bf v})=\dfrac{\sum_{i\in{\cal I}_{s}}\widehat{\mu}_{a}({\bf X}_{i})\mathds{1}({\bf V}_{i}={\bf v})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})},\ {\bf x}\in{\cal X},{\bf v}\in{\cal V},$ (4.2)
where $\mathds{1}(\cdot)$ is the indicator function. In the discrete case, the estimators in equation (4.2) are consistent. For a general discussion on estimating $\rho_{a}$, see Section 4.3.
For inference, we recommend a nonparametric, percentile bootstrap (Efron,, 1979) where the source and the target data are resampled separately and we take the $\alpha/2$ and $21-\alpha/$ 2 quantiles of the bootstrapped estimates of $\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})$, denoted $\widehat{L}_{a}(\Gamma_{a};1-\alpha)$ and $\widehat{U}_{a}(\Gamma_{a};1-\alpha)$ respectively. These quantiles are used to construct a $(1-\alpha)$ confidence interval (CI), denoted as $\widehat{\text{CI}}_{\textnormal{OR,a}}(\Gamma_{a};1-\alpha)=\left[\widehat{L}_{a}(\Gamma_{a};1-\alpha),\widehat{U}_{a}(\Gamma_{a};1-\alpha)\right]$.
Suppose $\rho_{a}({\bf v})$ is indexed by a finite-dimensional parameter $\boldsymbol{\eta}_{a}$. Theorem 4.1 shows that under regularity conditions, the plug-in, OR estimator $\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})$ in equation (4.1) is consistent and the nonparametric, percentile bootstrap leads to a valid CI.
Theorem 4.1 (Theoretical properties of the OR estimator and bootstrapped CI)
Suppose Assumptions 2.1 and 2.2 hold and $\theta_{a}(\Gamma_{a})\in\Theta$ where $\Theta$ is open and compact. Also suppose $\rho({\bf v};\boldsymbol{\eta}_{a})$ is twice differentiable with respect to $\boldsymbol{\eta}_{a}$. If $\widehat{\boldsymbol{\eta}}_{a}$ is an asymptotically linear estimate of $\boldsymbol{\eta}_{a}$ and $n_{s}\asymp n_{t}$, then $\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})\to_{p}\theta_{a}(\Gamma_{a})$. Furthermore, if regularity conditions (B1)-(B4) in Section B of the Appendix hold, the bootstrap interval $\widehat{\textnormal{CI}}_{\textnormal{OR,a}}(\Gamma_{a};1-\alpha)$ for $\alpha\in(0,0.5)$ satisfies $\mathbb{P}(\theta_{a}(\Gamma_{a})\in\widehat{\textnormal{CI}}_{\textnormal{OR,a}}\left(\Gamma_{a};1-\alpha)\right)\to 1-\alpha$.
4.2 Efficient influence function
The analysis pipeline in this section is based on the efficient influence function (EIF) and is more broadly applicable than that in Section 4.1, especially when ${\cal X}$ is not discrete and the propensity score is unknown. However, the EIF-based estimator is more complex and requires estimating multiple nuisance functions.
To motivate the estimator, we first derive the EIF of $\theta_{a}(\Gamma_{a})$ in Theorem 4.2.
Theorem 4.2 (Efficient Influence Function of $\theta_{a}(\Gamma_{a})$)
Under Assumptions 2.1 and 2.2, the EIF of $\theta_{a}(\Gamma_{a})$ is
$\textnormal{EIF}({\bf O}_{i},\theta_{a}(\Gamma_{a}))=$ $\dfrac{S_{i}w({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\dfrac{\Gamma_{a}}{[\Gamma_{a}\rho_{a}({\bf V}_{i})+1-\rho_{a}({\bf V}_{i})]^{2}}\left[\left\{\dfrac{A_{i}}{\pi({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\pi({\bf X}_{i})}\right\}\{Y_{i}-\mu_{a}({\bf X}_{i})\}\right.$
$+\mu_{a}({\bf X}_{i})-\rho_{a}({\bf V}_{i})\bigg{]}+\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\left[\dfrac{\Gamma_{a}\rho_{a}({\bf V}_{i})}{\Gamma_{a}\rho_{a}({\bf V})+1-\rho_{a}({\bf V}_{i})}-\theta_{a}(\Gamma_{a})\right],$
Also, if the propensity score $\pi({\bf X}_{i})$ is known, the EIF of $\theta_{a}(\Gamma_{a})$ remains unchanged.
We remark that when transportability holds, i.e., $\Gamma_{a}=1$, Theorem 4.2 reduces to the EIF in Zeng et al., (2023).
Following the modern trend in causal inference, we use cross-fitting and the EIF (e.g., Chernozhukov et al., (2017); Kennedy, (2022)) to estimate $\theta_{a}(\Gamma_{a})$. Specifically, we randomly partition the source and target sample indices ${\cal I}_{s}$ and ${\cal I}_{t}$ into $K$ disjoint sets, ${\cal I}_{s}^{(k)}$ and ${\cal I}_{t}^{(k)}$, respectively, for $k=1,2,\cdots,K$, and let ${\cal I}^{(k)}={\cal I}_{s}^{(k)}\cup{\cal I}_{t}^{(k)}$. For each $k$, the nuisance functions are estimated with data in ${\cal I}\setminus{\cal I}^{(k)}$ and they are denoted as $\widehat{\pi}^{(k)}({\bf x})$, $\widehat{\mu}_{a}^{(k)}({\bf x})$, $\widehat{w}^{(k)}({\bf v})$ and $\widehat{\rho}_{a}^{(k)}({\bf v})$. We then plug them into the “uncentered” EIF and evaluate it with the data in ${\cal I}^{(k)}$:
$\widehat{\theta}_{\textnormal{EIF,a}}^{(k)}(\Gamma_{a})=$ $\dfrac{1}{|{\cal I}_{s}^{(k)}|}\sum_{i\in{\cal I}_{s}^{(k)}}\dfrac{\Gamma_{a}\widehat{w}^{(k)}({\bf V}_{i})}{[\Gamma_{a}\widehat{\rho}_{a}^{(k)}({\bf V}_{i})+1-\widehat{\rho}_{a}^{(k)}({\bf V}_{i})]^{2}}\left[\left\{\dfrac{A_{i}}{\widehat{\pi}^{(k)}({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\widehat{\pi}^{(k)}({\bf X}_{i})}\right\}\{Y_{i}-\widehat{\mu}_{a}^{(k)}({\bf X}_{i})\}\right.$
$\left.\quad{}\quad{}\quad{}\quad{}+\widehat{\mu}_{a}^{(k)}({\bf X}_{i})-\widehat{\rho}_{a}^{(k)}({\bf V}_{i})\right]+\dfrac{1}{|{\cal I}_{t}^{(k)}|}\sum_{i\in{\cal I}_{t}^{(k)}}\dfrac{\Gamma_{a}\widehat{\rho}_{a}^{(k)}({\bf V}_{i})}{\Gamma_{a}\widehat{\rho}_{a}^{(k)}({\bf V}_{i})+1-\widehat{\rho}_{a}^{(k)}({\bf V}_{i})}.$
Finally, we take an average of $\widehat{\theta}_{\textnormal{EIF,a}}^{(k)}(\Gamma_{a})$ to arrive at the EIF-based, cross-fitting estimator of $\theta_{a}(\Gamma_{a})$, which we denote as $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})=K^{-1}\sum_{k=1}^{K}\widehat{\theta}_{\textnormal{EIF,a}}^{(k)}(\Gamma_{a})$. A step-by-step algorithm can be found from Section C of the Appendix. Theorem 4.3 shows that under conditions, $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})$ is consistent, asymptotically normal, and semiparametrically efficient.
Theorem 4.3 (Theoretical properties of the EIF-based estimator)
Suppose Assumptions 2.1 and 2.2 hold and there exist $c,C>0$ such that $c<\widehat{\pi}^{(k)}({\bf x})$, $\widehat{w}^{(k)}({\bf v})<C$ and $\widehat{\rho}_{a}^{(k)}({\bf v})\in[0,1]$ for ${\bf v}\in{\cal V}$ and ${\bf x}\in{\cal X}$. Then, the following holds:
(i) [Conditional Double Robustness]. Suppose $\widehat{\rho}^{(k)}_{a}$ is a consistent estimator of $\rho^{(k)}_{a}$ (i.e., $\lVert\widehat{\rho}^{(k)}_{a}({\bf V}_{i})-\rho^{(k)}_{a}({\bf V}_{i})\rVert=o_{p}(1)$). Then, $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})\to_{p}\theta_{a}(\Gamma_{a})$ if
$\lVert\widehat{\pi}^{(k)}({\bf X}_{i})-\pi^{(k)}({\bf X}_{i})\rVert\cdot\lVert\widehat{\mu}_{a}^{(k)}({\bf X}_{i})-\mu_{a}^{(k)}({\bf X}_{i})\rVert$ $=o_{p}(1),$ (4.3)
(ii) [Asymptotic normality and Semiparametric Efficiency] Suppose $\widehat{\rho}^{(k)}_{a}$, $\widehat{\mu}_{a}^{(k)}$, $\widehat{w}^{(k)}$, and $\widehat{\pi}^{(k)}$ are consistent estimators with the following rates:
$\lVert\widehat{\pi}^{(k)}({\bf X}_{i})-\pi^{(k)}({\bf X}_{i})\rVert\cdot\lVert\widehat{\mu}_{a}^{(k)}({\bf X}_{i})-\mu_{a}^{(k)}({\bf X}_{i})\rVert=o_{p}(n^{-1/2}),$ (4.4a)
$\lVert\widehat{w}^{(k)}({\bf V}_{i})-w^{(k)}({\bf V}_{i})\rVert\cdot\lVert\widehat{\rho}^{(k)}_{a}({\bf V}_{i})-\rho^{(k)}_{a}({\bf V}_{i})\rVert=o_{p}(n^{-1/2}),\text{ and}$ (4.4b)
$\lVert\widehat{\rho}^{(k)}_{a}({\bf V}_{i})-\rho^{(k)}_{a}({\bf V}_{i})\rVert^{2}=o_{p}(n^{-1/2}).$ (4.4c)
Then, $\sqrt{n}\left\{\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})-\theta_{a}(\Gamma_{a})\right\}\to_{d}N\left(0,\sigma^{2}_{\textnormal{EIF,a}}(\Gamma_{a})\right)$ where $\sigma^{2}_{\textnormal{EIF,a}}(\Gamma_{a})=\mathbb{E}[\{\textnormal{EIF}({\bf O}_{i},\theta_{a}(\Gamma_{a}))\}^{2}]$.
(iii) [Consistent Estimator of Standard Error] Suppose the same assumptions in (ii) hold. Then, $\widehat{\sigma}^{2}_{\textnormal{EIF,a}}(\Gamma_{a})\to_{p}\sigma^{2}_{\textnormal{EIF,a}}(\Gamma_{a})$, where $\widehat{\sigma}_{\textnormal{EIF,a}}^{2}(\Gamma_{a})=K^{-1}\sum_{k=1}^{K}\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\left\{\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))\right\}^{2}$ and $\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))$ is the empirical counterpart of $\textnormal{EIF}^{(k)}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))$ with plug-in estimates of the nuisance parameters $\widehat{\pi}^{(k)}$, $\widehat{\rho}_{a}^{(k)}$, $\widehat{w}^{(k)}$, and $\widehat{\mu}_{a}^{(k)}$.
Part (i) of Theorem 4.3 states that $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})$ is conditionally doubly robust in that if $\widehat{\rho}_{a}^{(k)}$ is consistent, $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})$ is consistent when either $\widehat{\pi}^{(k)}({\bf x})$ or $\widehat{\mu}_{a}^{(k)}$, but not necessarily both, is consistent. Part (ii) states that if all the nuisance functions are estimated consistently at the rates in equations (4.4a)-(4.4c), $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})$ is asymptotically normal and semiparametrically efficient. We remark that when transportability holds, our result recovers Theorem 5 of Zeng et al., (2023), which does not require equation (4.4c); see below for more discussions. Finally, Theorem 4.3 implies that an asymptotically valid, $1-\alpha$ CI of $\theta_{a}(\Gamma_{a})$ is $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})\pm z_{1-\alpha/2}\sqrt{\widehat{\sigma}_{\textnormal{EIF,a}}^{2}(\Gamma_{a})}$ where $z_{1-\alpha/2}$ is the $21-\alpha/$ 2 quantile of the standard normal distribution.
Equation (4.4c) requires that we not only consistently estimate the outcome regression $\rho_{a}({\bf v})$, but also estimate it at a sufficiently fast rate. If the true $\rho_{a}({\bf v})$ is a parametric function as in Theorem 4.1, equation (4.4c) is satisfied with a parametric estimator. Otherwise, we cannot estimate $\rho_{a}({\bf v})$ at a slow, nonparametric rate in hopes that another estimator of the nuisance function can “compensate” for the slow rate; this is referred to as the mixed bias property or rate double robustness (e.g., Rotnitzky et al., (2020), Kennedy, (2022)). In contrast, one approach to satisfy equation (4.4a) is to obtain data from an RCT where the propensity score is known a priori and estimate the outcome regression using a supervised machine learning method, which may converge slowly. More broadly, equation (4.4c) can be viewed as the cost of violating transportability. Intuitively, we incur this cost because the sensitivity model (3.1) is based on shifting the outcome distribution and if $\rho_{a}$ is poorly estimated, the sensitivity model is also incorrectly specified, which ultimately leads to a poor estimate of the TATE; see Section C of the Appendix for more details.
4.3 Estimation of nuisance functions
This section briefly discusses estimation of the nuisance parameters, specifically $\rho_{a}$ and $\omega$. For the other, “classical” nuisance functions (i.e., propensity score $\pi$ and the outcome regression function $\mu_{a}$), we echo the modern recommendation of using the investigator’s favorite classification and regression models. Note that if the source data is from an RCT, investigators should use the design of the RCT to estimate $\pi$.
The regression function $\rho_{a}$ can be estimated in a couple of different ways and we highlight each approach through the equalities below:
$\rho_{a}({\bf v})$ $=\mathbb{E}\{\mu_{a}({\bf X}_{i})\mid{\bf V}_{i}={\bf v},S_{i}=1\}$ (4.5)
$=\mathbb{E}\left[\left\{\dfrac{A_{i}\mathds{1}(A_{i}=a)}{\pi({\bf X}_{i})}+\dfrac{(1-A_{i})\mathds{1}(A_{i}=1-a)}{1-\pi({\bf X}_{i})}\right\}Y_{i}\mid{\bf V}_{i}={\bf v},S_{i}=1\right]$ (4.6)
$=\mathbb{E}\left[\left\{\dfrac{A_{i}\mathds{1}(A_{i}=a)}{\pi({\bf X}_{i})}+\dfrac{(1-A_{i})\mathds{1}(A_{i}=1-a)}{1-\pi({\bf X}_{i})}\right\}\{Y_{i}-\mu_{a}({\bf X}_{i})\}+\mu_{a}({\bf X}_{i})\mid{\bf V}_{i}={\bf v},S_{i}=1\right].$ (4.7)
The first equality (4.5) suggests estimating $\rho_{a}$ by regressing the predicted outcome $\widehat{\mu}_{a}({\bf X}_{i})$ on ${\bf V}_{i}$. The second equality (4.6) suggests regressing an inverse-probability-weighted (IPW) outcome $\left[A_{i}\mathds{1}(A_{i}=a)/\widehat{\pi}({\bf X}_{i})+(1-A_{i})\mathds{1}(A_{i}=1-a)/\{1-\widehat{\pi}({\bf X}_{i})\}\right]Y_{i}$ on ${\bf V}_{i}$. The third equality (4.7) suggests regressing an augmented IPW outcome,
$\left[A_{i}\mathds{1}(A_{i}=a)/\widehat{\pi}({\bf X}_{i})+(1-A_{i})\mathds{1}(A_{i}=1-a)/\{1-\widehat{\pi}({\bf X}_{i})\}\right]\{Y_{i}-\widehat{\mu}_{a}({\bf X}_{i})\}+\widehat{\mu}_{a}({\bf X}_{i})$
on ${\bf V}_{i}$. Under the first and the third approaches, the rate of convergence of $\widehat{\rho}_{a}$ is dependent of the rate of $\widehat{\mu}_{a}$ (Kennedy,, 2023). Under the second approach, the rate of convergence of $\widehat{\rho}_{a}$ is independent of the rate of convergence of $\widehat{\mu}_{a}$. We remark that when all covariates are discrete and $\pi$ and $\mu_{a}$ are estimated by taking means within subgroups defined by the covariates, the three approaches are equivalent.
For estimating the density ratio $w({\bf v})$, we recommend entropy balancing methods (Hainmueller,, 2012; Josey et al.,, 2022; Chen et al.,, 2023), which obtains $\widehat{w}({\bf V}_{i})$ as solutions to the following constrained optimization problem,
$\begin{split}&\underset{w_{i}}{\rm argmin}\sum_{i\in{\cal I}_{s}}w_{i}{\textnormal{log}}(w_{i}),\quad{}\textnormal{s.t. }\dfrac{1}{n_{s}}\sum_{i\in{\cal I}_{s}}w_{i}{\bf V}_{i}=\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}{\bf V}_{i}.\end{split}$ (4.8)
If the true $\mathbb{P}(S_{i}=1\mid{\bf V}_{i})$ is a logistic regression model and the parameters of the model are identified, the probability limit of the weights in (4.8) is equal to $\mathbb{P}(S_{i}=0\mid{\bf V}_{i})\mathbb{P}(S_{i}=1)/\{\mathbb{P}(S_{i}=1\mid{\bf V}_{i})\mathbb{P}(S_{i}=0)\}$. Otherwise, the weights in (4.8) generally have favorable, finite-sample properties (e.g., Chen et al., (2023)). For more discussions on estimating $w({\bf v})$, see Section C of the Appendix.
5 Calibrating sensitivity parameters
5.1 Definition of a sensitive effect and motivation for calibration
Section 4 provides two procedures to estimate the TATE for a given value of $(\Gamma_{0},\Gamma_{1})$ and allows investigators to study the change of the TATE as it moves away from $(\Gamma_{0},\Gamma_{1})=(1,1)$, i.e., the setting where transportability holds. Traditionally, investigators consider several $(\Gamma_{0},\Gamma_{1})$s that are not equal to $(1,1)$ and assess whether the statistical conclusion about the TATE changes between $(\Gamma_{0},\Gamma_{1})=(1,1)$ and other $(\Gamma_{0},\Gamma_{1})$s. A bit more formally, let ${\cal C}\subset\mathbb{R}^{+}\times\mathbb{R}^{+}\neq\{(1,1)\}$ denote the set of $(\Gamma_{0},\Gamma_{1})$s that the investigator is considering for the sensitivity analysis. Following the literature, we say the TATE is sensitive to transportability if the decision to reject the null hypothesis of no effect at the significance level $\alpha$ changed between $(\Gamma_{0},\Gamma_{1})=(1,1)$ and another value of $(\Gamma_{0},\Gamma_{1})\in{\cal C}$.
Definition 5.1 (Sensitivity to Transportability)
Consider a significance level $\alpha$ and the set ${\cal C}\subset\mathbb{R}^{+}\times\mathbb{R}^{+}\neq\{(1,1)\}$. For a given $(\Gamma_{0},\Gamma_{1})$, let $\widehat{\textnormal{CI}}(\Gamma_{0},\Gamma_{1};1-\alpha)$ denote a $1-\alpha$ CI of the TATE from Section 4. The TATE is sensitive to transportability in ${\cal C}$ if there exists $(\Gamma_{0},\Gamma_{1})\in{\cal C}$ such that either of the following holds:
$\begin{split}\text{(i) from significant to insignificant:}\quad{}&0\not\in\widehat{\textnormal{CI}}(1,1;1-\alpha)\text{ and }0\in\widehat{\textnormal{CI}}(\Gamma_{0},\Gamma_{1};1-\alpha);\\ \text{(ii) from insignificant to significant:}\quad{}&0\in\widehat{\textnormal{CI}}(1,1;1-\alpha)\text{ and }0\not\in\widehat{\textnormal{CI}}(\Gamma_{0},\Gamma_{1};1-\alpha).\end{split}$
If neither (i) nor (ii) holds, the TATE is insensitive to transportability in ${\cal C}$.
Some investigators have a well-defined ${\cal C}$ based on their belief about the unmeasured difference between the source and the target populations in the odds ratio scale. But in general, specifying a reasonable, “reference” set of sensitivity parameters has been a long-standing question in the literature on sensitivity analysis and this task is often referred to as calibration or benchmarking (Cinelli and Hazlett,, 2020; Huang, 2024b, ). One popular approach to generate the reference values is to omit an observed covariate (e.g., Hsu and Small, (2013); Cinelli and Hazlett, (2020); Ek and Zachariah, (2023); Huang, 2024b ) and conduct the sensitivity analysis with the values of the sensitivity parameters that are comparable to the effects of the omitted covariate on the outcome or the treatment. But, as discussed in Section 6.2 of Cinelli and Hazlett, (2020), this can lead to a misleading understanding of the magnitude of unmeasured confounding, especially when the omitted variable is correlated with other confounders.
In this section, we present a data-driven calibration procedure that generates a reference, calibrated set of the sensitivity parameters using an idea from design sensitivity (Rosenbaum,, 2004, 2020). Briefly, design sensitivity is used to benchmark designs of observational studies in terms of robustness against unmeasured confounding by measuring the limiting power to accept a particular type of alternative hypothesis, referred to as a “favorable situation” (Chapter 15 of Rosenbaum, (2020)). While Rosenbaum created the favorable situation from parametric models, we create it using the source data. Also, our calibrated set is a finite-sample, two-dimensional analog of Rosenbaum’s design sensitivity parameter in that the sensitivity parameters in the calibrated set lead to “accepting” the favorable situation created from data. Importantly, our calibration procedure avoids the issue from the omitted variables approach above by using the same covariates for both sensitivity analysis and calibration.
We state the calibration procedure in Section 5.2 and describe the rationale in Section 5.3.
5.2 Calibration procedure
The calibration procedure is divided into three steps. The first step partitions the source data from Aggarwal et al., (2023) into the rust belt states (PA, MI, WI), denoted as ${\cal I}_{s_{1}}$, and the sun belt states (AZ, NC), denoted as ${\cal I}_{s_{2}}$. The second step temporarily treats the voters in the sun belt states as the “proxy” target population and constructs two $1-\alpha$ CIs of the TATE for it:
  • (Our Transfer Learning Approach): We treat the rust belt states as the “proxy” source population and use the methods in Section 4 to infer the TATE in the sun belt states (i.e., the proxy target population). We denote the resulting confidence interval as ${\widehat{\text{CI}}}_{s_{1}\to s_{2}}(\Gamma_{0},\Gamma_{1};1-\alpha)$.
  • (The Standard Approach): Using the data from the sun belt states only (i.e., the proxy target population), we compute a valid $1-\alpha$ CI of the TATE, say the Wald confidence interval based on the difference-in-means estimator, and denote it as $\widehat{{\text{CI}}}_{s_{2}}(1-\alpha)$.
The third step keeps the values of $(\Gamma_{0},\Gamma_{1})$ where the CIs from both approaches overlap, or formally, $\mathcal{C}_{1}=\{(\Gamma_{0},\Gamma_{1})\mid\widehat{{\text{CI}}}_{s_{1}\to s_{2}}(\Gamma_{0},\Gamma_{1};1-\alpha)\cap\widehat{{\text{CI}}}_{s_{2}}(1-\alpha)\neq\emptyset\}$. We repeat the three steps above, but with the roles of the proxy target and proxy source populations reversed, yielding another set of sensitivity parameters ${\cal C}_{2}$. The intersection of the two sets, ${\cal C}={\cal C}_{1}\cap{\cal C}_{2}$, is the data-driven, calibration set of sensitivity parameters. Further computational details and a step-by-step algorithm are provided in Section D of the Appendix.
5.3 The rationale behind the calibration procedure
The first step partitions the source data into two subsets ${\cal I}_{s_{1}}$ and ${\cal I}_{s_{2}}$ such that there are scientifically meaningful, unobserved differences between them. For example, there are meaningful differences in socioeconomic status, labor markets, and region-specific politics between the sun belt states and the rust belt states and these differences are not measured by ${\bf V}_{i}$. We remark that investigators can choose other partitions that are interpretable; see below for more discussion.
After partitioning the data, the next two steps find $(\Gamma_{0},\Gamma_{1})$s that quantify the unmeasured differences between ${\cal I}_{s_{1}}$ and ${\cal I}_{s_{2}}$. This is accomplished by finding a set of $(\Gamma_{0},\Gamma_{1})$s such that the transported ATE (i.e., “Our Transfer Learning Approach” above) is close to the true ATE of the proxy target population, up to sampling error. Note that the true TATE in the proxy target population can be inferred with only Assumption 2.1, specifically using the $1-\alpha$ CI from the standard approach above. Then, the resulting set ${\cal C}$ represents the magnitude of the unmeasured differences between the two subsets ${\cal I}_{s_{1}}$ and ${\cal I}_{s_{2}}$ as the values in ${\cal C}$ correctly transported from the proxy source data to match the true ATE of the proxy target population.
It’s important to recognize that the set ${\cal C}$ obtained from the calibration procedure is not the true unmeasurable differences between the original source population and the original target population, for instance the unmeasured differences between the 2020 and 2024 elections. Similarly, using the calibrated set ${\cal C}$ in the sensitivity analysis does not imply that the true unmeasured differences between the original source and the target populations can be estimated from the unmeasured differences between the two subsets of the original source data. As mentioned in Section 3, the $(\Gamma_{0},\Gamma_{1})$ that parametrizes the unmeasured differences between the original source and the target populations cannot be identified or estimated.
Instead, akin to the usual approach of conducting sensitivity analysis based on a set ${\cal C}$ informed by the investigator’s prior belief, the calibrated set $\mathcal{C}$ is a data-driven approach to generate another interpretable set of sensitivity parameters. For example, if the TATE of the 2024 election is sensitive with respect to the set $\mathcal{C}$ generated from the investigator’s prior belief, it suggests that the unmeasured differences that are as large as those hypothesized by the investigator can overturn the conclusion about the TATE in the 2024 election. Similarly, if the TATE of the 2024 election is sensitive with respect to the calibrated set $\mathcal{C}$, it suggests that the unmeasured differences that are as large as those between the sun belt states and the rust best states in the 2020 election can overturn the conclusion about the TATE in the 2024 election. In short, our calibration procedure is another approach to understand and interpret the sensitivity parameters that is based on the observed data.
We also briefly mention a subtle point about the sample size and sampling uncertainty in the calibration procedure. Technically speaking, the partitioning step has a different sample size than that for the original analysis, which leads to different magnitudes of sampling uncertainty. Section D of the Appendix discusses how we re-scale the standard errors and conduct downsampling in the calibration procedure so that the sampling uncertainty is comparable between the calibration procedure and the original analysis. A broader discussion about sampling uncertainty in transfer learning is in Section 7.
Finally, as mentioned in the beginning of this section, investigators can choose other partitions of the source data in the first step. But, some partitions are more useful than others. For example, a random partition of the source data such that there are no unmeasurable differences between the two subsets is not meaningful. Nevertheless, between two non-random partitions, some investigators may find one partition to be more interpretable than the other. In fact, the investigators’ unrestricted ability to choose a partition is a useful feature of our calibration procedure compared to the omitted variable approach where the investigators are restricted to the list of observed confounders or usually focus on the “strongest” confounder. In general, compared to the calibration procedure based on omitting a covariate, we believe creating dissimilar partitions of the source is a promising way to study unobservable differences between the source and the target populations.
6 Ad effects in Pennsylvania for the 2024 election
6.1 Setup
We apply our approach to study the main empirical question from the paper, i.e., what is the effect of running a negative, digital ad campaign against Trump among registered voters in PA for the 2024 U.S. presidential election? The target data is from the PA’s voter registration database as of April 15, 2024, which initially contained $8,716,343$ registered voters. To harmonize with the source data by Aggarwal et al., (2023), we took a subset of voters in the PA database who are between 18 and 55 years old. We also recoded age, political party registration, and voting history in the PA database to match the definitions in the source data. In the end, we had $n_{t}=4,880,729$ registered voters in the target data and the shared covariates ${\bf V}_{i}$ included gender, age groups, party, and a subset of the voting history. The source covariates ${\bf X}_{i}$ included ${\bf V}_{i}$, race, and a richer set of voting history from Aggarwal et al., (2023) and there were $n_{s}=1,999,282$ registered voters from the source data. Figure 6.1 visualizes all of the covariates. For more details on the data description and data cleaning, see Section E of the Appendix.
For all 67 counties of PA, we estimate the ad effect in Section 6.2. We also conduct a subgroup analysis by gender, urbanicity, and education in Section 6.3. Due to page constraints and since all covariates are discrete, we present the results from the OR estimator and discuss the results from the EIF estimator in Section E of the Appendix; except for few discrepancies noted in Section 6.4, the two estimators reach the same conclusion. The regression function $\rho_{a}({\bf v})$ is estimated by regressing $\widehat{\mu}_{a}({\bf x})$ on ${\bf v}$. Following Aggarwal et al., (2023), $\mu_{a}({\bf x})$ is estimated by weighted least squares where the weights are the inverse propensity scores. The density ratio $w^{(k)}({\bf v})$ is estimated by entropy balancing in (4.8). As discussed above, we obtain calibrated sensitivity parameters by partitioning the source data into the rust belt states (i.e., PA, WI, MI) and the sun belt states (i.e., AZ, NC). Following Aggarwal et al., (2023), $\widehat{\rm CI}_{s_{1}}$ and $\widehat{\rm CI}_{s_{2}}$ in the calibration procedure are based on weighted least squares that regresses the outcome on the treatment and pre-treatment covariates and the weights are the inverse of the propensity scores.
Throughout the analysis, the significance level is $\alpha=0.05$. Also, as a reminder, a positive effect means that running negative ads against Trump increased voter turnout whereas a negative effect means that running negative ads decreased voter turnout.
Figure 6.1: Covariate distributions from the 2020 RCT by Aggarwal et al., (2023) (i.e., the source data), 2024 PA voters (i.e., the target data), and selected counties in PA during 2024. “NA” means the corresponding variable is missing.
6.2 Ad effect by counties
6.2.1 Ad effect under transportability
When $\Gamma_{0}=\Gamma_{1}=1$ (i.e., transportability holds), the ad effect is negative (i.e., decreased voter turnout when assigned to anti-Trump ads) and barely significant in Fulton county (95% CI: $[-1.64\%,-0.04\%]$, $p=0.04$). In all other 66 counties, the ad effects are insignificant; see panel A of Figure 6.2 for a visual illustration and Section E of the Appendix for the exact numbers. In other words, if the difference in voter turnout between PA voters in 2024 and the voters in 2020 can be completely adjusted with ${\bf V}_{i}$, then the negative ads will be ineffective in almost all counties for the 2024 election, except for Fulton county.
6.2.2 Ad effect with pre-specified $(\Gamma_{0},\Gamma_{1})$s
Next, we study the ad effects for different values of $(\Gamma_{0},\Gamma_{1})$. For brevity, we present two values of $(\Gamma_{0},\Gamma_{1})$ in panel B of Figure 6.2, and defer other values to Section E of the Appendix. We remark that this section mirrors a “traditional” sensitivity analysis discussed in Section 5.1 where the investigator pre-specifies $(\Gamma_{0},\Gamma_{1})$s based on existing, prior beliefs about the unmeasured differences between the 2020 and the 2024 elections.
Suppose $\Gamma_{0}=1.01$ and $\Gamma_{1}=0.99$, i.e., in the control arm, the counterfactual odd of voting in 2024 is $1.01$ times the counterfactual odd in 2020 and in the treated arm, the counterfactual odd of voting in 2024 is $0.99$ times the counterfactual odd in 2020. The ad effect is significant and negative in 51 counties and insignificant in 16 counties. The p-value is the smallest in Fulton ($p=0.021$), followed by Bedford ($p=0.027$) and Juniata ($p=0.042$).
Conversely, suppose $\Gamma_{0}=0.99$ and $\Gamma_{1}=1.01$, i.e., in the control arm, the counterfactual odd of voting in 2024 is $0.99$ times that in 2020 and in the treated arm, the counterfactual odd of voting in 2024 is $1.01$ times that in 2020. The ad effect is significant and positive in Philadelphia county ($p=0.044)$ and insignificant in other counties.
6.2.3 Ad effect with the calibrated set
We use the sensitivity parameters from the calibrated set ${\cal C}$ in Section 5.2 to conduct the sensitivity analysis. With a slight abuse of notation, we use ${\cal C}$ to denote the calibrated set for every county; see the discussion on sampling uncertainty in Section 5.3 and panel C of Figure 6.2 for examples of ${\cal C}$. An illustration of the calibrated results is shown in panel A of Figure 6.3.
Following Definition 5.1, 61 counties are sensitive to transportability within the calibration set. Philadelphia county is sensitive in that its result changed from an insignificant ad effect under transportability to a significant and positive effect when transportability is violated by the amount in the calibrated set ${\cal C}$; we refer to this type of sensitivity as sensitive for a positive effect. Bedford county is sensitive in that its result changed from an insignificant effect under transportability to a significant and negative effect when transportability is violated by the amount in the calibrated set ${\cal C}$; we refer to this type of sensitivity as sensitive for a negative effect. Fulton county is sensitive to transportability in that its result changed from a significant and negative effect under transportability to an insignificant effect when transportability is violated by the amount in the calibrated set ${\cal C}$; we refer to this type of sensitivity as sensitive for an insignificant effect. Overall, 59 counties are sensitive for a positive effect, one county is sensitive for a negative effect, and one county is sensitive for an insignificant effect. The other remaining six counties are insensitive.
In words, the conclusions of the 2024 ad effect can change in 61 counties from their corresponding conclusions under transportability if we consider the magnitudes of unmeasured differences between the sun belt and the rust belt states in 2020. Also, the conclusions of the 2024 ad effect remain unchanged in six counties if we consider the magnitudes of unmeasured differences between the sun belt and the rust belt states in 2020.
Figure 6.2: County-by-county analysis results for 2024 PA voters. Panel A: results under transportability. Panel B: results under two values of $(\Gamma_{0},\Gamma_{1})$. Panel C: an illustration of the analysis pipeline for Philadelphia county and Fulton county. The left column of panel C expands the analysis in panel B for various values of $(\Gamma_{0},\Gamma_{1})$ and each point represents the statistical significance of the ad effect for each $(\Gamma_{0},\Gamma_{1})$; only a few $(\Gamma_{0},\Gamma_{1})$s are displayed for visualization purposes. Gray represents an insignificant effect, blue represents a significant and positive effect, and red represents a significant and negative effect. The middle column of panel C conducts the calibration procedure and the green area is the calibrated set ${\cal C}$. The right column of panel C is the overlap of the two plots and represents the results of the sensitivity analysis with the sensitivity parameters in the calibrated set ${\cal C}$.
Figure 6.3: County-by-county analysis results. Panel A: results of sensitivity analysis under the calibrated set ${\cal C}$. The downward (upward) arrow represents a sensitive effect where the effect changed from an insignificant effect under transportability to a significant and negative (positive) effect. The two gray arrows pointing together represents a change from a significant effect to an insignificant effect. Counties without an arrow are insensitive. Panels B and C: Trump’s share (%) of votes in the 2020 and 2024 U.S. presidential elections, respectively.
6.2.4 Summary of the results and interpretations
From the sensitivity analyses in Section 6.2.2, a small, unmeasured difference between 2020 and 2024 leads to different conclusions about the ad effect in many counties compared to their conclusions under $(\Gamma_{0},\Gamma_{1})=(1,1)$ (i.e., when transportability holds). For example, a small, $0.01$ change in the odds of voting between 2020 and 2024, specifically from $(\Gamma_{0},\Gamma_{1})=(1,1)$ to $(\Gamma_{0},\Gamma_{1})=(1.01,0.99)$, yields many more significant conclusions across counties in PA. Similarly, the sensitivity analysis in Section 6.2.3 based on the calibrated set also suggests that many effects will be sensitive if the odds of voting changed by the amount in the calibrated set. If either the values of the sensitivity parameters in Section 6.2.2 or Section 6.2.3 are plausible, then our paper provides some evidence for the conjecture from Aggarwal et al., (2023) that the ad effect from their 2020 experiment will not generalize to most counties in PA for the 2024 election. Also, based on the direction of the sensitive effects, we believe that with the exception of Philadelphia county, the negative ads against Trump will generally decrease voter turnout in the 2024 election. More generally, because the original effects by Aggarwal et al., (2023) were close to null, we believe our analysis framework is the first to empirically illustrate and validate a simple, but under-appreciated point by Rosenbaum, (2010) that “small effects are sensitive to small [unmeasured] biases” in the context of transfer learning.
Similar to other sensitivity analyses, whether the $(\Gamma_{0},\Gamma_{1})$s considered in Sections 6.2.2 and 6.2.3 are reasonable, unmeasured differences between 2020 and 2024 is at the investigator’s discretion. The two sections provide different ways to interpret the sensitivity analysis and how conclusions of the TATE would change when transportability is violated by a certain amount. We also repeat two cautionary notes from Section 5.3 in that (a) the true unmeasured differences between 2020 and 2024 are not equal to the values of the sensitivity parameters in the two sections, especially the calibrated set from the calibration procedure, and (b) investigators can use different partitions in the calibration procedure to tailor the interpretability of the sensitivity parameter to their specific needs.
Finally, we emphasize that the direction of the ad effect on voter turnout does not equate to whether the ads will lead to less (or more) votes for Trump. This is because Aggarwal et al., (2023) did not measure information about whom a voter voted for. Nevertheless, we can make well-educated conjectures based on comparing the estimated TATEs in Section 6.2 with the shares of votes for Trump across each county; see Figure 6.3. In general, we see that the direction of the sensitive effect roughly corresponds to Trump’s share of votes in the 2020 and 2024 U.S. presidential elections. Philadelphia county, which was declared to be sensitive for a positive effect, has a history of voting for Democratic presidential candidates by large margins. Also, Bedford, Juniata, and Somerset counties, which were declared to be sensitive for a negative effect, voted for Trump by large numbers; in 2020, Trump received 83%, 80%, and 77% of the votes from Bedford, Juniata, and Somerset counties, respectively. However, we caution readers from over-interpreting this connection as Aggarwal et al., (2023) did not measure which candidate a voter voted for.
6.3 Subgroup analysis
After overturning of Roe v. Wade in 2022, many argued that voter turnout will vary substantially by gender and urbanicity, especially compared to past elections (e.g., Shea and Jacobs, (2023)). To study whether the ad effect will also vary by voter demographics, we estimate the effect of running a negative, digital ad campaign against Trump among 20 subgroups of voters. The 20 subgroups are defined by a three-way interaction between gender (female versus not female), urbanicity (rural versus urban), and education attainment (five levels). We use the U.S. Census to obtain information about whether (a) a PA voter lives in a rural or an urban census tract and (b) a PA voter lives in a zipcode with a certain level of educational attainment. Education attainment is categorized by the percentage of people with a Bachelor’s degree or higher and is in increments of 20% (i.e., $(0,20\%],(20,40\%],(40,60\%],(60,80\%],(80,100\%]$). Section E of the Appendix contains further details about the subgroups.
Figure 6.4 summarizes the results. Under transportability, we find some variations in the ad effect among different subgroups of voters, but none of the estimated effects are statistically significant. Voters in urban areas have positive ad effects (i.e., increased voter turnout) regardless of gender and educational attainment and the effects roughly increase with educational attainment. Among voters in rural areas, the ad effect is positive among females living in areas with high educational attainment and the magnitude of this effect is comparable to voters who live in urban areas. The ad effect is most negative (i.e., decreased voter turnout) among female voters living in rural areas with low educational attainment.
When transportability is violated and we conduct a sensitivity analysis with the calibrated set ${\cal C}$, the ad effect is sensitive for a negative effect among female voters living in rural areas with moderate to low educational attainment. The ad effect is sensitive for a positive effect among non-female voters living in urban areas with high educational attainment. The ad effect is sensitive in both directions among non-female voters living in a rural area with high educational attainment. Overall, for the unmeasured differences considered in the calibration set $\mathcal{C}$, the digital ads against Trump will be sensitive among 6 of the 20 subgroups of PA voters.
Figure 6.4: Subgroup analysis by gender, urbanicity, and education attainment in a voter’s zipcode. The horizontal bar represents the 95% CI of the ad effect under transportability. The colored boxes represent the results of the sensitivity analysis with the calibrated set ${\cal C}$.
6.4 Summary of diagnostics and robustness checks
We briefly highlight two additional analyses that we conducted to strengthen our conclusions above. A complete list of all the diagnostics and robustness checks can be found in Sections E to G of the Appendix. In particular, Sections E and F of the Appendix discuss robustness checks related to decisions we made during data pre-processing. Section G conducts a simulation analysis on semi-synthetic data.
First, the analyses based on the OR estimator and the EIF-based estimator were similar, but not identical. For example, in Figure 6.5, we see that for all 67 counties, the point estimates between the OR estimator and the EIF-based estimator fall closely to the 45 degree line and all the 95% confidence intervals generated from the two estimators overlap; note that the widths of the CIs from the two estimators did not uniformly dominate one another. Also, the subgroup analysis based on the EIF estimator was identical to that in Section 6.3 under transportability, and yielded a total of eight sensitive effects, one more for a positive, sensitive effect and one more for a negative sensitive effect. Given the simplicity of the OR estimator and the discreteness of $\mathcal{X}$, we decided to present our findings based on the OR estimator.
Second, the statistical theory that underpins our data analysis assumed that the target and source samples are independent and there are no overlapping voters between the two samples. But in our analysis, it’s plausible that a registered voter in PA for the 2020 election remained a registered voter in PA for the 2024 election. Unfortunately, the source data from Aggarwal et al., (2023) does not identify the voter’s residence exactly. Nevertheless, to allay concerns on potentially overlapping voters, we repeated our analysis with a restricted source data consisting of $n_{s}=662,225$ voters from NC and AZ only. The results from this analysis follow the same trends as above, but with less statistically significant results due to a much smaller sample size. Specifically, the county-by-county analysis results in no counties that are significant under transportability. Even after calibration, no counties are sensitive for positive effects and three fewer counties are sensitive for negative effects than those in Section 6.2. Also, the subgroup analysis did not yield any significant effects after calibration.
While restricting the source data to NC and AZ removes concerns about overlapping voters, it makes transportability less plausible since the target population is less similar to the restricted source data than the original source data that includes PA. Since the source and the target population should be as similar as possible to minimize bias, we decided to report the analysis where the source data contains voters from five states including PA.
Figure 6.5: Comparison between the OR and EIF-based estimators for estimating ad effects for every county in PA. In each panel, x- and y- axes represent results from the OR and the EIF-based estimator, respectively. The points represent point estimates and the gray bars represent 95% CIs. The dashed line represents the 45 degree line through origin (i.e., $y=x$).
7 Discussion and future work
This paper proposes a framework to evaluate political ads based on transfer learning with sensitivity analysis and we use the framework to address whether running a digital ad campaign against Trump is effective in changing voter turnout in PA for the 2024 U.S. presidential election. While not ideal compared to running a randomized trial during the 2024 election, the proposed approach is considerably cheaper as it leverages existing, large-scale experimental data from Aggarwal et al., (2023) and uses sensitivity analysis to account for unmeasurable shifts in context and voter demographics between elections. We present two estimation procedures for the TATE, one based on OR modeling with bootstrapped CIs (i.e., the recommended procedure) and another based on the EIF. For each procedure, we show that it leads to consistent estimates of the TATE and asymptotically valid $1-\alpha$ CIs. Finally, inspired by ideas from design sensitivity, we present a calibration procedure based on partitioning the source population and use it to generate a set of reference magnitudes of the sensitivity parameters for the sensitivity analysis.
Beyond elections, our framework provides statistically valid solutions to important, practical issues that arise in transportability and generalizability, such as dealing with mis-matched covariates between the source and the target population, addressing violation of transportability under $\mathcal{X}\neq\mathcal{V}$, providing a theoretical basis for a commonly used bootstrap procedure in transfer learning, and proposing a new calibration procedure without omitting a covariate; see Section 1.2 for a full list of contributions. However, we point out one important problem we did not address in this paper. Our framework assumes that the units in the target and the source data are sampled from an infinite population of voters. But, in some settings, including our election data, it may be more appropriate to treat the source and the target populations as a finite population. These questions about sampling from a finite population raised several interesting, theoretical questions and due to space constraints and this paper’s emphasis on application, we address them in an upcoming paper.
Finally, as we were finalizing the manuscript during the summer of 2024, the incumbent President Joe Biden has dropped out of the 2024 U.S. presidential election in late July of 2024; our original analysis plan assumed that President Biden is the Democratic Party’s nominee for the presidency. While we believe the interpretations from our analysis is still plausible since Trump was the nominee for the Republication party and the digital ad campaign consisted of negative ads against Trump, we caution readers from over-interpreting the results. Notably, our calibration procedure based on the rust belt and the sun belt states could under-estimate the dramatic shift in electoral context after Biden dropped out of the race and the consequences of this unprecedented event in American politics.
Appendix
Appendix A Extensions and Interpretations of the Sensitivity Model
A.1 Exponential Tilting for Continuous Outcomes
The proposed sensitivity model (2) is not limited to binary outcomes. It can be equivalently expressed for a general, possibly continuous outcome with support ${\cal Y}$. To ease communication, we let $\gamma_{a}={\textnormal{log}}(\Gamma_{a})\in\mathbb{R}$. Suppose the conditional density of the potential outcome on the target population is shifted from that of the source by an exponential tilting shift,
$p_{Y^{(a)}\mid{\bf V},S=0}(y_{a}\mid{\bf v},S_{i}=0)\propto\exp(\gamma_{a}y_{a})\cdot p_{Y^{(a)}\mid{\bf V},S=1}(y_{a}\mid{\bf v},S_{i}=1),\quad\forall{\bf v}\in{\cal V},$ (A.1)
where $\propto$ represents “proportional to” and $p_{Y^{(a)}\mid{\bf V},S=s}$ represents the conditional probability density function of $Y^{(a)}_{i}\mid{\bf V}_{i},S_{i}=s$ for $s=0,1$. When $\gamma_{a}=0$ (i.e., $\Gamma_{a}=1$), (A.1) reduces to $p_{Y^{(a)}\mid{\bf V},S=0}(y_{a}\mid{\bf v},S_{i}=0)=p_{Y^{(a)}\mid{\bf V},S=0}(y_{a}\mid{\bf V}_{i}={\bf v},S_{i}=1)$ and thereby transportability (Assumption 2.3) holds. When $\gamma_{a}\neq 0$, $\gamma_{a}$ measures the violation to the transportability assumption by the degree in shifts of the conditional densities.
Under (A.1) and for a given $\gamma_{a}$, the expected potential outcome under treatment level $a$ can be identified as follows.
Lemma A.1 (Identification of TATE for A General Outcome Under Sensitivity Model)
Suppose Assumptions 2.1, 2.2 and the sensitivity model in equation (A.1) hold. For a given $\gamma_{a}\in\mathbb{R}$, the expected potential outcome under treatment level $a\in\{0,1\}$ is
$\mathbb{E}[Y^{(a)}_{i}\mid S_{i}=0]$ $=\mathbb{E}\left(\dfrac{\mathbb{E}[\mathbb{E}\{\exp(\gamma_{a}Y_{i})Y_{i}\mid{\bf X}_{i},A_{i}=a,S_{i}=1\}\mid{\bf V}_{i},S_{i}=1]}{\mathbb{E}[\mathbb{E}\{\exp(\gamma_{a}Y_{i})\mid{\bf X}_{i},A_{i}=a,S_{i}=1\}\mid{\bf V}_{i},S_{i}=1]}\mid S_{i}=0\right),$ (A.2)
$=\theta_{a}(\gamma_{a}).$
For a binary outcome, Lemma A.1 reduces to Lemma 3.1. When ${\cal X}={\cal V}$, Lemma A.1 recovers the identification result in Dahabreh et al., (2022). When $\gamma_{a}=0$, i.e., transportabiligy holds, Lemma A.1 recovers the identification result in Zeng et al., (2023).
From (A.1), the difference between the two conditional densities at $y_{a}\in{\cal Y}$ is quantified by $\exp(\gamma_{a}y_{a})$ up to some normalizing constant. An extension is to replace $\exp(\gamma_{a}y_{a})$ with $\exp\{\gamma_{a}\delta(y_{a},{\bf v})\}$ where $\delta(y_{a},{\bf v})$ [^3] If $\delta(y_{a},{\bf v},\gamma_{a})$ can be factorized to $\delta_{1}(y_{a},\gamma_{a})\delta_{2}({\bf v},\gamma_{a})$ then it can be replaced with $\delta_{1}(y_{a},\gamma_{a})$. is a statistic including $y_{a}$ and ${\bf v}$. One may also further generalize $\gamma_{a}$ to a vector or generalize the exponential function to other forms based on experts’ knowledge. We note that the choice should ensure the density $p_{Y^{(a)}\mid{\bf V},S=0}$ is well-defined and we refer readers to Franks et al., (2020); Scharfstein et al., (2021) for practical choices.
A.2 Selection Model
An alternative view to the sensitivity model is via the selection to the source, in particular, via the probability of $S_{i}=1$. From this perspective, sensitivity model (A.1) implies a partially linear logistic regression model (Carroll et al.,, 1997) on the selection of $S_{i}$:
$\mathbb{P}(S_{i}=1\mid Y^{(a)}_{i}=y_{a},{\bf V}_{i}={\bf v})=\text{expit}\left(-\gamma_{a}y_{a}-\eta({\bf v},\gamma_{a})\right),\quad\forall y_{a}\in{\cal Y},{\bf v}\in{\cal V},$ (A.3)
$\eta({\bf v},\gamma_{a})={\textnormal{log}}\left(\frac{\mathbb{P}(S_{i}=0)}{\mathbb{P}(S_{i}=1)}\frac{w({\bf v})}{\mathbb{E}\left\{\exp(\gamma_{a}Y^{(a)}_{i})\mid{\bf V}_{i}={\bf v},S_{i}=1\right\}}\right),$
where $\text{expit}(t)=1/\{1+\exp(-t)\}$ for any $t\in\mathbb{R}$ is known as the logistic function. The selection model (A.3) indicates that the participation $S_{i}$ is determined by both the potential outcome and the covariate ${\bf V}_{i}$. After the logistic transformation, the selection probability is associated with $Y^{(a)}_{i}$ linearly with coefficient $\gamma_{a}$. If $\gamma_{a}=0$, then the selection will depend on ${\bf V}_{i}$ only, which reduces to the case where the difference between the target and the source is fully characterized by ${\bf V}_{i}$, i.e., when the transportability holds.
A.3 Estimation for a Continuous Outcome
The identification condition (A.2) directs an OR estimator through
$\widehat{\theta}_{\textnormal{OR,a}}^{\textnormal{cont}}(\gamma_{a})$ $=\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\widehat{\mathbb{E}}\left\{\exp(\gamma_{a}Y^{(a)}_{i})Y^{(a)}_{i}\mid{\bf V}_{i},S_{i}=1\right\}}{\widehat{\mathbb{E}}\left\{\exp(\gamma_{a}Y^{(a)}_{i})\mid{\bf V}_{i},S_{i}=1\right\}}.$
To motivate an EIF-based estimator, we present the EIF in Theorem A.2, which is a generalization of Theorem 4.2 to continuous outcomes.
Theorem A.2
Under Assumptions 2.1 and 2.2 and sensitivity model (A.1), the EIF for $\theta_{a}(\gamma_{a})$ is
$\textnormal{EIF}^{\textnormal{cont}}({\bf O}_{i},\theta_{a}(\gamma_{a}))$
$=$ $\dfrac{S_{i}w({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\left\{\dfrac{A_{i}}{\pi({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\pi({\bf X}_{i})}\right\}\left[\dfrac{\exp(\gamma_{a}Y_{i})Y_{i}}{\mathbb{E}\{\exp(\gamma_{a}Y^{(a)}_{i})\mid{\bf V}_{i},S_{i}=1\}}-\dfrac{\mathbb{E}\{\exp(\gamma_{a}Y_{i})Y_{i}\mid{\bf X}_{i},A=A_{i},S_{i}=1\}}{\mathbb{E}\{\exp(\gamma_{a}Y^{(a)}_{i})\mid{\bf V}_{i},S_{i}=1\}}\right.$
$\left.-\dfrac{\exp(\gamma_{a}Y_{i})\mathbb{E}\{\exp(\gamma_{a}Y^{(a)}_{i})Y^{(a)}_{i}\mid{\bf V}_{i},S_{i}=1\}}{[\mathbb{E}\{\exp(\gamma_{a}Y^{(a)}_{i})\mid{\bf V}_{i},S_{i}=1\}]^{2}}+\dfrac{\mathbb{E}\{\exp(\gamma_{a}Y_{i})\mid{\bf X}_{i},A=A_{i},S_{i}=1\}\mathbb{E}\{\exp(\gamma_{a}Y^{(a)}_{i})Y^{(a)}_{i}\mid{\bf V}_{i},S_{i}=1\}}{[\mathbb{E}\{\exp(\gamma_{a}Y^{(a)}_{i})\mid{\bf V}_{i},S_{i}=1\}]^{2}}\right]$
$+\dfrac{S_{i}w({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\left(\dfrac{\mathbb{E}\{e^{\gamma_{a}Y_{i}}Y_{i}\mid{\bf X}_{i},A=A_{i},S_{i}=1\}}{\mathbb{E}\left\{\exp(\gamma_{a}Y^{(a)}_{i})\mid{\bf V}_{i},S_{i}=1\right\}}-\dfrac{\mathbb{E}\{e^{\gamma_{a}Y^{(a)}_{i}}Y^{(a)}_{i}\mid{\bf V}_{i},S_{i}=1\}\mathbb{E}\{e^{\gamma_{a}Y_{i}}\mid{\bf X}_{i},A=A_{i},S_{i}=1\}}{[\mathbb{E}\{\exp(\gamma_{a}Y^{(a)}_{i})\mid{\bf V}_{i},S_{i}=1\}]^{2}}\right)$
$+\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\left[\dfrac{\mathbb{E}\left\{\exp(\gamma_{a}Y^{(a)}_{i})Y^{(a)}_{i}\mid{\bf V}_{i},S_{i}=1\right\}}{\mathbb{E}\left\{\exp(\gamma_{a}Y^{(a)}_{i})\mid{\bf V}_{i},S_{i}=1\right\}}-\theta_{a}(\gamma_{a})\right].$
$\textnormal{EIF}^{\textnormal{cont}}({\bf O}_{i},\theta_{a}(\gamma_{a}))$ reduces to $\textnormal{EIF}({\bf O}_{i},\theta_{a}(\gamma_{a}))$ in Theorem 4.2 for a binary outcome. It motivates the following EIF-based cross-fitting estimator:
$\widehat{\theta}_{\textnormal{EIF,a}}^{\textnormal{cont}}(\gamma_{a})=\dfrac{1}{K}\sum_{k=1}^{K}\widehat{\theta}_{\textnormal{EIF,a}}^{\textnormal{cont},(k)}(\gamma_{a}),$
where $\widehat{\theta}_{\textnormal{EIF,a}}^{\textnormal{cont},(k)}(\gamma_{a})$ is the estimate at $k$-th partition of the cross-fitting procedure as described in Section 4.2,
$\widehat{\theta}_{\textnormal{EIF,a}}^{\textnormal{cont},(k)}(\gamma_{a})$
$=$ $\dfrac{1}{|{\cal I}_{s}^{(k)}|}\sum_{i\in{\cal I}_{s}^{(k)}}\widehat{w}^{(k)}({\bf V}_{i})\left(\left\{\dfrac{A_{i}}{\widehat{\pi}^{(k)}({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\widehat{\pi}^{(k)}({\bf X}_{i})}\right\}\left[\dfrac{\exp(\gamma_{a}Y_{i})Y_{i}}{\widehat{\mathbb{E}}^{(k)}\{e^{\gamma_{a}Y^{(a)}_{i}}\mid{\bf V}_{i},S_{i}=1\}}-\dfrac{\widehat{\mathbb{E}}^{(k)}\{e^{\gamma_{a}Y_{i}}Y_{i}\mid{\bf X}_{i},A_{i},S_{i}=1\}}{\widehat{\mathbb{E}}^{(k)}\{e^{\gamma_{1}Y^{(a)}_{i}}\mid{\bf V}_{i},S_{i}=1\}}\right.\right.$
$\left.-\dfrac{e^{\gamma_{a}Y_{i}}\widehat{\mathbb{E}}^{(k)}\{e^{\gamma_{a}Y^{(a)}_{i}}Y^{(a)}_{i}\mid{\bf V}_{i},S_{i}=1\}}{[\widehat{\mathbb{E}}^{(k)}\{e^{\gamma_{a}Y^{(a)}_{i}}\mid{\bf V}_{i},S_{i}=1\}]^{2}}+\dfrac{\widehat{\mathbb{E}}^{(k)}\{e^{\gamma_{a}Y_{i}}\mid{\bf X}_{i},A_{i},S_{i}=1\}\widehat{\mathbb{E}}^{(k)}\{e^{\gamma_{a}Y^{(a)}_{i}}Y^{(a)}_{i}\mid{\bf V}_{i},S_{i}=1\}}{\left[\widehat{\mathbb{E}}\{e^{\gamma_{a}Y^{(a)}_{i}}\mid{\bf V}_{i},S_{i}=1\}\right]^{2}}\right]$
$\left.+\dfrac{\widehat{\mathbb{E}}^{(k)}\{e^{\gamma_{a}Y_{i}}Y_{i}\mid{\bf X}_{i},A_{i},S_{i}=1\}\widehat{\mathbb{E}}^{(k)}\{e^{\gamma_{a}Y^{(a)}_{i}}\mid{\bf V}_{i},S_{i}=1\}-\widehat{\mathbb{E}}^{(k)}\{e^{\gamma_{a}Y^{(a)}_{i}}Y^{(a)}_{i}\mid{\bf V}_{i},S_{i}=1\}\widehat{\mathbb{E}}^{(k)}\{e^{\gamma_{a}Y_{i}}\mid{\bf X}_{i},A_{i},S_{i}=1\}}{\left[\widehat{\mathbb{E}}^{(k)}\{\exp(\gamma_{1}Y^{(a)}_{i})\mid{\bf V}_{i},S_{i}=1\}\right]^{2}}\right)$
$+\dfrac{1}{|{\cal I}_{t}^{(k)}|}\sum_{i\in{\cal I}_{t}^{(k)}}\dfrac{\widehat{\mathbb{E}}^{(k)}\left\{\exp(\gamma_{1}Y^{(a)}_{i})Y^{(a)}_{i}\mid{\bf V}_{i},S_{i}=1\right\}}{\widehat{\mathbb{E}}^{(k)}\left\{\exp(\gamma_{1}Y^{(a)}_{i})\mid{\bf V}_{i},S_{i}=1\right\}}.$
Appendix B Details and Proofs for the Outcome Regression Based Estimation
This section provides details and proofs for the inference procedure with the OR estimator proposed in Section 4.2. We detail the bootstrap procedure in Section B.2, state regularity conditions for the bootstrap consistency in Section B.3, and prove Theorem 4.1 in Section B.4.
B.1 Estimation of $\rho_{a}$
We verify that when ${\cal X}$ and ${\cal V}$ are discrete and $\pi$ and $\mu_{a}$ are estimated by group averages, then estimators motivated from (8), (9) and (10) are equivalent. To be explicit, for a given ${\bf x}\in{\cal X}$, the estimates of $\pi$ and $\mu_{a}$ are
$\widehat{\pi}({\bf x})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}A_{i}\mathds{1}({\bf X}_{i}={\bf x})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf X}_{i}={\bf x})},$
$\widehat{\mu}_{a}({\bf x})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}\mathds{1}(A_{i}=a,{\bf X}_{i}={\bf x})Y_{i}}{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{i}=a,{\bf X}_{i}={\bf x})},$
respectively. The equalities in (8), (9), (10) suggest an outcome regression typed estimator $\widehat{\rho}_{a}^{\text{OR}}$, an inverse probability weighting estimator $\widehat{\rho}_{a}^{\text{IPW}}$, and The equality in (8) suggests an outcome regression typed estimator that we denote as an augmented inverse probability weighting estimator $\widehat{\rho}_{a}^{\text{AIPW}}$, respectively, where for ${\bf v}\in{\cal V}$,
$\widehat{\rho}_{a}^{\textnormal{OR}}({\bf v})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})\widehat{\mu}_{a}({\bf X}_{i})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})},$
$\widehat{\rho}_{a}^{\textnormal{IPW}}({\bf v})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}\left\{\dfrac{A_{i}Y_{i}}{\widehat{\pi}({\bf X}_{i})}+\dfrac{(1-A_{i})Y_{i}}{1-\widehat{\pi}({\bf X}_{i})}\right\}\mathds{1}({\bf V}_{i}={\bf v})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})},$
$\widehat{\rho}_{a}^{\textnormal{AIPW}}({\bf v})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}\left[\left\{\dfrac{A_{i}}{\widehat{\pi}({\bf X}_{i})}+\dfrac{(1-A_{i})}{1-\widehat{\pi}({\bf X}_{i})}\right\}\{Y_{i}-\widehat{\mu}_{a}({\bf X}_{i})\}+\widehat{\mu}_{a}({\bf X}_{i})\right]\mathds{1}({\bf V}_{i}={\bf v})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}.$
Lemma B.1
When ${\cal X}$ and ${\cal V}$ are discrete, $\widehat{\rho}_{a}^{\textnormal{OR}}({\bf v})=\widehat{\rho}_{a}^{\textnormal{IPW}}({\bf v})=\widehat{\rho}_{a}^{\textnormal{AIPW}}({\bf v})$ for any ${\bf v}\in{\cal V}$.
Proof of Lemma B.1. Without loss of generality, we prove for the case when $a=1$.
First, we show that $\widehat{\rho}_{1}^{\textnormal{OR}}({\bf v})=\widehat{\rho}_{1}^{\textnormal{IPW}}({\bf v})$. We can simplify $\widehat{\rho}_{1}^{\textnormal{OR}}({\bf v})$ as
$\widehat{\rho}_{1}({\bf v})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})\widehat{\mu}_{a}({\bf X}_{i})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}$
$=\dfrac{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})\cdot\dfrac{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})Y_{j}}{\sum_{k\in{\cal I}_{s}}\mathds{1}(A_{k}=1,{\bf X}_{k}={\bf X}_{i})}}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}$
$=\dfrac{1}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}\sum_{i\in{\cal I}_{s}}\left\{\dfrac{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})Y_{j}}{\sum_{k\in{\cal I}_{s}}\mathds{1}(A_{k}=1,{\bf X}_{k}={\bf X}_{i})}\right\}.$ (B.1)
We can simplify $\widehat{\rho}_{1}^{\textnormal{IPW}}({\bf v})$ as
$\widehat{\rho}_{1}^{\textnormal{IPW}}({\bf v})$ $=\dfrac{\sum_{i\in{\cal I}_{s}}\left\{\dfrac{\mathds{1}(A_{i}=1)Y_{i}}{\widehat{\pi}({\bf X}_{i})}\right\}\mathds{1}({\bf V}_{i}={\bf v})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}$
$=\dfrac{1}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}\cdot\sum_{i\in{\cal I}_{s}}\left[\dfrac{\mathds{1}(A_{i}=1)Y_{i}}{\widehat{\pi}({\bf X}_{i})}\right]$
$=\dfrac{1}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}\cdot\sum_{i\in{\cal I}_{s}}\left\{\dfrac{\mathds{1}(A_{i}=1)Y_{i}}{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1)\mathds{1}({\bf X}_{j}={\bf X}_{i})/\left\{\sum_{k\in{\cal I}_{s}}\mathds{1}({\bf X}_{k}={\bf X}_{i})\right\}}\right\}$
$=\dfrac{1}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})}\cdot\sum_{i\in{\cal I}_{s}}\left\{\dfrac{\sum_{k\in{\cal I}_{s}}\mathds{1}(A_{i}=1)Y_{i}\mathds{1}({\bf X}_{k}={\bf X}_{i})}{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1)\mathds{1}({\bf X}_{j}={\bf X}_{i})}\right\}$ (B.2)
[…]
Next, we show that $\widehat{\rho}_{1}^{\textnormal{AIPW}}({\bf v})=\widehat{\rho}_{1}^{\textnormal{IPW}}({\bf v})$.
$\widehat{\rho}_{1}^{\textnormal{AIPW}}({\bf v})-\widehat{\rho}_{1}^{\textnormal{IPW}}({\bf v})=$ $\dfrac{\sum_{i\in{\cal I}_{s}}\left\{-\dfrac{A_{i}}{\widehat{\pi}({\bf X}_{i})}+1\right\}\widehat{\mu}_{a}({\bf X}_{i})\mathds{1}({\bf V}_{i}={\bf v})}{\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})},$
where the numerator is
$\sum_{i\in{\cal I}_{s}}\left\{-\dfrac{\mathds{1}(A_{i}=1)}{\widehat{\pi}({\bf X}_{i})}+1\right\}\widehat{\mu}_{a}({\bf X}_{i})\mathds{1}({\bf V}_{i}={\bf v})$
$=$ $\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})\widehat{\mu}_{a}({\bf X}_{i})\left\{\dfrac{-\mathds{1}(A_{i}=1)+\dfrac{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})}{\sum_{k\in{\cal I}_{s}}}\mathds{1}({\bf X}_{k}={\bf X}_{i})}{\dfrac{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})}{\sum_{k\in{\cal I}_{s}}\mathds{1}({\bf X}_{k}={\bf X}_{i})}}\right\}$
$=$ $\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})\widehat{\mu}_{a}({\bf X}_{i})\dfrac{-\mathds{1}(A_{i}=1)\sum_{k\in{\cal I}_{s}}\mathds{1}({\bf X}_{k}={\bf X}_{i})+\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})}{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})}$
$=$ $\sum_{i\in{\cal I}_{s}}\mathds{1}({\bf V}_{i}={\bf v})\widehat{\mu}_{a}({\bf X}_{i})\dfrac{-\sum_{k\in{\cal I}_{s}}\mathds{1}(A_{i}=1)\mathds{1}({\bf X}_{k}={\bf X}_{i})+\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})}{\sum_{j\in{\cal I}_{s}}\mathds{1}(A_{j}=1,{\bf X}_{j}={\bf X}_{i})}$
$=$ $0$
Therefore, $\widehat{\rho}_{1}^{\textnormal{AIPW}}({\bf v})=\widehat{\rho}_{1}^{\textnormal{IPW}}({\bf v})$.
$\square$
B.2 Details for the Bootstrap
We detail the nonparametric, percentile bootstrap for the inference with the OR estimator. In each bootstrap iteration, we resample with replacement the source and target samples, respectively, to have sizes $n_{s}$ and $n_{t}$, and construct an OR estimator with the resampled data. After repeating the bootstrap iterations for a large number of times, say $B$ times, we calculate the $\alpha/2$ and $21-\alpha/$ 2 quantiles of the resulting bootstrap estimates, denoted as $\widehat{L}_{a}(\Gamma_{a};1-\alpha)$ and $\widehat{U}_{a}(\Gamma_{a};1-\alpha)$. By Theorem 4.1, the interval $\widehat{\text{CI}}_{\textnormal{OR,a}}(\Gamma_{a})=[\widehat{L}_{a}(\Gamma_{a};1-\alpha),\widehat{U}_{a}(\Gamma_{a};1-\alpha)]$ is a consistent confidence interval for $\theta_{a}(\Gamma_{a})$. A step-by-step procedure is provided in Algorithm 1.
We note that underlying true quantiles of the bootstrap estimates are estimated by their empirical counterparts ($\widehat{L}_{a}(\Gamma_{a};1-\alpha)$ and $\widehat{U}_{a}(\Gamma_{a};1-\alpha)$). This estimation step introduces an additional random error. Since this error can be made arbitrarily small by resampling the data for sufficiently many times, our proof supposes that $\widehat{L}_{a}(\Gamma_{a};1-\alpha)$ and $\widehat{U}_{a}(\Gamma_{a};1-\alpha)$ are the exact quantiles of bootstrap estimates. This argument follows the approach in Chapter 23 of van der Vaart, (1998). For numerical results throughout the paper, the bootstrap iterations are repeated for $B=1000$ times.
Algorithm 1 Outcome regression estimator with nonparametric, percentile bootstrap
1:Sensitivity parameters $\Gamma_{a}$, confidence level $1-\alpha$, bootstrap iteration $B$.
2:Step 1: Estimate $\widehat{\rho}_{a}({\bf v})$ using the source data.
3:Step 2: Estimate $\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})$ as in (4).
4:Step 3: Nonparametric, percentile bootstrap
5:for $b$ in $1,\cdots B$ do
6:     Resample source and target data with replacement at sizes $n_{s}$ and $n_{t}$, respectively.
7:     With the resampled data, obtain $\widehat{\theta}_{\textnormal{OR,a}}^{*,b}(\Gamma_{a})$.
8:end for
9:Calculate the $\alpha/2$ and $21-\alpha/$ 2 quantiles of $\left\{\widehat{\theta}_{\textnormal{OR,a}}^{*,b}(\Gamma_{a})\right\}_{b=1}^{B}$, denoted as $\widehat{L}_{a}(\Gamma_{a};1-\alpha)$ and $\widehat{U}_{a}(\Gamma_{a};1-\alpha)$ where
$\widehat{L}_{a}(\Gamma_{a};1-\alpha)=\widehat{Q}^{*}(\alpha/2),\quad\widehat{U}_{a}(\Gamma_{a};1-\alpha)=\widehat{Q}^{*}(1-\alpha/2),$
$\widehat{Q}^{*}(\tau)=\inf_{t}\left\{\frac{1}{B}\sum_{b=1}^{B}\mathds{1}(\widehat{\theta}_{\textnormal{OR}}^{*,b}(\Gamma_{a})\leq t)\geq\tau\right\},\forall\tau\in(0,1).$
10:The OR estimator $\widehat{\theta}_{\textnormal{OR}}(\Gamma_{a})$ with a $(1-\alpha)$ confidence interval $\widehat{\text{CI}}_{\textnormal{OR,a}}(\Gamma_{a};1-\alpha)=[\widehat{L}_{a}(\Gamma_{a};1-\alpha),\widehat{U}_{a}(\Gamma_{a};1-\alpha)]$.
B.3 Regularity Conditions for the Bootstrap
Recall that we suppose the $\rho_{a}({\bf v})$ is indexed by a finite-dimensional parameter $\boldsymbol{\eta}_{a}$. Specifically, suppose the parameter $\boldsymbol{\eta}_{a}$ is estimated through an estimating equation,
$\dfrac{1}{n_{s}}\sum_{i\in{\cal I}_{s}}{\bf S}({\bf O}_{i},\widehat{\boldsymbol{\eta}}_{a})={\bf 0}$
with a known ${\bf S}({\bf O}_{i},\boldsymbol{\eta}_{a})$. Let ${\boldsymbol{\beta}}_{a}(\Gamma_{a})=[\boldsymbol{\eta}_{a}^{\rm T},\theta_{a}(\Gamma_{a})]^{\rm T}$ and
$\boldsymbol{\phi}_{a}({\bf O}_{i},{\boldsymbol{\beta}}_{a}(\Gamma_{a}))$ $=\left[\dfrac{S_{i}}{\mathbb{P}(S_{i}=1)}{\bf S}({\bf O}_{i},\boldsymbol{\eta}_{a})^{\rm T},\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\phi_{a}({\bf V}_{i},\theta_{a}(\Gamma_{a}),\boldsymbol{\eta}_{a})\right]^{\rm T},\text{where}$
$\phi_{a}({\bf V}_{i},\theta_{a}(\Gamma_{a}),\boldsymbol{\eta}_{a})$ $=\dfrac{\Gamma_{a}\rho_{a}({\bf V}_{i},\boldsymbol{\eta}_{a})}{\Gamma_{a}\rho_{a}({\bf V}_{i},\boldsymbol{\eta}_{a})+1-\rho_{a}({\bf V}_{i},\boldsymbol{\eta}_{a})}-\theta_{a}(\Gamma_{a}).$
Then $\widehat{\boldsymbol{\beta}}_{a}(\Gamma_{a})=[\widehat{\boldsymbol{\eta}}_{a}^{\rm T},\widehat{\theta}_{a}(\Gamma_{a})]^{\rm T}$ can be alternatively expressed as the solution to the estimating equation
$\dfrac{1}{n}\sum_{i=1}^{n}\boldsymbol{\phi}_{a}({\bf O}_{i},\widehat{\boldsymbol{\beta}}(\Gamma_{a}))={\bf 0}.$
We define the bootstrap estimator $\widehat{\boldsymbol{\beta}}^{*}_{a}(\Gamma_{a})$ as the solution to
$\dfrac{1}{n}\sum_{i=1}^{n}W_{n,i}\boldsymbol{\phi}({\bf O}_{i},\widehat{\boldsymbol{\beta}}^{*}_{a}(\Gamma_{a}))={\bf 0},$
where $(W_{n,1},\cdots,W_{n,n_{s}})\sim\text{Multinomial}(n_{s};1/n_{s},\cdots,1/n_{s})$ and
$(W_{n,n_{s}+1},\cdots,W_{n,n})\sim\text{Multinomial}(n_{t};1/n_{t},\cdots,1/n_{t})$.
We assume the following regularity conditions.
  • (B1)
    $\mathbb{E}\{\boldsymbol{\phi}({\bf O}_{i},{\boldsymbol{\beta}}_{a}(\Gamma_{a}))\}={\bf 0}$ with a unique solution ${\boldsymbol{\beta}}(\Gamma_{a})$.
  • (B2)
    Parameter ${\boldsymbol{\beta}}_{a}(\Gamma_{a})$ is contained in a compact parameter space $\Xi$ and $\mathbb{E}\sup_{{\boldsymbol{\beta}}_{a}(\Gamma_{a})\in\Xi}\lVert\boldsymbol{\phi}\rVert_{1}<\infty$.
  • (B3)
    $\mathbb{E}\left(\sup_{{\boldsymbol{\beta}}_{a}(\Gamma_{a})\in\Xi}\lVert\partial\boldsymbol{\phi}^{2}_{a}/\partial{\boldsymbol{\beta}}_{a}(\Gamma_{a})^{2}\rVert\right)<\infty$.
  • (B4)
    The function class $\{\boldsymbol{\phi}_{a}({\bf O}_{i},{\boldsymbol{\beta}}_{a}(\Gamma_{a})),{\boldsymbol{\beta}}_{a}(\Gamma_{a})\in\Xi\}$ is $\mathbb{P}$-Donsker and $\mathbb{E}\lVert\boldsymbol{\phi}_{a}({\bf O}_{i},\widetilde{\boldsymbol{\beta}}(\Gamma_{a}))-\boldsymbol{\phi}_{a}({\bf O}_{i},{\boldsymbol{\beta}}_{a}(\Gamma_{a}))\rVert^{2}\to 0$ as long as $\lVert\widetilde{\boldsymbol{\beta}}_{a}(\Gamma_{a})-{\boldsymbol{\beta}}_{a}(\Gamma_{a})\rVert\to 0$.
Condition (B1) is essentially assuming $\mathbb{E}\{{\bf S}({\bf O},\boldsymbol{\eta}_{a})\}={\bf 0}$ with the unique solution being the true parameter $\boldsymbol{\eta}_{a}$. Condition (B2) guarantees that $\boldsymbol{\phi}_{a}$ is $\mathbb{P}$-Gilvenko-Cantelli by Wellner, (2005, Lemma 6.1). Condition (B3) and (B4) are standard regularity conditions for the complexity of the function class and the smoothness of the estimating equation.
B.4 Proof of Theorem 4.1
Before proving Theorem 4.1, we state the asymptotic Normality of the OR estimator in Theorem B.2. Next in Theorem B.3, we show that the bootstrap estimator is also asymptotically Normal with the same asymptotic variance. Finally we prove the bootstrap CI consistency in Theorem 4.1.
Theorem B.2 (OR estimator)
Suppose Assumptions 2 and 3 hold and $n_{s}\asymp n_{t}$. Also suppose $\rho_{a}({\bf v},\boldsymbol{\eta}_{a})$ is twice differentiable with respect to $\boldsymbol{\eta}_{a}$ and $\widehat{\boldsymbol{\eta}}_{a}$ is an asymptotically linear estimate of $\boldsymbol{\eta}_{a}$ with some influence function ${\bf g}_{a}({\bf O},\boldsymbol{\eta}_{a})$; i.e., $\sqrt{n}(\widehat{\boldsymbol{\eta}}_{a}-\boldsymbol{\eta}_{a})=\dfrac{1}{\sqrt{n}}\sum_{i=1}^{n}{\bf g}({\bf O}_{i},\boldsymbol{\eta}_{a})+o_{p}(1)$. If $\theta_{a}(\Gamma_{a})\in\Theta$ where $\Theta$ is open and compact, then $\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})\to_{p}\theta_{a}(\Gamma_{a})$ and $\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})$ is asymptotically linear with influence function
$\psi_{a}({\bf O}_{i},\theta_{a}(\Gamma_{a}),\boldsymbol{\eta}_{a})$ $=\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\phi_{a}({\bf V}_{i},\theta_{a}(\Gamma_{a}),\boldsymbol{\eta}_{a})+\mathbb{E}\left(\partial\phi_{a}/\partial\boldsymbol{\eta}_{a}^{\rm T}\mid S_{i}=0\right){\bf g}_{a}({\bf O}_{i},\boldsymbol{\eta}_{a}).$
Consequently,
$\sqrt{n}(\widehat{\theta}_{\textnormal{OR,a}}-\theta_{a})$ $=\dfrac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{a}({\bf O}_{i},\theta_{a}(\Gamma_{a}),\boldsymbol{\eta}_{a})+o_{p}(1)\to_{d}N(0,\sigma_{\textnormal{OR,a}}^{2}(\Gamma_{a})),\text{ where}$
$\sigma_{\textnormal{OR,a}}^{2}(\Gamma_{a})$ $=\mathbb{E}\{\psi^{2}_{a}({\bf O}_{i},\theta_{a}(\Gamma_{a}),\boldsymbol{\eta}_{a})\}.$
Proof of Theorem B.2. Without loss of generality, we prove the results for $\theta_{1}(\Gamma_{1})$. We suppress the dependence of $\theta_{1}$ on $\Gamma_{1}$ for notation simplicity.
Since $\Theta$ is compact and $\rho_{1}({\bf v})$ is between zero and one, by Newey and McFadden, (1994, Lemma 2.4), we have that
$\sup_{\theta_{1}\in\Theta}\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{1}\rho_{1}({\bf V}_{i})}{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}-\theta_{1}\right\rVert=o_{p}(1).$
In addition, we note that by the asymptotic linearity of $\widehat{\boldsymbol{\eta}}$,
$\left\lVert\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\right\rVert=\partial\rho_{1}/\partial\boldsymbol{\eta}_{1}^{\rm T}(\boldsymbol{\eta}_{1}-\widehat{\boldsymbol{\eta}}_{1})+o_{p}(1)=o_{p}(1).$ (B.3)
Now we establish consistency by (van der Vaart,, 1998, Theorem 5.9). Note that
$\sup_{\theta_{1}\in\Theta}\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})}{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})}-\theta_{1}\right\rVert$
$\leq$ $\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})}{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})}-\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{1}\rho_{1}({\bf V}_{i})}{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}\right\rVert$
$+\sup_{\theta_{1}\in\Theta}\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{1}\rho_{1}({\bf V}_{i})}{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}-\theta_{1}\right\rVert$
$\leq$ $\dfrac{\Gamma_{1}}{\min\{1,\Gamma_{1}\}}\left\lVert\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\widehat{\rho}_{1}({\bf V}_{i})-\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\rho_{1}({\bf V}_{i})\right\rVert+o_{p}(1),$
$=$ $o_{p}(1),$
where the first inequality follows from triangle inequality, the second inequality follows from the boundedness of $\rho_{1}({\bf v})$ and the compactness of the parameter space, and the last inequality follows from (B.3). By van der Vaart, (1998, Theorem 5.9), $\widehat{\theta}_{1}$ is consistent for $\theta_{1}$.
Finally we prove the asymptotic Normality. With Taylor expansion, we have
$0=$ $\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\phi_{1}({\bf V}_{i},\widehat{\theta}_{1},\widehat{\boldsymbol{\eta}}_{1})$
$=$ $\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\phi_{1}({\bf V}_{i},\theta_{1},\boldsymbol{\eta}_{1})+\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\partial\phi_{1}}{\partial\theta_{1}}(\widehat{\theta}_{1}-\theta_{1})$
$+\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}\dfrac{\partial\phi_{1}}{\partial\boldsymbol{\eta}_{1}^{\rm T}}(\widehat{\boldsymbol{\eta}}_{1}-\boldsymbol{\eta}_{1})+\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}(\widetilde{\boldsymbol{\eta}}_{1}-\boldsymbol{\eta}_{1})^{\rm T}\dfrac{\partial^{2}\phi_{1}}{\partial\boldsymbol{\eta}_{1}\partial\boldsymbol{\eta}_{1}^{\rm T}}(\widetilde{\boldsymbol{\eta}}_{1}-\boldsymbol{\eta}_{1})/2,$
where $\widetilde{\boldsymbol{\eta}}_{1}$ is between $\boldsymbol{\eta}_{1}$ and $\widehat{\boldsymbol{\eta}}_{1}$. Multiplying both sides with $\sqrt{n}$ and rearranging terms, we have
$\sqrt{n}(\widehat{\theta}_{1}-\theta_{1})=$ $\sqrt{n}\dfrac{1-S_{i}}{\widehat{\mathbb{P}}(S_{i}=0)}\phi_{1}({\bf V}_{i},\theta_{1},\boldsymbol{\eta}_{1})+\dfrac{1}{\sqrt{n}}\sum_{i=1}^{n}\dfrac{\mathbb{P}(S_{i}=0)}{\widehat{\mathbb{P}}(S_{i}=0)}\mathbb{E}\left(\partial\phi_{1}/\partial\boldsymbol{\eta}_{1}^{\rm T}\mid S_{i}=0\right){\bf g}_{1}({\bf O}_{i},\boldsymbol{\eta}_{1})+o_{p}(1).$
Since ^$\widehat{\mathbb{P}}(S_{i}=0)=n_{t}/n$ converges to $\mathbb{P}(S_{i}=0)$ almost surely, we have
$\sqrt{n}(\widehat{\theta}_{1}-\theta_{1})=$ $\sqrt{n}\dfrac{S_{i}}{\mathbb{P}(S_{i}=1)}\phi_{1}({\bf V}_{i},\theta_{1},\boldsymbol{\eta}_{1})+\dfrac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbb{E}\left(\partial\phi_{1}/\partial\boldsymbol{\eta}_{1}^{\rm T}\mid S_{i}=0\right){\bf g}_{1}({\bf O}_{i},\boldsymbol{\eta}_{1})+o_{p}(1).$
The proof is completed. $\square$
Next we consider the asymptotic properties for the bootstrap estimator. The resampling procedure during each bootstrap iteration can be viewed as using a weighted sample, where the weights are determined by Multinomial distributions. Therefore, for a bootstrap quantity, for example $\widehat{\theta}_{\textnormal{OR,a}}^{*}(\Gamma_{a})$, there are two sources of randomness: the randomness from the observed data and the randomness from the bootstrap weights. To distinguish between them, until the end of this subsection we denote by $\mathbb{P}_{{\bf O}}$ the probability measure for the observed data and $\mathbb{P}_{W}$ the probability measure for bootstrap weights, and $\mathbb{P}_{{\bf O}W}$ the probability measure on the product space (recall that the bootstrap weights are independent of data). Similar rules apply to the notation of expectations: $\mathbb{E}_{{\bf O}}$, $\mathbb{E}_{W}$ and $\mathbb{E}_{{\bf O}W}$, respectively. A formal treatment of these notations can be found from Cheng and Huang, (2010).
Theorem B.3 (Bootstrap Consistency)
Suppose conditions in Theorem B.2 as well as conditions (B1) and (B2) hold, then $\widehat{\theta}_{a}^{*}(\Gamma_{a})\to\theta_{a}(\Gamma_{a})$ in $\mathbb{P}_{{\bf O}W}$-probability. Suppose additionally conditions (B3) and (B4), then conditional on observations, the bootstrap estimate $\widehat{\theta}_{\textnormal{OR,a}}^{*}(\Gamma_{a})$ satisfies
[…]
Proof of Theorem B.3. We start by proving the consistency, i.e., $\widehat{\theta}_{a}^{*}(\Gamma_{a})\to\theta_{a}(\Gamma_{a})$ in $\mathbb{P}_{{\bf O}W}$-probability. By Lemma 6.1 of Wellner, (2005), condition (B2) guarantees that $\boldsymbol{\phi}_{a}$ is $\mathbb{P}$-Gilvenko-Cantelli. Together with condition (B1), by the multiplier Gilvenko-Cantelli theorem (Vaart and Wellner,, 1996, 3.6.16),
$\sup_{{\boldsymbol{\beta}}_{a}(\Gamma_{a})\in\Xi}\left|\dfrac{1}{n}\sum_{i=1}^{n}W_{n,i}\boldsymbol{\phi}_{a}(\theta_{a}(\Gamma_{a}),\boldsymbol{\eta}_{a})-\mathbb{P}_{{\bf O}}\boldsymbol{\phi}_{a}(\theta_{a}(\Gamma_{a}),\boldsymbol{\eta}_{a})\right|\to 0\text{ in }\mathbb{E}_{{\bf O}W}\text{ probability}.$
Then the consistency for $\widehat{\theta}^{*}_{\textnormal{OR,a}}(\Gamma_{a})$ follows from Corollary 3.2.3 of Vaart and Wellner, (1996).
Next, to prove the asymptotic Normality, it’s sufficient to show
$\sqrt{n}(\widehat{\boldsymbol{\beta}}^{*}_{a}(\Gamma_{a})-\widehat{\boldsymbol{\beta}}_{a}(\Gamma_{a}))\mid\{{\bf O}_{i}\}_{i=1}^{n}\to_{d}N\left(0,{\boldsymbol{\Sigma}}_{a}(\Gamma_{a})\right),$
in $\mathbb{P}_{{\bf O}}$-probability, where
${\boldsymbol{\Sigma}}_{a}(\Gamma_{a})=\mathbb{E}_{{\bf O}}\left\{\dfrac{\partial\boldsymbol{\phi}_{a}({\bf O},{\boldsymbol{\beta}}_{a}(\Gamma_{a}))}{\partial{\boldsymbol{\beta}}_{a}(\Gamma_{a})}\right\}^{-1}\mathbb{E}_{{\bf O}}\{\boldsymbol{\phi}_{a}({\bf O},{\boldsymbol{\beta}}_{a}(\Gamma_{a}))\boldsymbol{\phi}_{a}({\bf O},{\boldsymbol{\beta}}_{a}(\Gamma_{a}))^{\rm T}\}\left[\mathbb{E}_{{\bf O}}\left\{\dfrac{\partial\boldsymbol{\phi}_{a}({\bf O},{\boldsymbol{\beta}}_{a}(\Gamma_{a}))}{\partial{\boldsymbol{\beta}}_{a}(\Gamma_{a})}\right\}^{-1}\right]^{\rm T}.$
From there, the asymptotic Normality of $\widehat{\theta}_{\textnormal{OR,a}}^{*}(\Gamma_{a})$ follows from Delta Method.
To show (B.4), we follow Wellner and Zhan, (1996) or Cheng and Huang, (2010). In particular, the asymptotic Normality in (B.4) holds under regularity conditions (B1) to (B4) and additional conditions (W1) to (W3) on the bootstrap weights:
  • (W1)
[…]
  • (W2)
    $\lim_{\lambda\to\infty}\lim\sup_{n\to\infty}\sup_{t\geq\lambda}t^{2}\mathbb{P}_{W}(W_{ni}\geq t)=0$.
  • (W3)
    $\sum_{i=1}^{n}(W_{ni}-1)^{2}/n\to c$ for some constant $c$.
We are left to verify (W1)-(W3), which can be implied from conditions (W1’)-(W3’) by Lemma 3.1 of Præstgaard and Wellner, (1993).
  • (W1’)
    $\lim\sup_{n\to\infty}\mathbb{E}_{W}(W_{n,i}^{4})<\infty$.
  • (W2’)
    There exists a constant $c$ such that $\mathbb{E}_{W}(W_{ni}^{2})\to 1+c^{2}$.
  • (W3’)
    $\textnormal{Cov}_{W}(W_{n,i}^{2},W_{n,j}^{2})\leq 0$, $i\neq j$.
Finally we verity (W1’)-(W3’). Let $n^{(k)}=n(n-1)\cdots(n-k+1)$ for integer $k$. Without loss of generality suppose $i,j\in{\cal I}_{s}$.
$\mathbb{E}_{W}(W_{n,i}^{2})=$ $2-1/n_{s}\to 2,$
$\mathbb{E}_{W}(W_{n,i}^{4})=$ $1+7n_{s}^{(2)}/n_{s}^{2}+6n_{s}^{(3)}/n_{s}^{3}+n_{s}^{(4)}/n_{s}^{4}\leq 15,$
$\textnormal{Cov}_{W}(W_{ni}^{2},W_{nj}^{2})=$ $\dfrac{1}{n_{s}^{4}}\left[\left\{n_{s}^{(4)}-\left(n_{s}^{(2)}\right)^{2}\right\}+2n_{s}\left\{n_{s}^{(3)}-n_{s}\cdot n_{s}^{(2)}\right\}+n_{s}^{2}\left\{n_{s}^{(2)}-n_{s}^{2}\right\}\right]$
$\leq$ $0$
Hence, (W1’)-(W3’) are satisfied. $\square$
Now we are ready to prove the confidence interval consistency result in Theorem 4.1. This proof resembles the classic proofs for bootstrap CI consistency (Shao and Tu,, 1995; van der Vaart,, 1998).
Proof of Theorem 4.1. The consistency of $\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})$ has been proven in Theorem B.2. Here we prove the bootstrap confidence interval consistency.
Let $\Psi_{a}$ be the cumulative distribution function (c.d.f.) of $N(0,\sigma_{\textnormal{OR,a}}^{2}(\Gamma_{a}))$. Let $\widehat{\Psi}_{a}$ and $\widehat{\Psi}^{*}_{a}$ be the empirical distribution functions of $\sqrt{n}(\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})-\theta_{a}(\Gamma_{a}))$ and $\sqrt{n}(\widehat{\theta}_{\textnormal{OR,a}}^{*}(\Gamma_{a})-\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a}))$, respectively. Then $\widehat{\Psi}_{a}\to_{d}\Psi_{a}$ by Theorem B.2 and $\widehat{\Psi}^{*}_{a}\mid\{{\bf O}_{i}\}_{i=1}^{n}\to_{d}\Psi_{a}$ in $\mathbb{P}_{{\bf O}}$-probability by Theorem B.3. For the latter, there exists a subsequence that converges almost surely. For simplicity we assume the whole sequence converges almost surely; similar arguments have been made in Lemma 23.3 of van der Vaart, (1998) and Cheng and Huang, (2010). Applying the quantile convergence theorem (van der Vaart,, 1998, Lemma 21.2) onto the random distribution functions $\widehat{\Psi}^{*}_{a}$, we have $(\widehat{\Psi}^{*})^{-1}_{a}(\tau)$ converges to $\Psi_{a}^{-1}(\tau)$ almost surely for any $\tau\in(0,1)$. By Slutsky’s theorem,
$\sqrt{n}(\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})-\theta_{a}(\Gamma_{a}))-(\widehat{\Psi}^{*})^{-1}(\alpha/2)\to_{d}N(0,\sigma^{2}_{\textnormal{OR,a}}(\Gamma_{a}))-\Psi^{-1}(\alpha/2).$
Further noting $\sqrt{n}\left(\widehat{L}_{a}(\Gamma_{a})-\widehat{\theta}_{a}(\Gamma_{a})\right)=(\widehat{\Psi}^{*})^{-1}(\alpha/2)$, we have
$\mathbb{P}\left(\widehat{L}_{a}(\Gamma_{a})\leq\theta_{a}(\Gamma_{a})\right)=$ $\mathbb{P}\left(\sqrt{n}\{\widehat{L}_{a}(\Gamma_{a})-\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})\}\leq\sqrt{n}(\theta_{a}(\Gamma_{a})-\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a}))\right)$ (B.4)
$=$ $\mathbb{P}\left((\widehat{\Psi}^{*})^{-1}(\alpha/2)\leq\sqrt{n}\{\theta_{a}(\Gamma_{a})-\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})\}\right)$ (B.5)
$=$ $\mathbb{P}\left(\sqrt{n}\{\theta_{a}-\widehat{\theta}_{\textnormal{OR,a}}(\Gamma_{a})\}\leq-(\widehat{\Psi}^{*})^{-1}(\alpha/2)\right)$ (B.6)
$\to$ $1-\alpha/2\text{ as }n\to\infty.$ (B.7)
The proof of $\mathbb{P}\left(\widehat{U}_{a}(\Gamma_{a})\geq\theta_{a}(\Gamma_{a})\right)\to 1-\alpha/2$ follows similarly and is therefore omitted. The confidence interval consistency follows. $\square$
Appendix C Details and Proofs for the EIF-Based Estimation
In this section we provide details and proofs for the EIF-based estimator $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})$ proposed in Section 4.2 of the main text.
C.1 Implementation Details
For the EIF-based estimator $\widehat{\theta}_{\textnormal{EIF,a}}^{(k)}(\Gamma_{a})$, we remark that the second sum may be replaced by a counterpart that does not use sample splitting, i.e., $\dfrac{1}{|{\cal I}_{t}|}\sum_{i\in{\cal I}_{t}}\dfrac{\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})}{\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})}$, and the resulting estimator will have the same asymptotic distribution as $\widehat{\theta}^{(k)}_{\textnormal{EIF,a}}$. For simplicity of presentation and implementation, we focus our attention on $\widehat{\theta}^{(k)}_{\textnormal{EIF,a}}$ throughout as the first sum in the EIF estimator requires sample splitting.
A step-by-step implementation of the EIF-based estimation is provided in Algorithm 2.
Algorithm 2 EIF-Based Estimation with Cross-Fitting
1:${\cal I}_{s}$, ${\cal I}_{t}$; integer $K\geq 2$; sensitivity parameters $\Gamma_{1},\Gamma_{0}$; confidence level $(1-\alpha)$.
2:Step 1 (Partitioning): Randomly split ${\cal I}_{s}$ and ${\cal I}_{t}$ to ${\cal I}_{s,k}$ and ${\cal I}_{t,k}$, respectively, $1\leq k\leq K$.
3:Step 2 (Cross Fitting):
4:for $k$ in $1,2,\cdots K$ do
5:     With $\{{\cal I}_{s}\backslash{\cal I}_{s,k}\}\cup\{{\cal I}_{t}\backslash{\cal I}_{t,k}\}$, obtain $\widehat{\pi}^{(k)}({\bf X}_{i})$, $\widehat{w}^{(k)}({\bf V}_{i})$, $\widehat{\mu}^{(k)}_{a}({\bf X}_{i})$ and $\widehat{\rho}^{(k)}({\bf V}_{i})$.
6:     With ${\cal I}_{s,k}\cup{\cal I}_{t,k}$, calculate $\widehat{\theta}_{\textnormal{EIF}}^{(k)}(\Gamma_{0},\Gamma_{1})=\widehat{\theta}_{\textnormal{EIF,1}}^{(k)}(\Gamma_{1})-\widehat{\theta}_{\textnormal{EIF,0}}^{(k)}(\Gamma_{0})$ with $\widehat{\theta}_{\textnormal{EIF,a}}^{(k)}(\Gamma_{a})$ in Section 4.2.
7:end for
8:Step 3 (Building Estimator): Construct the estimator $\widehat{\theta}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})=\dfrac{1}{K}\sum_{k=1}^{K}\widehat{\theta}_{\textnormal{EIF}}^{(k)}(\Gamma_{0},\Gamma_{1})$.
9:Step 4 (Variance Estimation): Construct the variance estimator
$\widehat{\sigma}^{2}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})=\dfrac{1}{K}\sum_{k=1}^{K}\left[\dfrac{1}{|{\cal I}_{k}|}\sum_{i\in{\cal I}_{k}}\left\{\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,1}}(\Gamma_{1}))-\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,0}}(\Gamma_{0}))\right\}^{2}\right].$
10:The EIF-based estimator $\widehat{\theta}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})$ with a $(1-\alpha)$ confidence interval
$\left(\widehat{\theta}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})-z_{\alpha/2}\widehat{\sigma}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})/\sqrt{n},\ \widehat{\theta}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})+z_{1-\alpha/2}\widehat{\sigma}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})/\sqrt{n}\right),$
where $z_{\beta}$ is the $\beta$ quantile of the standard Normal distribution for any $\beta\in(0,1)$ and $\widehat{\sigma}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})=\sqrt{\widehat{\sigma}^{2}_{\textnormal{EIF}}(\Gamma_{0},\Gamma_{1})}$.
C.2 Estimating the Density Ratio
The estimation of the density ratio can proceed in two methods falling into two categories. The first category is to recognize the relationship between $w({\bf v})$ and $\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})$ via the Bayes rule, i.e.,
$w({\bf v})$ $=\dfrac{\mathbb{P}(S_{i}=1)}{\mathbb{P}(S_{i}=0)}\dfrac{\mathbb{P}(S_{i}=0\mid{\bf V}_{i}={\bf v})}{\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})},$ (C.1)
and estimate $w({\bf v})$ by estimating $\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})$ with a binary classifier and estimating $\mathbb{P}(S_{i}=1)$ as $n_{s}/n$; see Kallus and Mao, (2024) and Zeng et al., (2023). For example, when ${\cal V}$ is discrete, one may estimate this probability for any ${\bf v}\in{\cal V}$ by calculating the proportion of source samples among all samples with the same covariate:
^$\widehat{\mathbb{P}}(S_{i}=1\mid{\bf V}_{i}={\bf v})=\dfrac{\sum_{i=1}^{n}\mathds{1}(S_{i}=1,{\bf V}_{i}={\bf v})}{\sum_{i=1}^{n}\mathds{1}({\bf V}_{i}={\bf v})},\ \ \widehat{w}({\bf v})=\dfrac{n_{s}}{n_{t}}\dfrac{\widehat{\mathbb{P}}(S_{i}=0\mid{\bf V}_{i}={\bf v})}{\widehat{\mathbb{P}}(S_{i}=1\mid{\bf V}_{i}={\bf v})}.$ (C.2)
Equation (C.1) also reveals the necessity of having a sufficiently large target sample (i.e., the second part of Assumption 2.2). Intuitively, a substantially small target sample will make estimation of $\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})$ challenging due to class imbalance. Also, when $\mathbb{P}(S_{i}=0)$ is close to zero, $w({\bf v})$ can be large in magnitude, which will generally increase the bias and variance of the estimated TATE.
The second category is to use principles behind covariate balance to estimate $w({\bf v})$. Specifically, $w$ serves as a balancing score between the source and the target population, i.e.,
$\mathbb{E}\{f({\bf V}_{i})w({\bf V}_{i})\mid S_{i}=1\}=\mathbb{E}\{f({\bf V}_{i})\mid S_{i}=0\},\text{ any measurable }f.$
Han et al., (2021) considered this connection to construct an exponential tilting estimator of $w({\bf V}_{i})$. Relatedly, Josey et al., (2022); Chen et al., (2023) used entropy balancing of Hainmueller, (2012) to estimate $w({\bf V}_{i})$.
To account for the possible imbalance between the source and target samples (i.e., $n_{s}$ and $n_{t}$ may differ a lot) and to enable covariate balancing, we proceed with the entropy balancing method in (12) that falls into the second category. The solutions $\widehat{w}_{i}$ of entropy balancing are characterized in Lemma C.1.
Lemma C.1
The solution of (12) is $\widehat{w}_{i}=\exp(\widehat{\alpha}+\widehat{\boldsymbol{\beta}}^{\rm T}{\bf V}_{i})$, where $(\widehat{\alpha},\widehat{\boldsymbol{\beta}})$ is solution to
$\min_{\alpha,{\boldsymbol{\beta}}}\dfrac{1}{n_{s}}\sum_{i\in{\cal I}_{s}}\exp(\alpha+{\boldsymbol{\beta}}^{\rm T}{\bf V}_{i})-\alpha-\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}{\boldsymbol{\beta}}^{\rm T}{\bf V}_{i}.$ (C.3)
Lemma C.1 is a special case of Proposition 1 of Chen et al., (2023). The dual problem (C.3) is an unconstrained convex optimization problem and numeric solutions can be efficiently solved by algorithms like the Newton-Raphson method. The implementation is performed using the optim function in R.
C.3 Proof of Theorem 4.2 and Theorem A.2
We prove Theorem A.2, the EIF for a general outcome under sensitivity model (A.1). It includes Theorem 4.2 as a special case for a binary outcome. To simplify notation, we suppress the dependence of the TATE on $\Gamma_{a}$ and denote the expected potential outcome on the target population at treatment level $a$ as $\theta_{a}$ for $a=0,1$. We also drop the subscript $i$ and denote by ${\bf O}$ a generic random variable, which consists of $({\bf X},Y,S=1)$ for the source and $({\bf V},S=0)$ for the target. We recall that we have defined $\gamma_{a}={\textnormal{log}}(\Gamma_{a})$.
We start with the case where $\pi({\bf X})$ is unknown and therefore considered as a nuisance parameter. For clarify we denote its true value as $\pi_{0}({\bf X})$. Denote by $p_{{\bf V}\mid S=1}$, $p_{{\bf X}\mid{\bf V},S=1}$, $p_{Y\mid{\bf X},A,S=1}$ the density functions of the conditional distributions of ${\bf V}\mid S$, ${\bf X}\mid{\bf V},S=1$ and $Y\mid{\bf X},A,S=1$, respectively. For a generic observation ${\bf O}$, the log-likelihood can be written as
$l({\bf O})=$ $(1-S){\textnormal{log}}\left(p_{{\bf V}\mid S=0}({\bf V}\mid S=0)\right)+S{\textnormal{log}}\left(p_{{\bf V}\mid S=1}({\bf V}\mid S=1)\right)$
$+S{\textnormal{log}}\left(p_{{\bf X}\mid{\bf V},S=1}({\bf X}\mid{\bf V},S=1)\right)+AS{\textnormal{log}}\left(\pi({\bf X})\right)+S(1-A){\textnormal{log}}(1-\pi({\bf X}))$
$+SA{\textnormal{log}}\left(p_{Y\mid{\bf X},A=1,S=1}(Y\mid{\bf X},A=1,S=1)\right)+S(1-A){\textnormal{log}}\left(p_{Y\mid{\bf X},A=0,S=1}(Y\mid{\bf X},A=0,S=1)\right).$
Consider the Hilbert space $\Lambda$ that contains all one-dimensional zero-mean measurable functions of the observed data with finite variance. Consider $p_{Y\mid{\bf X},A=0,S=1}$, $p_{Y\mid{\bf X},A=1,S=1}$, $\pi({\bf X})$, $p_{{\bf X}\mid{\bf V},S=1}$, $p_{{\bf V}\mid S=0}$ and $p_{{\bf V}\mid S=1}$ as nuisance functions and denote their nuisance tangent spaces as $\Lambda_{Y\mid{\bf X},A=1,S=1}$, $\Lambda_{Y\mid{\bf X},A=0,S=1}$, $\Lambda_{\pi}$, $\Lambda_{{\bf X}\mid S=1}$, $\Lambda_{{\bf V}\mid S=1}$ and $\Lambda_{{\bf V}\mid S=0}$, respectively. Then
$\Lambda=\Lambda_{Y\mid{\bf X},A=1,S=1}\oplus\Lambda_{Y\mid{\bf X},A=0,S=1}\oplus\Lambda_{\pi}\oplus\Lambda_{{\bf X}\mid S=1}\oplus\Lambda_{{\bf V}\mid S=1}\oplus\Lambda_{{\bf V}\mid S=0},$
where $\oplus$ is the direct sum between orthogonal spaces, and
$\Lambda_{Y\mid{\bf X},A=1,S=1}$ =$=\left\{SA{\bf b}_{1}(Y,{\bf X}):\mathbb{E}\left[{\bf b}_{1}(Y,{\bf X})\mid{\bf X},A=1,S=1\right]={\bf 0}\right\},$
$\Lambda_{Y\mid{\bf X},A=0,S=1}$ =$=\left\{S(1-A){\bf b}_{2}(Y,{\bf X}):\mathbb{E}\left[{\bf b}_{2}(Y,{\bf X})\mid{\bf X},A=0,S=1\right]={\bf 0}\right\},$
$\Lambda_{\pi}$ =$=\left\{S[A-\pi_{0}({\bf x})]{\bf b}_{3}({\bf X}):0<\pi_{0}({\bf X})<1\right\},$
$\Lambda_{{\bf X}\mid S=1}$ =$=\left\{S{\bf b}_{4}({\bf X}):\mathbb{E}\left[{\bf b}_{4}({\bf X})\mid{\bf V},S=1\right]={\bf 0}\right\},$
$\Lambda_{{\bf V}\mid S=1}$ =$=\left\{S{\bf b}_{5}({\bf V}):\mathbb{E}\left[{\bf b}_{5}({\bf V})\mid S=1\right]={\bf 0}\right\},$
$\Lambda_{{\bf V}\mid S=0}$ =$=\left\{(1-S){\bf b}_{6}({\bf V}):\mathbb{E}\left[{\bf b}_{6}({\bf V})\mid S=0\right]={\bf 0}\right\}.$
Without loss of generality, we derive the EIF for $\theta_{1}$. The EIF for $\theta_{0}$ is analogous and thus omitted for brevity. Consider parametric submodels indexed by parameter ${\boldsymbol{\alpha}}$ where ${\boldsymbol{\alpha}}={\bf 0}$ represents the true data generating process. We re-express the log-likelihood under the parametric submodel,
$l({\bf O},{\boldsymbol{\alpha}})=$ $(1-S){\textnormal{log}}p_{{\bf V}\mid S=0}({\bf V}\mid S=0;{\boldsymbol{\alpha}})+S{\textnormal{log}}p_{{\bf V}\mid S=1}({\bf V}\mid S=1;{\boldsymbol{\alpha}})$
$+S{\textnormal{log}}p_{{\bf X}\mid{\bf V},S=1}({\bf X}\mid{\bf V},S=1;{\boldsymbol{\alpha}})+AS{\textnormal{log}}\pi({\bf x};{\boldsymbol{\alpha}})+S(1-A){\textnormal{log}}(1-\pi({\bf X};{\boldsymbol{\alpha}}))$
$+SA{\textnormal{log}}p_{Y\mid{\bf X},A=1,S=1}(Y\mid{\bf X},A=1,S=1;{\boldsymbol{\alpha}})$
$+S(1-A){\textnormal{log}}p_{Y\mid{\bf X},A=0,S=1}(Y\mid{\bf X},A=0,S=1;{\boldsymbol{\alpha}}).$
Define the score function
${\bf S}({\bf O})=$ $\dfrac{\partial l({\bf O},{\boldsymbol{\alpha}})}{\partial{\boldsymbol{\alpha}}}\Bigg{|}_{{\boldsymbol{\alpha}}={\bf 0}}$
$=$ $SA{\bf S}_{1}(Y,{\bf X})+S(1-A){\bf S}_{2}(Y,{\bf X})+S\dfrac{\partial\left[\{A{\textnormal{log}}(\pi({\bf X};{\boldsymbol{\alpha}}))+(1-A){\textnormal{log}}(1-\pi({\bf X};{\boldsymbol{\alpha}}))\right]}{\partial{\boldsymbol{\alpha}}}\Bigg{|}_{{\boldsymbol{\alpha}}={\bf 0}}$
$+S{\bf S}_{4}({\bf X})+S{\bf S}_{5}({\bf V})+(1-S){\bf S}_{6}({\bf V}),\text{ where}$
${\bf S}_{1}(Y,{\bf X})=$ $\dfrac{\partial{\textnormal{log}}p_{Y\mid{\bf X},A=1,S=1}(Y\mid{\bf X},A=1,S=1;{\boldsymbol{\alpha}})}{\partial{\boldsymbol{\alpha}}}\Bigg{|}_{{\boldsymbol{\alpha}}={\bf 0}},$
${\bf S}_{2}(Y,{\bf X})=$ $\dfrac{\partial{\textnormal{log}}p_{Y\mid{\bf X},A=0,S=1}(Y\mid{\bf X},A=0,S=1;{\boldsymbol{\alpha}})}{\partial{\boldsymbol{\alpha}}}\Bigg{|}_{{\boldsymbol{\alpha}}={\bf 0}},$
${\bf S}_{4}({\bf X})=$ $\dfrac{\partial{\textnormal{log}}p_{{\bf X}\mid{\bf V},S=1}({\bf X}\mid{\bf V},S=1;{\boldsymbol{\alpha}})}{\partial{\boldsymbol{\alpha}}}\Bigg{|}_{{\boldsymbol{\alpha}}={\bf 0}},$
${\bf S}_{5}({\bf V})=$ $\dfrac{\partial{\textnormal{log}}p_{{\bf V}\mid S=1}({\bf V}\mid S=1;{\boldsymbol{\alpha}})}{\partial{\boldsymbol{\alpha}}}\Bigg{|}_{{\boldsymbol{\alpha}}={\bf 0}},$
${\bf S}_{6}({\bf V})=$ $\dfrac{\partial{\textnormal{log}}p_{{\bf V}\mid S=0}({\bf V}\mid S=0;{\boldsymbol{\alpha}})}{\partial{\boldsymbol{\alpha}}}\Bigg{|}_{{\boldsymbol{\alpha}}={\bf 0}},$
and $SA{\bf S}_{1}(Y,{\bf X})\in\Lambda_{Y\mid{\bf X},A=1,S=1}$, $S(1-A){\bf S}_{2}(Y,{\bf X})\in\Lambda_{Y\mid{\bf X},A=0,S=1}$, $S{\bf S}_{4}({\bf X})\in\Lambda_{{\bf X}\mid S=1}$, $S{\bf S}_{5}({\bf V})\in\Lambda_{{\bf V}\mid S=1}$, $(1-S){\bf S}_{6}({\bf V})\in\Lambda_{{\bf V}\mid S=0}$.
Next, we show that
$\mathbb{E}\left[\phi^{\textnormal{cont}}_{1}({\bf O},\theta_{1}){\bf S}({\bf O})\right]$ $=\dfrac{\partial\theta_{1}}{\partial{\boldsymbol{\alpha}}}\Bigg{|}_{{\boldsymbol{\alpha}}={\bf 0}},$ (C.4)
where
$\phi^{\textnormal{cont}}_{1}({\bf O},\theta_{1}(\gamma_{1}))$
$=$ $\dfrac{Sw({\bf V})}{\mathbb{P}(S=1)\pi({\bf X})}\left[\dfrac{\exp(\gamma_{1}Y)Y}{\mathbb{E}\{\exp(\gamma_{1}Y^{(1)})\mid{\bf V},S=1\}}-\dfrac{\mathbb{E}\{\exp(\gamma_{1}Y)Y\mid{\bf X},A=1,S=1\}}{\mathbb{E}\{\exp(\gamma_{1}Y^{(1)})\mid{\bf V},S=1\}}\right.$
$\left.-\dfrac{\exp(\gamma_{1}Y)\mathbb{E}\{\exp(\gamma_{1}Y^{(1)})Y^{(1)}\mid{\bf V},S=1\}}{[\mathbb{E}\{\exp(\gamma_{1}Y^{(1)})\mid{\bf V},S=1\}]^{2}}+\dfrac{\mathbb{E}\{\exp(\gamma_{1}Y)\mid{\bf X},A=1,S=1\}\mathbb{E}\{\exp(\gamma_{1}Y^{(1)})Y^{(1)}\mid{\bf V},S=1\}}{[\mathbb{E}\{\exp(\gamma_{1}Y^{(1)})\mid{\bf V},S=1\}]^{2}}\right]$
$+\dfrac{Sw({\bf V})}{\mathbb{P}(S=1)}\dfrac{\mathbb{E}\{e^{\gamma_{1}Y}Y\mid{\bf X},A=1,S=1\}\mathbb{E}\{e^{\gamma_{1}Y^{(1)}}\mid{\bf V},S=1\}-\mathbb{E}\{e^{\gamma_{1}Y^{(1)}}Y^{(1)}\mid{\bf V},S=1\}\mathbb{E}\{e^{\gamma_{1}Y}\mid{\bf X},A=1,S=1\}}{[\mathbb{E}\{\exp(\gamma_{1}Y^{(1)})\mid{\bf V},S=1\}]^{2}}$
$+\dfrac{1-S}{\mathbb{P}(S=0)}\left[\dfrac{\mathbb{E}\left\{\exp(\gamma_{1}Y^{(1)})Y^{(1)}\mid{\bf V},S=1\right\}}{\mathbb{E}\left\{\exp(\gamma_{1}Y^{(1)})\mid{\bf V},S=1\right\}}-\theta_{1}\right].$
To show (C.4), we calculate its right-hand side:
$\dfrac{\partial\theta_{1}}{\partial{\boldsymbol{\alpha}}}\Bigg{|}_{{\boldsymbol{\alpha}}={\bf 0}}=$ $\mathbb{E}\left(w({\bf V})\mathbb{E}\left[\mathbb{E}\{B_{1}(Y^{(1)},{\bf X}){\bf S}_{1}(Y,{\bf X})\mid{\bf X},A=1,S=1\}\mid{\bf V},S=1\right]\mid S=1\right)$ (C.5)
$+\mathbb{E}\left[\mathbb{E}\left\{w({\bf V})B_{4}({\bf X}){\bf S}_{4}({\bf X})\mid{\bf V},S=1\right\}\mid S=1\right]$ (C.6)
$+\mathbb{E}\left\{\mathbb{E}(Y^{(1)}{\bf S}_{6}({\bf V})\mid S=0)\right\},$ (C.7)
where
$B_{1}(Y^{(1)},{\bf X})=$ $\dfrac{e^{\gamma_{1}Y^{(1)}}Y^{(1)}}{\mathbb{E}\{e^{\gamma_{1}Y^{(1)}}\mid{\bf V},S=1\}}-\dfrac{e^{\gamma_{1}Y^{(1)}}\mathbb{E}\{e^{\gamma_{1}Y^{(1)}}Y^{(1)}\mid{\bf V},S=1\}}{[\mathbb{E}\{e^{\gamma_{1}Y^{(1)}}\mid{\bf V},S=1\}]^{2}},$
$B_{4}({\bf X})=$ $\dfrac{\mathbb{E}\{e^{\gamma_{1}Y}Y\mid{\bf X},A=1,S=1\}\mathbb{E}\{e^{\gamma_{1}Y^{(1)}}\mid{\bf V},S=1\}}{[\mathbb{E}\{e^{\gamma_{1}Y^{(1)}}\mid{\bf V},S=1\}]^{2}}$
$-\dfrac{\mathbb{E}\{e^{\gamma_{1}Y^{(1)}}Y^{(1)}\mid{\bf V},S=1\}\mathbb{E}\{e^{\gamma_{1}Y}\mid{\bf X},A=1,S=1\}}{[\mathbb{E}\{e^{\gamma_{1}Y^{(1)}}\mid{\bf V},S=1\}]^{2}}.$
Further, note that
[…]
$=\mathbb{E}\left(\dfrac{SAw({\bf V})}{\mathbb{P}(S=1)\pi({\bf X})}\left[B_{1}(Y^{(1)},{\bf X})-\mathbb{E}\{B_{1}(Y^{(1)},{\bf X})\mid{\bf X},A=1,S=1\}\right]{\bf S}({\bf O})\right),$
[…]
$=\mathbb{E}\left\{\dfrac{1-S}{\mathbb{P}(S=0)}\left[\dfrac{\mathbb{E}\left\{\exp(\gamma_{1}Y^{(1)})Y^{(1)}\mid{\bf v},S=1\right\}}{\mathbb{E}\left\{\exp(\gamma_{1}Y^{(1)})\mid{\bf v},S=1\right\}}-\theta_{1}\right]{\bf S}({\bf O})\right\},$
we have
[…]
Finally, we verify that $\phi^{\textnormal{cont}}_{1}({\bf O},\theta_{1})\in\Lambda$ since
$\dfrac{SAw({\bf V})}{\mathbb{P}(S=1)\pi({\bf X})}\left[B_{1}(y^{(1)},{\bf x})-\mathbb{E}\{B_{1}(Y^{(1)},{\bf X})\mid{\bf x},A=1,S=1\}\right]\in\Lambda_{Y\mid{\bf X},A=1,S=1},$
$\dfrac{S}{\mathbb{P}(S=1)}\left\{B_{4}({\bf X})-\mathbb{E}[B_{4}({\bf X})\mid{\bf v},S=1]\right\}\in\Lambda_{{\bf X}\mid S=1},\text{ and}$
$\dfrac{1-S}{\mathbb{P}(S=0)}\left[\dfrac{\mathbb{E}\left\{\exp(\gamma_{1}Y^{(1)})Y^{(1)}\mid{\bf V},S=1\right\}}{\mathbb{E}\left\{\exp(\gamma_{1}Y^{(1)})\mid{\bf V},S=1\right\}}-\theta_{1}\right]\in\Lambda_{{\bf V}\mid S=0}.$
Therefore, $\phi^{\textnormal{cont}}_{1}({\bf O},\theta_{1})$ is the EIF in Theorem A.2, i.e., $\textnormal{EIF}^{\textnormal{cont}}_{1}({\bf O},\theta_{1})$. Moreover, if the outcome is binary, we can re-express the followings:
$\mathbb{E}\{\exp(\gamma_{1}Y)Y\mid{\bf X},A=1,S=1\}=$ $\Gamma_{1}\mu_{1}({\bf X}),$
$\mathbb{E}\{\exp(\gamma_{1}Y)\mid{\bf X},A=1,S=1\}=$ $\Gamma_{1}\mu_{1}({\bf X})+1-\mu_{1}({\bf X}),$
$\mathbb{E}\{\exp(\gamma_{1}Y^{(1)})Y^{(1)}\mid{\bf V},S=1\}=$ $\Gamma_{1}\rho_{1}({\bf V}),$
$\mathbb{E}\{\exp(\gamma_{1}Y^{(1)})\mid{\bf V},S=1\}=$ $\Gamma_{1}\rho_{1}({\bf V})+1-\rho_{1}({\bf V}).$
Plugging in them to $\textnormal{EIF}^{\textnormal{cont}}({\bf O},\theta_{1})$ yields $\textnormal{EIF}({\bf O},\theta_{1})$ as the expression of the EIF for a binary outcome in Theorem 4.2.
Next, we suppose $\pi({\bf X})$ is known as its true value $\pi_{0}({\bf X})$. Then $\pi({\bf X})$ is no longer considered as a nuisance function and the Hilbert space $\Lambda$ can now be decomposed as
$\Lambda=\Lambda_{Y\mid{\bf X},A=1,S=1}\oplus\Lambda_{Y\mid{\bf X},A=0,S=1}\oplus\Lambda_{{\bf X}\mid S=0}\oplus\Lambda_{{\bf V}\mid S=1}\oplus\Lambda_{{\bf V}\mid S=0}.$
Under the parametric submodel, the log-likelihood becomes
$l({\bf O},{\boldsymbol{\alpha}})=$ $(1-S){\textnormal{log}}p_{{\bf V}\mid S=0}({\bf V}\mid S=0;{\boldsymbol{\alpha}})+S{\textnormal{log}}p_{{\bf V}\mid S=1}({\bf V}\mid S=1;{\boldsymbol{\alpha}})$
$+S{\textnormal{log}}p_{{\bf X}\mid{\bf V},S=1}({\bf X}\mid{\bf V},S=1;{\boldsymbol{\alpha}})+AS{\textnormal{log}}\pi_{0}({\bf X})+S(1-A){\textnormal{log}}(1-\pi_{0}({\bf X}))$
$+SA{\textnormal{log}}p_{Y\mid{\bf X},A=1,S=1}(Y\mid{\bf X},A=1,S=1;{\boldsymbol{\alpha}})$
$+S(1-A){\textnormal{log}}p_{Y\mid{\bf X},A=0,S=1}(Y\mid{\bf X},A=0,S=1;{\boldsymbol{\alpha}}).$
Then the score function becomes
${\bf S}({\bf O})=$ $\dfrac{\partial l({\bf O},{\boldsymbol{\alpha}})}{\partial{\boldsymbol{\alpha}}}\Bigg{|}_{{\boldsymbol{\alpha}}={\bf 0}}$
$=$ $SA{\bf S}_{1}(Y,{\bf X})+S(1-A){\bf S}_{2}(Y,{\bf X})+S{\bf S}_{4}({\bf X})+S{\bf S}_{5}({\bf V})+(1-S){\bf S}_{6}({\bf V}),$
where we still have $SA{\bf S}_{1}(Y,{\bf X})\in\Lambda_{Y\mid{\bf X},A=1,S=1}$, $S(1-A){\bf S}_{2}(Y,{\bf X})\in\Lambda_{Y\mid{\bf X},A=0,S=1}$, $S{\bf S}_{4}({\bf X})\in\Lambda_{{\bf X}\mid S=1}$, $S{\bf S}_{5}({\bf V})\in\Lambda_{{\bf V}\mid S=1}$, $(1-S){\bf S}_{6}({\bf V})\in\Lambda_{{\bf V}\mid S=0}$. Therefore, $\mathbb{E}\left[\phi^{\textnormal{cont}}_{1}({\bf O},\theta_{1}){\bf S}({\bf O})\right]=\dfrac{\partial\theta_{1}}{\partial{\boldsymbol{\alpha}}}\Bigg{|}_{{\boldsymbol{\alpha}}={\bf 0}}$ holds following the same argument as we’ve shown.
C.4 Lemma C.2
In this section we characterize the plug-in bias for the EIF-based estimator $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})$. For the generality of the conclusion and to avoid overloading the notation, we assume the nuisance functions are estimated from an independent sample. We introduce the general notation for the uncentered EIF,
$\varphi_{a}({\bf O}_{i})$
$=$ $\dfrac{S_{i}w({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\dfrac{\Gamma_{a}}{[\Gamma_{a}\rho_{a}({\bf V}_{i})+1-\rho_{a}({\bf V}_{i})]^{2}}\left[\left\{\dfrac{A_{i}}{\pi({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\pi({\bf X}_{i})}\right\}\{Y_{i}-\mu_{a}({\bf X}_{i})\}+\mu_{a}({\bf X}_{i})-\rho_{a}({\bf V}_{i})\right]$
$+\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\dfrac{\Gamma_{a}\rho_{a}({\bf V}_{i})}{\Gamma_{a}\rho_{a}({\bf V}_{i})+1-\rho_{a}({\bf V}_{i})},$
and its estimate
$\widehat{\varphi}_{a}({\bf O}_{i})$
$=$ $\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)}\dfrac{\Gamma_{a}}{[\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})]^{2}}\left[\left\{\dfrac{A_{i}}{\widehat{\pi}({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\widehat{\pi}({\bf X}_{i})}\right\}\{Y_{i}-\widehat{\mu}_{a}({\bf X}_{i})\}+\widehat{\mu}_{a}({\bf X}_{i})-\widehat{\rho}_{a}({\bf V}_{i})\right]$
$+\dfrac{1-S_{i}}{\widehat{\mathbb{P}}(S_{i}=0)}\dfrac{\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})}{\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})},$
where ^$\widehat{\mathbb{P}}(S_{i}=1)=n_{s}/n$, $\widehat{\mu}_{a}({\bf X}_{i})$, $\widehat{\rho}_{a}({\bf V}_{i})$, $\widehat{\pi}({\bf X}_{i})$ and $\widehat{w}({\bf V}_{i})$ are estimated from an independent sample.
Lemma C.2
There exists a constant $C$ such that
$|\mathbb{E}\{\widehat{\varphi}_{a}({\bf O}_{i})-\varphi_{a}({\bf O}_{i})\}|\leq$ $C\left(\lVert\widehat{\mu}_{a}({\bf X}_{i})-\mu_{a}({\bf X}_{i})\rVert\cdot\lVert\widehat{\pi}({\bf X}_{i})-\pi({\bf X}_{i})\rVert+\lVert\widehat{\rho}_{a}({\bf V}_{i})-\rho_{a}({\bf V}_{i})\rVert\cdot\lVert\widehat{w}({\bf V}_{i})-w({\bf V}_{i})\rVert\right.$
$+\left.\lVert\widehat{\rho}_{a}({\bf V}_{i})-\rho_{a}({\bf V}_{i})\rVert^{2}\right).$
In particular, if $\Gamma_{1}=\gamma_{0}=1$, there exists a constant $C$ such that
$|\mathbb{E}\{\widehat{\varphi}_{a}({\bf O}_{i})-\varphi_{a}({\bf O}_{i})\}|\leq$ $C\left(\lVert\widehat{\mu}_{a}({\bf X}_{i})-\mu_{a}({\bf X}_{i})\rVert\cdot\lVert\widehat{\pi}({\bf X}_{i})-\pi({\bf X}_{i})\rVert+\lVert\widehat{\rho}_{a}({\bf V}_{i})-\rho_{a}({\bf V}_{i})\rVert\cdot\lVert\widehat{w}({\bf V}_{i})-w({\bf V}_{i})\rVert\right).$
Proof of Lemma C.2: Without loss of generality, we prove the case for $a=1$.
$\mathbb{E}\{\widehat{\varphi}_{1}({\bf O}_{i})-\varphi_{1}({\bf O}_{i})\}$
$=$ $\mathbb{E}\{\widehat{\varphi}_{1}({\bf O}_{i})\}-\theta_{1}$
$=$ $\mathbb{E}\{\widehat{\varphi}_{1}({\bf O}_{i})\}-\mathbb{E}\left[\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\dfrac{\Gamma_{1}\rho_{1}({\bf V}_{i})}{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}\right]$
$=$ $\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)}\dfrac{\Gamma_{1}}{[\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\dfrac{A}{\widehat{\pi}({\bf X}_{i})}\{\mu_{1}({\bf X}_{i})-\widehat{\mu}_{1}({\bf X}_{i})\}\right]$
$+\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)}\dfrac{\Gamma_{1}}{[\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})]^{2}}\{\widehat{\mu}_{1}({\bf X}_{i})-\mu_{1}({\bf X}_{i})\}\right]$
$-\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)}\dfrac{\Gamma_{1}}{[\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\}\right]$
$+\mathbb{E}\left[\dfrac{1-S_{i}}{\widehat{\mathbb{P}}(S_{i}=0)}\dfrac{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})}{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})}-\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\dfrac{\Gamma_{1}\rho_{1}({\bf V}_{i})}{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}\right]$
$=$ $\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)\widehat{\pi}({\bf X}_{i})}\dfrac{\Gamma_{1}}{[\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\{\widehat{\pi}({\bf X}_{i})-\pi({\bf X}_{i})\}\{\widehat{\mu}_{1}({\bf X}_{i})-\mu_{1}({\bf X}_{i})\}\right]$
$-\mathbb{E}\left[\left\{\dfrac{\mathbb{P}(S_{i}=1)-\widehat{\mathbb{P}}(S_{i}=1)}{\widehat{\mathbb{P}}(S_{i}=1)\mathbb{P}(S_{i}=1)}+\dfrac{1}{\mathbb{P}(S_{i}=1)}\right\}S_{i}\widehat{w}({\bf V}_{i})\dfrac{\Gamma_{1}}{[\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\}\right]$
$+\mathbb{E}\left(\dfrac{(1-S_{i})\Gamma_{1}\left[\widehat{\rho}_{1}({\bf V}_{i})\{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}-\rho_{1}({\bf V}_{i})\{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}\right]}{\mathbb{P}(S_{i}=0)\widehat{\mathbb{P}}(S_{i}=0)\{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}\{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}}\right)$
$\leq$ $O(1)\cdot\mathbb{E}\left[\{\widehat{\pi}({\bf X}_{i})-\pi({\bf X}_{i})\}\{\widehat{\mu}_{1}({\bf X}_{i})-\mu_{1}({\bf X}_{i})\}\right]$
$-\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\dfrac{\Gamma_{1}}{[\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\}\right]$
$+\mathbb{E}\left[\dfrac{(1-S_{i})}{\mathbb{P}(S_{i}=0)}\dfrac{\Gamma_{1}}{\{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}\{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}}\left\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\right\}\right]$
$\leq$ $O(1)\cdot\mathbb{E}\left[\{\widehat{\pi}({\bf X}_{i})-\pi({\bf X}_{i})\}\{\widehat{\mu}_{1}({\bf X}_{i})-\mu_{1}({\bf X}_{i})\}\right]$
$+\mathbb{E}\left[\dfrac{S_{i}}{\mathbb{P}(S_{i}=1)}\dfrac{\Gamma_{1}\{1-\Gamma_{1}\}\left\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\right\}\left\{\widehat{w}({\bf V}_{i})-w({\bf V}_{i})\right\}}{\{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}^{2}\{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}}\right]$
$+\mathbb{E}\left[\dfrac{Sw({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\dfrac{\Gamma_{1}\{1-\Gamma_{1}\}\left\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\right\}^{2}}{\{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}^{2}\{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}}\right]$
$\leq$ $O(1)\cdot\mathbb{E}\left[\{\widehat{\pi}({\bf X}_{i})-\pi({\bf X}_{i})\}\{\widehat{\mu}_{1}({\bf X}_{i})-\mu_{1}({\bf X}_{i})\}\right]+O(1)\cdot\mathbb{E}\left[\{\widehat{w}({\bf V}_{i})-w({\bf V}_{i})\}\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\}\right]$
$+O(1)\cdot\mathbb{E}\left[\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\}^{2}\right]$
$\leq$ $O(1)\left\{\lVert\widehat{\mu}_{a}({\bf X}_{i})-\mu_{a}({\bf X}_{i})\rVert\cdot\lVert\widehat{\pi}({\bf X}_{i})-\pi({\bf X}_{i})\rVert+\lVert\widehat{\rho}_{a}({\bf V}_{i})-\rho_{a}({\bf V}_{i})\rVert\cdot\lVert\widehat{w}({\bf V}_{i})-w({\bf V}_{i})\rVert\right.$
$\left.+\lVert\widehat{\rho}_{a}({\bf V}_{i})-\rho_{a}({\bf V}_{i})\rVert^{2}\right\}$
When $\Gamma_{1}=0$, following the same procedure and using the fact that $\Gamma_{1}=1$, we have
$\mathbb{E}\{\widehat{\varphi}_{1}({\bf O}_{i})-\varphi_{1}({\bf O}_{i})\}$
$\leq$ $O(1)\left\{\lVert\widehat{\mu}_{a}({\bf X}_{i})-\mu_{a}({\bf X}_{i})\rVert\cdot\lVert\widehat{\pi}({\bf X}_{i})-\pi({\bf X}_{i})\rVert+\lVert\widehat{\rho}_{a}({\bf V}_{i})-\rho_{a}({\bf V}_{i})\rVert\cdot\lVert\widehat{w}({\bf V}_{i})-w({\bf V}_{i})\rVert\right\}.$
C.5 Proof of Theorem 4.3
The EIF-based estimator is $\widehat{\theta}_{\textnormal{EIF}}(\Gamma_{a})=\widehat{\theta}_{\textnormal{EIF,1}}-\widehat{\theta}_{\textnormal{EIF,0}}$ with $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})=\dfrac{1}{K}\sum_{k=1}^{K}\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\widehat{\varphi}({\bf O}_{i})$. Without loss of generality, we consider the proof for $\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a})$ and drop $\Gamma_{a}$ in notation for simplicity. We have
$\widehat{\theta}_{\textnormal{EIF,a}}-\theta_{a}$ =$=\left\{\dfrac{1}{K}\sum_{k=1}^{K}\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\widehat{\varphi}_{a}^{(k)}({\bf O}_{i})-\theta_{a}\right\}$ (C.8)
=$=\left\{\dfrac{1}{K}\sum_{k=1}^{K}\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})\right\}$ (C.9)
$=\dfrac{1}{n}\sum_{i=1}^{n}\textnormal{EIF}({\bf O}_{i},\theta_{a})+\left\{\dfrac{1}{K}\sum_{k=1}^{K}\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}\textnormal{EIF}({\bf O}_{i})\right\}$ (C.10)
$=\dfrac{1}{n}\sum_{i=1}^{n}\textnormal{EIF}({\bf O}_{i},\theta_{a})+\left\{\dfrac{1}{K}\sum_{k=1}^{K}\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\left[\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\textnormal{EIF}({\bf O}_{i},\theta_{a})\right]\right\}.$
We define
$R_{k}=\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\left\{\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\textnormal{EIF}({\bf O}_{i},\theta_{a})\right\},\text{ for }k=1,\cdots K.$
C.5.1 Part (i)
Since $K$ is independent of data, to show that $\widehat{\theta}_{\textnormal{EIF,a}}$ is consistent, it suffices to show
$R_{1}=o_{p}(1).$
From Lemma C.2,
$\mathbb{E}(R_{1})\leq O(1)$$\cdot\left\{\lVert\widehat{w}^{(k)}({\bf V}_{i})-w^{(k)}({\bf V}_{i})\rVert\cdot\lVert\widehat{\rho}^{(k)}_{a}({\bf V}_{i})-\rho^{(k)}_{a}({\bf V}_{i})\rVert+\lVert\widehat{\rho}^{(k)}_{a}({\bf V}_{i})-\rho^{(k)}_{a}({\bf V}_{i})\rVert^{2}\right\}$
$+O(1)\cdot\lVert\widehat{\pi}^{(k)}({\bf X}_{i})-\pi^{(k)}({\bf X}_{i})\rVert\cdot\lVert\widehat{\mu}_{a}^{(k)}({\bf X}_{i})-\mu_{a}^{(k)}({\bf X}_{i})\rVert$
$\leq$ $o_{p}(1),$
where the second inequality follows from the conditions that $\lVert\widehat{\rho}^{(k)}_{a}({\bf V}_{i})-\rho^{(k)}_{a}({\bf V}_{i})\rVert=o_{p}(1)$ and (6). Next, we show $R_{1}-\mathbb{E}(R_{1})=o_{p}(1)$. Conditioning on ${\cal I}_{k}^{c}={\cal I}\backslash{\cal I}_{k}$, we calculate the mean and variance for $R_{1}-\mathbb{E}(R_{1})$:
$\mathbb{E}\{R_{1}-\mathbb{E}(R_{1})\mid{\cal I}_{k}^{c}\}$ $=\mathbb{E}\left[\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\mathbb{E}\{\widehat{\phi}^{(k)}_{\textnormal{EIF,a}}({\bf O}_{i},\theta_{a})\}\mid{\cal I}_{k}^{c}\right]-\mathbb{E}\left[\textnormal{EIF}({\bf O}_{i},\theta_{a})-\mathbb{E}\{\textnormal{EIF}({\bf O}_{i},\theta_{a})\}\right]$
$=0,$
$\textnormal{Var}(R_{1}-\mathbb{E}(R_{1})\mid{\cal I}_{k}^{c})$ $=\textnormal{Var}(R_{1}\mid{\cal I}_{k}^{c})\leq K\lVert\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\textnormal{EIF}({\bf O}_{i},\theta_{a})\rVert^{2}/n.$
Then for any $\varepsilon>0$, by Chebyshev’s inequality,
$\mathbb{P}\left(\dfrac{R_{1}-\mathbb{E}(R_{1})}{\lVert\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\textnormal{EIF}({\bf O}_{i},\theta_{a})\rVert/\sqrt{n}}\geq\varepsilon\right)$ $=\mathbb{E}\left\{\mathbb{P}\left(\dfrac{R_{1}-\mathbb{E}(R_{1})}{K\lVert\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\textnormal{EIF}({\bf O}_{i},\theta_{a})\rVert/\sqrt{n}}\geq\varepsilon\mid{\cal I}_{k}^{c}\right)\right\}$
$\leq 1/\varepsilon^{2}.$
Therefore,
$R_{1}-\mathbb{E}(R_{1})$ $=KO_{p}(\lVert\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\textnormal{EIF}({\bf O}_{i},\theta_{a})\rVert)/\sqrt{n}\leq O_{p}(1/\sqrt{n})=o_{p}(1).$
C.5.2 Part (ii)
The decomposition at the beginning of the proof suggests
$\sqrt{n}(\widehat{\theta}_{\textnormal{EIF,a}}-\theta_{a})$ $=\dfrac{1}{\sqrt{n}}\sum_{i=1}^{n}\textnormal{EIF}({\bf O}_{i},\theta_{a})+\sqrt{n}\left\{\dfrac{1}{K}\sum_{k=1}^{K}\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\left[\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\textnormal{EIF}({\bf O}_{i},\theta_{a})\right]\right\}$
Since $K$ is independent of the data, it suffices to show
$R_{1}=o_{p}(n^{-1/2}).$
From Lemma C.2 and the rate conditions (7a), (7b) and (7c) in Theorem 4.3, we have
$\mathbb{E}(R_{1})=o_{p}(n^{-1/2}).$
In what follows we show $R_{1}-\mathbb{E}(R_{1})=o_{p}(n^{-1/2})$. Conditioning on ${\cal I}_{k}^{c}={\cal I}\backslash{\cal I}_{k}$, we calculate the mean and variance for $R_{1}-\mathbb{E}(R_{1})$:
$\mathbb{E}\{R_{1}-\mathbb{E}(R_{1})\mid{\cal I}_{k}^{c}\}$ $=\mathbb{E}\left[\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\mathbb{E}\{\widehat{\phi}^{(k)}_{\textnormal{EIF,a}}({\bf O}_{i},\theta_{a})\}\mid{\cal I}_{k}^{c}\right]-\mathbb{E}\left[\textnormal{EIF}({\bf O}_{i},\theta_{a})-\mathbb{E}\{\textnormal{EIF}({\bf O}_{i},\theta_{a})\}\right]$
$=0,$
$\textnormal{Var}(R_{1}-\mathbb{E}(R_{1})\mid{\cal I}_{k}^{c})$ $=\textnormal{Var}(R_{1}\mid{\cal I}_{k}^{c})\leq K\lVert\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\textnormal{EIF}({\bf O}_{i},\theta_{a})\rVert^{2}/n.$
Then for any $\varepsilon>0$, by Chebyshev’s inequality,
$\mathbb{P}\left(\dfrac{R_{1}-\mathbb{E}(R_{1})}{\lVert\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\textnormal{EIF}({\bf O}_{i},\theta_{a})\rVert/\sqrt{n}}\geq\varepsilon\right)$ $=\mathbb{E}\left\{\mathbb{P}\left(\dfrac{R_{1}-\mathbb{E}(R_{1})}{K\lVert\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\textnormal{EIF}({\bf O}_{i},\theta_{a})\rVert/\sqrt{n}}\geq\varepsilon\mid{\cal I}_{k}^{c}\right)\right\}$
$\leq 1/\varepsilon^{2}.$
Since all nuisance parameters are consistently estimated by assumption (i.e., $\lVert\widehat{\rho}^{(k)}_{a}({\bf V}_{i})-\rho^{(k)}_{a}({\bf V}_{i})\rVert=o_{p}(1)$, $\lVert\widehat{\mu}_{a}^{(k)}({\bf X}_{i})-\mu_{a}^{(k)}({\bf X}_{i})\rVert=o_{p}(1)$, $\lVert\widehat{w}^{(k)}({\bf V}_{i})-w^{(k)}({\bf V}_{i})\rVert=o_{p}(1)$, $\lVert\widehat{\pi}^{(k)}({\bf X}_{i})-\pi^{(k)}({\bf X}_{i})\rVert=o_{p}(1)$), Lemma C.2 suggests that $\lVert\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\textnormal{EIF}({\bf O}_{i},\theta_{a})\rVert=o_{p}(1)$. Therefore,
$R_{1}-\mathbb{E}(R_{1})$ $=KO_{p}(\lVert\widehat{\textnormal{EIF}}^{(k)}({\bf O}_{i},\theta_{a})-\textnormal{EIF}({\bf O}_{i},\theta_{a})\rVert)/\sqrt{n}=o_{p}(1/\sqrt{n}).$
C.5.3 Part (iii)
In order to show
$\widehat{\sigma}^{2}_{\textnormal{EIF,a}}(\Gamma_{a})-\sigma^{2}_{\textnormal{EIF,a}}(\Gamma_{a})=\dfrac{1}{K}\sum_{k=1}^{K}\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\widehat{\textnormal{EIF}}^{2}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))-\mathbb{E}\{\textnormal{EIF}^{2}({\bf O}_{i},\theta_{a}(\Gamma_{a}))\}=o_{p}(1),$
it’s sufficient to show
$R_{k,1}-R_{k,2}=\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\widehat{\textnormal{EIF}}^{2}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))-\mathbb{E}\{\textnormal{EIF}^{2}({\bf O}_{i},\theta_{a}(\Gamma_{a}))\}=o_{p}(1),$ (C.11)
where
$R_{k,1}$ $=\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\left\{\widehat{\textnormal{EIF}}^{2}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))-\textnormal{EIF}^{2}({\bf O}_{i},\theta_{a}(\Gamma_{a}))\right\},$
$R_{k,2}$ $=\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\left[\textnormal{EIF}^{2}({\bf O}_{i},\theta_{a}(\Gamma_{a}))-\mathbb{E}\{\textnormal{EIF}^{2}({\bf O}_{i},\theta(\Gamma_{a}))\}\right].$
(C.11) can be concluded since $R_{k,2}=O_{p}(n^{-1/2})$ by $\mathbb{E}\{\textnormal{EIF}^{4}({\bf O}_{i},\theta(\Gamma_{a}))\}<\infty$, and $R_{k,2}=O_{p}(n^{-1/2})$ by the following argument. Note that
$|R_{k,1}|\leq$ $\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\left|\widehat{\textnormal{EIF}}^{2}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))-\textnormal{EIF}^{2}({\bf O}_{i},\theta_{a}(\Gamma_{a}))\right|$
$=$ $\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\left|\widehat{\textnormal{EIF}}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))-\textnormal{EIF}({\bf O}_{i},\theta_{a}(\Gamma_{a}))\right|\cdot\left|\widehat{\textnormal{EIF}}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))+\textnormal{EIF}({\bf O}_{i},\theta_{a}(\Gamma_{a}))\right|$
$\leq$ $\sqrt{\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\left|\widehat{\textnormal{EIF}}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))-\textnormal{EIF}({\bf O}_{i},\theta(\Gamma_{a}))\right|^{2}}\sqrt{\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\left|\widehat{\textnormal{EIF}}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))+\textnormal{EIF}({\bf O}_{i},\theta(\Gamma_{a}))\right|^{2}}$
$\leq$ $\sqrt{\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\left|\widehat{\textnormal{EIF}}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))-\textnormal{EIF}({\bf O}_{i},\theta_{a}(\Gamma_{a}))\right|^{2}}\left(\sqrt{\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\left|\widehat{\textnormal{EIF}}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))-\textnormal{EIF}({\bf O}_{i},\theta(\Gamma_{a}))\right|^{2}}\right.$
$\left.+\sqrt{\frac{4}{|{\cal I}_{k}|}\sum_{i\in{\cal I}_{k}}\phi^{2}_{\textnormal{EIF,a}}({\bf O}_{i},\theta(\Gamma_{a}))}\right),$
we have
$R_{k,1}^{2}\lesssim R_{n}\left\{\frac{4}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}_{k}}\textnormal{EIF}^{2}({\bf O}_{i},\theta_{a}(\Gamma_{a}))+R_{n}\right\}$
where $R_{n}=\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\left|\widehat{\textnormal{EIF}}({\bf O}_{i},\widehat{\theta}_{\textnormal{EIF,a}}(\Gamma_{a}))-\textnormal{EIF}({\bf O}_{i},\theta_{a}(\Gamma_{a}))\right|^{2}$. Since $\dfrac{1}{|{\cal I}^{(k)}|}\sum_{i\in{\cal I}^{(k)}}\textnormal{EIF}^{2}({\bf O}_{i},\theta_{a}(\Gamma_{a}))=O_{p}(1)$, it’s sufficient to show $R_{n}=O_{p}(n^{-1/2})$, which holds by the proof of Theorem 4.3.
Appendix D Details and Examples of the Calibration Procedure
This section provides details and illustrations for the calibration procedure introduced in Section 5.
D.1 Analysis Pipeline
We start with some remarks about the implementation of our calibration procedure. First, it’s important to have the ratio of the sample sizes between the proxy source and target data be equal to that of the original source and the target data. This can be accomplished by downsampling one of the two proxy data. Relatedly, to make the comparisons fairer, it’s useful to rescale the standard error estimate in the transported CI from the calibration procedure by multiplying it with $\sqrt{|{\cal I}_{s_{2}}|/n_{t}}$ in order to mimic the length of the CI for the original TATE. This was mentioned in Algorithm 1 under Step 1. Third, one should make sure the shared covariates in constructing $\widehat{\text{CI}}_{s\to t}(\Gamma_{0},\Gamma_{1};1-\alpha)$ should match the shared covariates ${\bf V}_{i}$ in the actual target sample. See Algorithm 3 for the implementation and Section D of the Supplementary Materials for more discussions. Algorithm 3 provides a step-by-step procedure for calibrating the sensitivity parameters. As an example, Figure D.1 illustrates ${\cal C}_{1}$, ${\cal C}_{2}$ and the final calibration region ${\cal C}$ for estimating the ad effect in Philadelphia in 2024.
Algorithm 3 Calibrating Sensitivity Parameters
1:Source data, confidence level $1-\alpha$, set ${\cal C}_{\text{all}}\in\mathbb{R}\times\mathbb{R}$.
2:Step 1 (Partition source data): Partition the source data into two parts and denote their corresponding indices as ${\cal I}_{s_{1}}$ and ${\cal I}_{s_{2}}$ where ${\cal I}_{s_{1}}\cup{\cal I}_{s_{2}}={\cal I}_{s}$ and ${\cal I}_{s_{1}}\cap{\cal I}_{s_{2}}=\emptyset$.
3:if ${\cal I}_{s_{1}}/{\cal I}_{s_{2}}>n_{s}/n_{t}$ then
4:     Randomly subset ${\cal I}_{s_{1}}$ of size $|{\cal I}_{s_{2}}|\cdot n_{s}.n_{t}$ and denote the resulting set of indices as ${\cal I}_{s_{1}}$.
5:else
6:     Randomly subset ${\cal I}_{s_{2}}$ of size $|{\cal I}_{s_{1}}|\cdot n_{t}/n_{s}$ and denote the resulting set of indices as ${\cal I}_{s_{2}}$.
7:end if
8:Step 2.1 (Construct CI via the standard approach: With data in ${\cal I}_{s_{2}}$, estimate the ATE and its $(1-\alpha)$ confidence interval, denoted as $\widehat{\text{CI}}_{s_{2}}(1-\alpha)$.
9:Step 2.2 (Construct CI via our transfer learning approach) :
10:With $\{({\bf X}_{i},A_{i},Y_{i},S_{i}=1):i\in{\cal I}_{s_{1}}\}\cup\{({\bf V}_{i},S_{i}=0):i\in{\cal I}_{s_{2}}\}$, estimate the ATE on ${\cal S}_{2}$ and its standard error with any $(\Gamma_{0},\Gamma_{1})\in{\cal C}_{\text{all}}$, denoted as $\widehat{\theta}_{s_{1}\to s_{2}}(\Gamma_{0},\Gamma_{1})$ and $\widehat{\text{SE}}_{s_{1}\to s_{2}}$. Denote the re-scaled confidence interval as
$\widehat{\text{CI}}_{s_{1}\to s_{2}}(\Gamma_{0},\Gamma_{1};1-\alpha)=\left[\widehat{\theta}_{s_{1}\to s_{2}}(\Gamma_{0},\Gamma_{1})\mp z_{1-\alpha/2}\cdot\widehat{\text{SE}}_{s_{1}\to s_{2}}(\Gamma_{0},\Gamma_{1})\cdot\sqrt{|{\cal I}_{s_{2}|}/n_{t}}\right].$ (D.1)
11:Step 3 (Find the plausible range) : Find the plausible range of sensitivity parameters when transporting from ${\cal S}_{1}$ to ${\cal S}_{2}$:
${\cal C}_{1}=\left\{(\Gamma_{0},\Gamma_{1})\in{\cal C}_{\text{all}}:\widehat{\text{CI}}_{s_{2}}\cap\widehat{\text{CI}}_{s_{1}\to s_{2}}(\Gamma_{0},\Gamma_{1})\neq\emptyset\right\}.$ (D.2)
12:Calibration in the other direction Exchange ${\cal S}_{1}$ and ${\cal S}_{2}$ and repeat Steps 1-3, resulting in the plausible range ${\cal C}_{2}$.
13:Intersect two plausible regions to construct the final region: ${\cal C}={\cal C}_{1}\cap{\cal C}_{2}$.
Figure D.1: The calibration procedure for the analysis in Philadelphia in 2024. Panels from left to right plots ${\cal C}_{1}$, ${\cal C}_{2}$ and ${\cal C}={\cal C}_{1}\cap{\cal C}_{2}$ in shadowed areas along $\Gamma_{1}$ in the y-axis and $\Gamma_{0}$ in the x-axis.
D.2 Interpretations
The sensitivity parameters $\Gamma_{0}$ and $\Gamma_{1}$ quantify the change in turnout from 2020 to 2024 in the control arm and the treatment arm, respectively, and different values of $\Gamma_{0}$ and $\Gamma_{1}$ will generally correspond to different effect sizes and direction. Some examples are listed below and Table D.1 enumerates more examples.
  1. 1.
    Suppose $\Gamma_{0}=1$ and $\Gamma_{1}>1$ (i.e., the $y>1$ part in Figure D.1). Then the turnout in 2024 if voters are not exposed to anti-Trump ads will be the same as that in 2020, but the turnout in 2024 if the voters are exposed to anti-Trump ads will be larger than that in 2020. Also, the ad effect in 2024 will be higher than that in 2020 and if $\Gamma_{1}$ is sufficiently large, the effect will be positive and statistically significant.
  2. 2.
    Suppose $\Gamma_{0}>1$ and $\Gamma_{1}=1$ (i.e., the $x>1$ part in Figure D.1). Then the turnout in 2024 if voters are not exposed to anti-Trump ads will be higher than that in 2020, but the turnout in 2024 if the voters are exposed to anti-Trump ads will be the same as that in 2020. Also, the ad effect in 2024 is likely smaller than that in 2020, and if $\Gamma_{0}$ is large enough, the effect may be negative and significant.
  3. 3.
    Suppose $\Gamma_{0}>1$ and $\Gamma_{1}>1$ (i.e., top right region of Figure D.1). Then, the odds of turnout in both treatment and controls will be higher in 2024 than those in 2020. In this case, the ad effect in 2024 may be similar to that in 2020, especially if the shift in the turnouts between 2024 and 2020 are comparable between the control and treatment arms. A similar phenomena would occur if $\Gamma_{0}<1$ and $\Gamma_{1}<1$ (i.e., bottom left region of Figure D.1).
  4. 4.
    Suppose $\Gamma_{0}<1$ and $\Gamma_{1}>1$ (i.e., top left region of Figure D.1). Then, the odd of turnout if voters are not exposed to anti-Trump ads will be lower in 2024 than that in 2020, but the odd of turnout if voter are exposed to anti-Trump ads will be higher than 2024 than that in 2020. Then, the combined effect of the changes in the odds would be a large and positive value of the ad effect in 2024.
  5. 5.
    Suppose $\Gamma_{0}>1$ and $\Gamma_{1}<1$ (i.e., bottom right region of Figure D.1). Then, the odd of turnout if voters are not exposed to anti-Trump ads will be higher in 2024 than that in 2020, but the odd of turnout if voter are exposed to anti-Trump ads will be lower than 2024 than that in 2020. Then, the combined effect of the changes in the odds would be a negative ad effect in 2024 that is large in magnitude.
Table D.1: Examples on the signs of $\Gamma_{0}$, $\Gamma_{1}$ and the ad effect in 2024 (i.e., TATE) compared with the ad effect in 2020.
$\Gamma_{0}$ $\Gamma_{1}$ Odd of turnout in 2024 if unexposed to negative ads (i.e., $Y^{(0)}$) Odd of turnout in 2024 if exposed to negative ads (i.e., $Y^{(1)}$) ATE in 2024 (i.e., TATE)
0 $>0$ same as the corresponding odd in 2020 higher than the corresponding odd in 2020 higher than ATE in 2020
$>1$ $=1$ the corresponding odd in 2020 same as the corresponding odd in 2020 lower than ATE in 2020
$>1$ $>1$ higher than the corresponding odd in 2020 higher than the corresponding odd in 2020 may be similar with ATE in 2020
$<1$ $>1$ lower than the corresponding odd in 2020 higher than the corresponding odd in 2020 higher than ATE in 2020
$>1$ $<1$ higher than the corresponding odd in 2020 lower than the corresponding odd in 2020 lower than ATE in 2020
As discussed in Section 5, not all values of $\Gamma_{0},\Gamma_{1}$ are meaningful and the calibration procedure, which produces the set $\mathcal{C}$ (i.e., the green area in Figure D.1) allows us to focus on values of $\Gamma_{0}$ and $\Gamma_{1}$ that are more interpretable.
Appendix E Supplementary Materials for the Ad Effect in Pennsylvania
E.1 Additional Data Description
Our analysis consists of two datasets, the source data derived from the 2020 RCT data from Aggarwal et al., (2023) and the target data derived from the 2024 PA voter database. Prior to analysis, we recoded the shared covariates ${\bf V}_{i}$ from these two datasets for them to match. A description is provided as follows.
The age was coded as four groups (18-24, 25-24, 35-39, and 40+) in the 2020 RCT data and as date of birth in the 2020 PA voter database. For the target data, we calculated their age by the year of 2024 and excluded voters above 55 years’ old to match the range of age in Aggarwal et al., (2023), and then constructed a variable of age groups according to the source data. The resulting age group variable for analysis is a discrete variable with four levels.
For each voter, their party information from the 2020 RCT data was coded as one of the four levels: Democratic, Republican, Unknown and Other, whereas in the 2024 PA voter database was one of fifty choices including Democratic and Republican. For analysis, we constructed a party variable with three levels: Democratic, Republican, and Other/Unknown, whereas voters that didn’t belong to the first two levels would have their party level being “Other/Unknown”. We note that the party information from the 2020 RCT data was inaccurate with 72% being unknown and we refer readers to Aggarwal et al., (2023) for details.
The gender was coded in two levels (female and other) in the 2020 RCT data and three levels (female, male, unknown) in the 2024 voter database. Our gender variable for analysis has two levels: female and non-female where the non-female level includes voters whose gender weren’t coded as female.
The voting history information from the 2020 RCT was coded as ten binary variables. Each variable indicated whether a voter has voted in a specific year for every other year between 2000 and 2018 (i.e., voted in 2000, voted in 2002, voted in 2004, voted in 2006, voted in 2008, voted in 2010, voted in 2012, voted in 2014, voted in 2016, voted in 2018). The voting history information from the PA voter database differed across counties and the availability is provided in Figure E.1. Later, to check robustness of the estimation results with respect to the coding of voting history, we also repeated the analysis with two alternative ways of coding the voting history. The total three coding types are summarized below. Unless specified, the voting history was coded as in (1), i.e., following Aggarwal et al., (2023).
  1. (1)
    Voting history is coded as in Aggarwal et al., (2023), i.e., as ten binary variables indicating voting participation every two year from 2000 to 2018.
  2. (2)
    Voting history is coded as ten binary variables indicating voting participation 2/4/6$\cdots$/20 years ago.
  3. (3)
    Voting history is coded as two binary variables indicating voting participation in past presidential mid-term elections.
In addition to the common covariates, the 2020 RCT data also contains the race information, which is a categorical variable with four levels: White, Black, Latinx, and Other. Finally after pre-processing, the source covariates ${\bf X}_{i}$ include age group, gender, party, race, and ten binary variables indicating voting participation from 2000 to 2018, among which the common covariates ${\bf V}_{i}$ include age group, gender, party, and part of the voting history. The availability of covariates across counties in PA is provided in Figure E.1. Figure E.1 also provides the sample size $n_{t}$ across counties in the x-axis. Table E.1 summarizes the covariates (which are all discrete) and their levels.
Covariate Levels Available from target?
Age group 18-24, 25-34, 35-39, 40+ Yes
Gender Female, non-female Yes
Party Democratic, Republican, Other/Unknown Yes
Race White, Black, Latinx, other No
Voted in 2000 0, 1 No
Voted in 2002 0, 1 No
Voted in 2004 0, 1 Available in some counties
Voted in 2006 0, 1 Available in some counties
Voted in 2008 0, 1 Available in some counties
Voted in 2010 0, 1 Yes
Voted in 2012 0, 1 Yes
Voted in 2014 0, 1 Yes
Voted in 2016 0, 1 Yes
Voted in 2018 0, 1 Yes
Table E.1: Descriptions on covariates in pooled data.
Figure E.1: Target data availability across Pennsylvania counties in 2024. The y-axis lists the source covariate ${\bf X}$ and the points indicate the availability of these covariates in the target covariate ${\bf V}$ across counties. The x-axis is county with the number of voters (i.e., $n_{t}$) in parentheses.
E.2 Details for County-Level Ad Effects in Section 6.2
E.2.1 Numeric Values for Figure 3
We provide a comprehensive result (i.e., the specific numbers of confidence intervals) for the results presented in Section 6.2. Specifically, Table LABEL:tab:pa.county.ate.voted lists the ad effect estimated by the OR estimator under the three cases in parts A ($\Gamma_{0}=\Gamma_{1}=1$) and B ($(\Gamma_{0},\Gamma_{1})=(0.99,1.01)$ and $(\Gamma_{0},\Gamma_{1})=(1.01,0.99)$) of Figure 3 for each county in Pennsylvania. {longtblr}[ caption = County-by-county ad effect with the OR estimator in PA under $(\Gamma_{0},\Gamma_{1})=(1,1)$, $(\Gamma_{0},\Gamma_{1})=(0.99,1.01)$, and $(\Gamma_{0},\Gamma_{1})=(1.01,0.99)$. Each cells lists the TATE with $95\%$ CI in parentheses. , label = tab:pa.county.ate.voted, ] colspec = X— X XX, County $\Gamma_{0}=\Gamma_{1}=1$ $\Gamma_{0}=0.99,\Gamma_{1}=1.01$ $\Gamma_{0}=1.01,\Gamma_{1}=0.99$
Adams -0.41 (-1.05, 0.22) -0.04 (-0.68, 0.59) -0.79 (-1.42, -0.16)
Allegheny 0.06 (-0.55, 0.67) 0.41 (-0.2, 1.03) -0.29 (-0.9, 0.32)
Armstrong -0.6 (-1.3, 0.1) -0.25 (-0.95, 0.45) -0.95 (-1.64, -0.25)
Beaver -0.26 (-0.86, 0.34) 0.09 (-0.51, 0.69) -0.61 (-1.2, -0.01)
Bedford -0.77 (-1.55, 0) -0.44 (-1.21, 0.34) -1.11 (-1.88, -0.34)
Berks -0.18 (-0.76, 0.4) 0.18 (-0.4, 0.76) -0.54 (-1.11, 0.04)
Blair -0.52 (-1.19, 0.15) -0.17 (-0.84, 0.5) -0.87 (-1.53, -0.2)
Bradford -0.57 (-1.26, 0.12) -0.21 (-0.9, 0.48) -0.94 (-1.63, -0.25)
Bucks -0.13 (-0.71, 0.46) 0.22 (-0.36, 0.81) -0.48 (-1.06, 0.11)
Butler -0.42 (-1.06, 0.22) -0.07 (-0.71, 0.57) -0.78 (-1.42, -0.13)
Cambria -0.4 (-1.05, 0.24) -0.05 (-0.7, 0.59) -0.75 (-1.39, -0.11)
Cameron -0.54 (-1.24, 0.16) -0.17 (-0.87, 0.53) -0.91 (-1.6, -0.21)
Carbon -0.34 (-0.97, 0.28) 0.04 (-0.59, 0.67) -0.72 (-1.34, -0.1)
Centre -0.06 (-0.62, 0.51) 0.32 (-0.25, 0.89) -0.43 (-1, 0.13)
Chester -0.05 (-0.62, 0.52) 0.3 (-0.28, 0.87) -0.4 (-0.97, 0.17)
Clarion -0.57 (-1.28, 0.13) -0.23 (-0.94, 0.48) -0.91 (-1.62, -0.21)
Clearfield -0.55 (-1.24, 0.14) -0.2 (-0.89, 0.49) -0.9 (-1.59, -0.21)
Clinton -0.45 (-1.1, 0.21) -0.08 (-0.74, 0.58) -0.81 (-1.47, -0.16)
Columbia -0.32 (-0.93, 0.3) 0.05 (-0.56, 0.67) -0.69 (-1.3, -0.07)
Crawford -0.45 (-1.13, 0.22) -0.1 (-0.78, 0.58) -0.81 (-1.48, -0.13)
Cumberland -0.17 (-0.77, 0.43) 0.19 (-0.41, 0.79) -0.53 (-1.13, 0.06)
Dauphin 0.04 (-0.56, 0.63) 0.41 (-0.19, 1) -0.33 (-0.93, 0.26)
Delaware 0.02 (-0.59, 0.64) 0.38 (-0.24, 1) -0.33 (-0.95, 0.29)
Elk -0.51 (-1.19, 0.16) -0.15 (-0.83, 0.53) -0.87 (-1.55, -0.2)
Erie -0.19 (-0.78, 0.39) 0.16 (-0.43, 0.74) -0.54 (-1.12, 0.04)
Fayette -0.39 (-1.04, 0.26) -0.02 (-0.67, 0.63) -0.76 (-1.4, -0.11)
Forest -0.61 (-1.34, 0.12) -0.26 (-0.99, 0.47) -0.97 (-1.7, -0.24)
Franklin -0.52 (-1.18, 0.15) -0.16 (-0.83, 0.51) -0.87 (-1.54, -0.21)
Fulton -0.84 (-1.64, -0.04) -0.5 (-1.3, 0.31) -1.19 (-1.98, -0.39)
Greene -0.49 (-1.15, 0.18) -0.13 (-0.8, 0.54) -0.84 (-1.51, -0.17)
Huntingdon -0.56 (-1.26, 0.15) -0.21 (-0.91, 0.5) -0.91 (-1.61, -0.2)
Indiana -0.39 (-1.04, 0.26) -0.02 (-0.68, 0.63) -0.75 (-1.39, -0.1)
Jefferson -0.63 (-1.36, 0.1) -0.27 (-1, 0.46) -0.99 (-1.71, -0.26)
Juniata -0.7 (-1.46, 0.05) -0.38 (-1.13, 0.38) -1.03 (-1.78, -0.28)
Lackawanna -0.09 (-0.71, 0.53) 0.27 (-0.35, 0.89) -0.44 (-1.06, 0.17)
Lancaster -0.3 (-0.9, 0.29) 0.04 (-0.55, 0.64) -0.65 (-1.24, -0.06)
Lawrence -0.39 (-1.02, 0.25) -0.03 (-0.66, 0.61) -0.75 (-1.39, -0.12)
Lebanon -0.35 (-0.98, 0.27) 0.01 (-0.61, 0.64) -0.72 (-1.34, -0.1)
Lehigh -0.02 (-0.59, 0.56) 0.36 (-0.21, 0.94) -0.39 (-0.97, 0.18)
Luzerne -0.19 (-0.79, 0.42) 0.17 (-0.43, 0.78) -0.55 (-1.15, 0.05)
Lycoming -0.45 (-1.09, 0.2) -0.1 (-0.74, 0.55) -0.8 (-1.44, -0.15)
McKean -0.54 (-1.22, 0.15) -0.16 (-0.85, 0.52) -0.91 (-1.59, -0.23)
Mercer -0.41 (-1.03, 0.21) -0.05 (-0.67, 0.57) -0.77 (-1.38, -0.15)
Mifflin -0.6 (-1.33, 0.12) -0.26 (-0.98, 0.47) -0.95 (-1.67, -0.23)
Monroe 0.06 (-0.53, 0.65) 0.46 (-0.13, 1.05) -0.34 (-0.92, 0.25)
Montgomery 0.03 (-0.58, 0.64) 0.38 (-0.22, 0.99) -0.32 (-0.93, 0.28)
Montour -0.25 (-0.87, 0.37) 0.14 (-0.48, 0.76) -0.63 (-1.25, -0.01)
Northampton -0.01 (-0.58, 0.56) 0.37 (-0.2, 0.94) -0.4 (-0.97, 0.17)
Northumberland -0.39 (-1.04, 0.26) -0.02 (-0.67, 0.63) -0.76 (-1.41, -0.11)
Perry -0.5 (-1.22, 0.22) -0.15 (-0.87, 0.58) -0.86 (-1.58, -0.14)
Philadelphia 0.38 (-0.35, 1.11) 0.75 (0.02, 1.48) 0.01 (-0.71, 0.74)
Pike -0.21 (-0.82, 0.39) 0.19 (-0.41, 0.79) -0.62 (-1.22, -0.01)
Potter -0.67 (-1.43, 0.09) -0.32 (-1.08, 0.44) -1.03 (-1.79, -0.27)
Schuylkill -0.43 (-1.08, 0.21) -0.07 (-0.72, 0.58) -0.79 (-1.44, -0.15)
Snyder -0.58 (-1.27, 0.12) -0.23 (-0.93, 0.47) -0.92 (-1.62, -0.23)
Somerset -0.64 (-1.35, 0.07) -0.31 (-1.02, 0.41) -0.97 (-1.68, -0.26)
Sullivan -0.53 (-1.26, 0.2) -0.17 (-0.9, 0.56) -0.88 (-1.61, -0.16)
Susquehanna -0.37 (-1.08, 0.33) -0.01 (-0.72, 0.7) -0.73 (-1.44, -0.03)
Tioga -0.55 (-1.25, 0.15) -0.19 (-0.89, 0.51) -0.91 (-1.61, -0.21)
Union -0.28 (-0.86, 0.31) 0.1 (-0.48, 0.69) -0.65 (-1.23, -0.07)
Venango -0.51 (-1.19, 0.16) -0.16 (-0.83, 0.51) -0.87 (-1.54, -0.2)
Warren -0.45 (-1.13, 0.22) -0.09 (-0.76, 0.58) -0.82 (-1.49, -0.15)
Washington -0.35 (-0.99, 0.28) 0.01 (-0.62, 0.65) -0.72 (-1.35, -0.09)
Wayne -0.41 (-1.08, 0.25) -0.05 (-0.71, 0.62) -0.78 (-1.44, -0.12)
Westmoreland -0.37 (-1, 0.27) -0.02 (-0.65, 0.62) -0.72 (-1.35, -0.09)
Wyoming -0.5 (-1.18, 0.17) -0.15 (-0.83, 0.52) -0.85 (-1.53, -0.18)
York -0.33 (-0.93, 0.28) 0.03 (-0.58, 0.63) -0.68 (-1.29, -0.08)
E.2.2 Additional Example Sensitivity Parameters
To supplement part B of Figure 3, Figure E.2 plot the conclusions under additional choices of $(\Gamma_{0},\Gamma_{1})$s.
Figure E.2: Conclusions of ad effects with given sensitivity parameters.
E.3 Comparison Between Estimators
In this section, we compare results from the the two estimators: the OR estimator as presented in Section 6.2 of the main text and the EIF-based estimator. Figure E.3 compares the results under transportability across the three ways of coding the voting history (see Section E.1 for the coding of voting history). In each panel, the x- and y-axes represent the results from the OR estimator and the EIF estimator; we find the points lie around the $y=x$ line which indicates that the point estimates are close. The CIs also have comparable lengths. Numeric values of the estimates (CIs) are given in Figures E.4 and E.6 for the OR and EIF-based estimators, respectively.
Figure E.3: Comparison between the OR and EIF-based estimators for estimating ad effects for every PA county under transportability. Each panel represents one way of coding the voting history; the left panel corresponds to the way presented in the main text. In each panel, x- and y- axes represent results from the OR and the EIF-based estimator, respectively. The points represent point estimates and gray bars represent 95% CIs. The dashed line represents $x=y$.
We zoom into the first way of coding the voting history, which corresponds to Aggarwal et al., (2023) and the main text; results for the other two ways are presented in Sections E.4 and E.5 for the OR and EIF-based estimators, respectively. When $\Gamma_{0}=\Gamma_{1}=1$, the EIF-based estimator yields no significant results while the OR estimator yields a significant and negative effect in Fulton. The conservative result from the EIF-based estimator is due to the small sample size of voters in the Fulton county ($n_{t}=4746$): the small $n_{t}$ leads to finite-sample violations to Assumption 2.2 and difficulty in estimating the density ratio $w({\bf V})$, which in turn yields higher variance estimates. We note that this phenomenon only happen for a few small counties, and in general, the widths of CIs for one estimator do not uniformly dominate the other.
When $\Gamma_{0}\neq 1$ or $\Gamma_{1}\neq 1$, we apply the calibration procedure for each estimator following the same procedure as in Section 6.2. Within the calibrated region, the analysis based on the EIF-based estimator produced nine more counties sensitive for a positive effect and two fewer counties sensitive for a negative effect than those in Section 6; see Figures E.5 and E.7.
E.4 Robustness Checks on County Level Ad Effects for the OR Estimator
In this section, we check robustness of the OR estimator across three ways of coding the voting history mentioned in Section E.1. From Figure E.4, we find the 95% CIs under transportability are similar across ways of coding the voting history with slight differences. The second way of coding the voting history gives no significance while the third way yields significance for Fulton, Potter and Bedford counties which all have a negative effect.
Figure E.5 provides the conclusions under transportability on the top panels and the changes in conclusions within the calibrated sensitivity analysis on the bottom panels. We note that the top left and bottom left panels are parts A and D of Figure 3 in the main text. For the calibrated sensitivity analysis, the results of 51 counties are the same across the three ways of coding voting history. It’s notable that the Philadelphia county continues to be only county sensitive for a positive effect. For the other 16 counties, discrepancy occurs mainly due to the differences in significance under transportability. It’s notable that the Monroe county can be insensitive, sensitive for a negative effect, sensitive for a positive effect depending on the way of coding the voting history.
Figure E.4: Robustness checks of the OR estimator with respect to the voting history coding types. Each panel plots the 95% CI under transportability on the x-axis with color indicating significance.
Figure E.5: Robustness checks of the OR estimator with respect to the voting history coding types. The top row represents results under transportability and the bottom row represents the change of conclusions after the calibrated sensitivity analysis.
E.5 Robustness Checks on County Level Ad Effects for the EIF-Based Estimator
In this section, we check robustness of the EIF-Based estimator across three ways of coding the voting history mentioned in Section E.1. From Figure E.6, we find the 95% CIs under transportability are similar across ways of coding the voting history. The ad effects are insignificant except for one case: the ad effect is significant and negative in Juniata county in the second way of coding the voting history.
Figure E.7 provides the conclusions under transportability on the top panels and the changes in conclusions within the calibrated sensitivity analysis on the bottom panels. For the calibrated sensitivity analysis, the results are consistent in 45 counties across ways of coding voting history. In the 21 counties with inconsistent results, we find the highest discrepancy in the Centre county and the Chester county. The Centre county is sensitive to a significant effect of either sign, to a significant and negative effect, and insensitive in the three ways of coding. The Chester county is sensitive to a significant and positive effect, to a significant and negative effect, and insensitive in the three cases.
Figure E.6: Robustness checks of the EIF estimator with respect to the voting history coding types. Each panel plots the 95% CI under transportability on the x-axis with color indicating significance.
Figure E.7: Robustness checks of the EIF estimator with respect to the voting history coding types. The top row represents results under transportability and the bottom row represents the change of conclusions after the calibrated sensitivity analysis.
E.6 Details on Data Pre-Processing for Subgroup Analysis
This section details the construction of variables regarding urbanicity and education attainment in the subgroup analysis presented in Section 6.3.
E.6.1 Percentage of Bachelor’s Degree Or Higher in ZIP Codes
To construct a variable as a proxy for a voter’s education attainment, we leverage the ZIP code information from the PA voter database. In specific, for every ZIP code in the PA voter database, we calculate the percentage of receiving a Bachelor’s degree or higher from the 2022 American Community Survey (ACS), which is a comprehensive census that represents the U.S. population. To preserve privacy, we excluded ZIP codes with fewer than 20 voters from the PA voter database or from the ACS data. This step removed 146 ZIP codes and 2149 voters. As a result, for each voter, we have the percentage of Bachelor’s degree or higher in their ZIP-code area. And for analysis, we divided the percentages into five groups by every 20 percent.
E.6.2 Urbanicty in Census Tracts
For urbanicity, we mapped a voter’s address with the 2020 U.S. census which classifies a census tract as urban or rural (i.e., not urban) based on characteristics including population, housing, and land area among others. We refer readers to the U.S. Census Bureau’s urban-rural classification for the criterion of classifying a census tract as urban or rural. Among all 4,880,729 voters, the addresses of 176,866 (0.04%) cannot be matched with a census tract. Their urbanicity was imputed by the proportion of urban voters with the same ZIP code (if the proportion is less than 50%, we imputed the urbanicity to be rural and vice versa), except for 1,147 whose urbanicity cannot be imputed because their ZIP codes are either missing or do not match with ZIP codes of other voters. These voters take 0.02% of the original voters and have been excluded from the analysis in Section 6.3.
E.7 Robustness Checks on Subgroup Analysis in Section 6.3
In this section, we provide the estimated ad effects in subgroups from both estimators across three ways of coding the voting history. Figure E.8 presents the estimation results of both estimators by the interaction between gender, urbanicity, and education attainment in 2022. Under the transportability assumption, point estimates and 95% confidence intervals by the EIF-based estimator is close to those by the OR estimator presented in the main text. The calibrated results by the EIF-based estimator also mostly coincide with the OR estimator.
Figure E.8: Ad effect estimates in subgroups defined by the interaction between gender, urbanicity, and percentage of Bachelor’s degree within the same ZIP-code area using both OR estimator and the EIF-based estimator. The results of the OR estimator are exactly those in Figure 5.
Appendix F Replication of Section 6 When Excluding Three States from the Source
In this section, we repeat the analysis in Section 6 when excluding voters from Pennsylvania (PA), Michigan (MI), and Wisconsin (WI) from the source data. The source data now consists of $n_{s}=662225$ voters from North Carolina (NC) and Arizona (AZ) in the 2020 experiment by Aggarwal et al., (2023). Table F.1 summarizes the voter demographics and turnout in (PA, MI, WI) and (NC, AZ).
Table F.1: Voter demographics for (PA, MI, WI) and (NC, AZ) in the RCT data from Aggarwal et al., (2023).
States (PA, MI, WI) (NC, AZ)
Size 1337057 662225
Gender = Other (%) 638840 (47.8) 382401 (57.7)
Age groups (%)
   18-24 288783 (21.6) 224352 (33.9)
   25-34 430123 (32.2) 208818 (31.5)
   35-39 161576 (12.1) 65365 ( 9.9)
   40+ 456575 (34.1) 163690 (24.7)
Party (%)
   Democrat 108810 ( 8.1) 74135 (11.2)
   Other 1195792 (89.4) 548670 (82.9)
   Republican 32455 ( 2.4) 39420 ( 6.0)
Voted in 2020 = Yes (%) 761181 (56.9) 329639 (49.8)
Sections F.1 and F.2 provide results for a county-by-county analysis and a subgroup analysis, respectively, which mirror Sections 6.2 and 6.3 of the main text.
F.1 Ad Effects by Counties
Figures F.1 and F.2 plot the results from the OR estimator and the EIF-based estimator, respectively. When $\Gamma_{0}=\Gamma_{1}=1$, i.e., under transportability, the ad effect is insignificant in all counties of PA for both estimators. When $\Gamma_{0}\neq 1$ or $\Gamma_{1}\neq 1$, after calibration, the ad effect is sensitive for a negative effect in 42 counties from the OR estimator and 10 counties from the EIF estimator. Results in the other counties are insensitive. We note that the result while restricting the source with data from NC, AZ alone give more conservative result than Section 6 due to the smaller sample size of the source population.
Figure F.1: County-by-county analysis with the OR estimator when the source data only consists of voters in NC and AZ. The left panel represents the insignificance of the result under transportability. In the right panel, the red downward arrow represents counties sensitive to a signifcant and negative effect.
Figure F.2: County-by-county analysis with the EIF estimator when the source data only consists of voters in NC and AZ. The left panel represents the insignificance of the result under transportability. In the right panel, the red downward arrow represents counties sensitive to a signifcant and negative effect.
F.2 Subgroup Analysis
We estimate the ad effect in 20 subgroups of gender, urbanicity, and education attainment for voters within the same ZIP-code area. Results are shown in Figure F.3. When $\Gamma_{1}=\Gamma_{0}=1$, i.e., under transportability, the effects are in general higher for non-female voters than female voters, and higher for urban voters than rural voters. When $\Gamma_{0}\neq 1$ or $\Gamma_{1}\neq 1$, after calibration, none of the subgroups are sensitive for a significant ad effect.
Figure F.3: Subgroup analysis by the interaction between gender, urbanicity, and education subgroups while including only the NC and AZ voters in the source data. The solid lines and dashed lines represent 95% CIs under transportability for the OR estimator and the EIF-based estimator, respectively. The gray squares represent that all subgroups are insensitive for a significant effect.
Appendix G Simulations
In this section, we validate asymptotic properties of our proposed estimators on simulated datasets generated according to the 2020 RCT data.
In order to generate data that mimics the 2020 RCT data, we let the source covariate ${\bf X}_{i}$ be gender, race, and age groups and set its distribution ${\bf X}_{i}\mid S_{i}=1$ to be the empirical distribution of these covariates in the 2020 RCT data. Given ${\bf x}\in{\cal X}$, the treatment is randomized within 18 strata mimicking the design in Aggarwal et al., (2023). The $\mu_{1}({\bf x})$ and $\mu_{0}({\bf x})$ are generated in two scenarios. In Scenario (A), they differ by $0.005$ or $-0.005$ whereas the overall average effect is close to zero, mimicking the real data where the overall ad effect is negligible despite small, heterogeneous effects in subgroups. In Scenario (B), the difference between $\mu_{1}({\bf x})$ and $\mu_{0}({\bf x})$ is larger in magnitude and more heterogeneous. The covariate distribution on the target population, $p_{{\bf X}\mid S=0}$ is generated such that $p_{{\bf X}\mid S=0}({\bf x})/p_{{\bf X}\mid S=1}({\bf x})$ is between $0.9$ and $1.1$. Table G.1 presents the values of this generation. The target covariate ${\bf V}_{i}$ is set to be the gender variable alone. The sensitivity parameter $\gamma_{0}$ is set to zero and $\gamma_{1}$ varies. The source sample size $n_{t}$ and target sample size $n_{t}$ are set equal.
Table G.1: Data generation in simulated datasets.
Gender Race Age Group $p_{{\bf X}\mid S=1}({\bf x})$ $p_{{\bf X}\mid S=0}({\bf x})$ $\pi({\bf x})$ Scenario (A) Scenario (B)
$\mu_{0}({\bf x})$ $\mu_{1}({\bf x})$ $\mu_{0}({\bf x})$ $\mu_{1}({\bf x})$
Female Black 18-24 0.0061 0.0055 0.6 0.4 0.35 0.2 0.6
Female Black 25-34 0.0077 0.0071 0.7 0.4 0.35 0.2 0.6
Female Black Other 0.0157 0.0150 0.8 0.5 0.45 0.7 0.2
Female Latinx 18-24 0.0073 0.0066 0.6 0.5 0.45 0.7 0.2
Female Latinx 25-34 0.0089 0.0083 0.8 0.4 0.35 0.3 0.3
Female Latinx Other 0.0147 0.0139 0.9 0.5 0.45 0.7 0.2
Female Other 18-24 0.1001 0.1042 0.6 0.6 0.55 0.3 0.5
Female Other 25-34 0.1271 0.1353 0.8 0.5 0.45 0.6 0.2
Female Other Other 0.2016 0.2218 0.9 0.6 0.55 0.3 0.5
Other Black 18-24 0.0197 0.0193 0.6 0.3 0.35 0.2 0.6
Other Black 25-34 0.0280 0.0285 0.8 0.2 0.25 0.2 0.6
Other Black Other 0.0397 0.0409 0.8 0.3 0.35 0.2 0.6
Other Latinx 18-24 0.0174 0.0169 0.6 0.3 0.35 0.25 0.55
Other Latinx 25-34 0.0201 0.0200 0.8 0.3 0.35 0.25 0.55
Other Latinx Other 0.0211 0.0212 0.9 0.4 0.45 0.25 0.55
Other Other 18-24 0.1061 0.1118 0.7 0.5 0.55 7 0.2
Other Other 25-34 0.1277 0.1375 0.8 0.4 0.45 0.25 0.55
Other Other Other 0.1310 0.1425 0.9 0.5 0.55 0.7 0.2
After generating datasets, the propensity score $\pi({\bf x})$ is estimated with the average proportion of treated units within each. The outcome regression functions $\mu_{a}({\bf x})$ and $\rho_{a}({\bf v})$ are estimated by reweighing samples with $S_{i}=1$ and $A_{i}=a$ as in (5). The density ratio $w({\bf v})$ is estimated with (C.2). For the OR estimator, the inference is based on 1000 bootstrap iterations. For the EIF-based estimator, the inference is based on the cross-fitting procedure with $K=2$ splits. The confidence level is set to $0.951-\alpha=$ 0.95. Simulation results are based on 1000 replicates.
From results in Table G.2, both estimators are consistent and their empirical standard deviation (SD) decays with $\sqrt{n}$. The estimated SEs are close to the empirical SDs and the coverage rate nears the nominal level $0.95$. These results validate bootstrap CI consistency in Theorem 4.1 as well as the asymptotic Normality of the EIF-based cross-fitting estimator in Theorem 4.2.
Table G.2: Simulation results. Bias, RMSE, empirical standard deviation (Emp.SD) and estimated standard error (Est.SE) have been multiplied with $1000$.
$\gamma_{1}=1$ Scenario (A) Scenario (B)
Estimator $n_{s}(=n_{t})$ Bias RMSE Emp.SD Est.SE Rate Bias RMSE Emp.SD Est.SE Rate
OR $10^{5}$0.135 4.317 4.317 4.275 0.943 0.076 4.169 4.171 4.123 0.952
OR 2×$\times 10^{5}$0.126 3.047 3.046 3.018 0.953 0.030 2.951 2.952 2.913 0.939
EIF $10^{5}$ 0.004 4.307 4.309 4.283 0.953 -0.0.082 3.943 3.944 4.135 0.955
EIF 2×$\times 10^{5}$0.008 3.029 3.030 3.024 0.953 -0.549 2.996 2.947 2.920 0.945
$\gamma_{1}=1.05$ Scenario (A) Scenario (B)
OR $10^{5}$0.136 4.318 4.318 4.276 0.945 0.076 4.178 4.180 4.130 0.953
OR 2×$\times 10^{5}$0.126 3.047 3.046 3.019 0.954 0.029 2.957 2.958 2.920 0.940
EIF $10^{5}$ 0.225 4.356 4.357 4.283 0.947 -0.265 4.152 4.145 4.142 0.948
EIF 2×$\times 10^{5}$ 0.095 3.028 3.028 3.024 0.943 -0.483 2.984 2.946 2.925 0.941